### v1

# A Perfectly Truthful Calibration Measure

Abstract
Calibration requires that predictions are conditionally unbiased and, therefore, reliably interpretable as probabilities. Calibration measures quantify how far a predictor is from perfect calibration. As introduced by Haghtalab et al. (2024), a calibration measure is truthful if it is minimized in expectation when a predictor outputs the ground-truth probabilities. Although predicting the true probabilities guarantees perfect calibration, in reality, when calibration is evaluated on a finite sample, predicting the truth is not guaranteed to minimize any known calibration measure. All known calibration measures incentivize predictors to lie in order to appear more calibrated on a finite sample. Such lack of truthfulness motivated Haghtalab et al. (2024) and Qiao and Zhao (2025) to construct approximately truthful calibration measures in the sequential prediction setting, but no perfectly truthful calibration measure was known to exist even in the more basic batch setting.
We design a perfectly truthful calibration measure in the batch setting: averaged two-bin calibration error (ATB). In addition to being truthful, ATB is sound, complete, continuous, and quadratically related to two existing calibration measures: the smooth calibration error smCal and the (lower) distance to calibration $\underline{\textsc{distCal}}$. The simplicity in our definition of ATB makes it efficient and straightforward to compute. ATB allows faster estimation algorithms with significantly easier implementations than smCal and $\underline{\textsc{distCal}}$, achieving improved running time and simplicity for the calibration testing problem studied by Hu et al. (2024). We also introduce a general recipe for constructing truthful measures based on the variance additivity of independent random variables, which proves the truthfulness of ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
1 Introduction
Probabilistic forecasting has become increasingly important in modern AI-assisted decision-making. Unlike deterministic classification, probabilistic forecasts provide uncertainty quantification, allowing assessment of risks. One desired property of probabilistic predictions is calibration, which requires predictions to be conditionally unbiased and, therefore, reliably interpretable as probabilities. For example, neural networks for tumor diagnosis are trained to output a prediction $r\in[0,1]$, ideally interpretable as the probability of a binary state $y\in\{0,1\}$: the tumor segment being malignant or not. The neural network is calibrated if, conditioned on the output being, say $40\%$, the probability that the tumor is malignant is also $40\%$.
[…]
There is a recent line of work studying the truthfulness of calibration measures (Haghtalab et al., 2024; Qiao and Zhao, 2025) in the sequential prediction setting. We focus on the more basic batch setting, where a calibration measure is truthful if it is minimized, in expectation over a random sample, when a predictor outputs the ground truth probabilities (we term this predictor the truthful predictor). Ideally, to assess calibration, a calibration measure should rank a calibrated predictor with the lowest expected error. This would imply that the calibration measure is truthful: the truthful predictor is always calibrated and should achieve minimum calibration error. However, no known calibration measure is truthful. It is true that assuming knowledge of the full joint distribution of the prediction and the state, all known calibration measures reach the minimum 0 if and only if a predictor is perfectly calibrated, but in practice, calibration error is evaluated on random samples rather than the unobserved underlying distribution. On a random sample of arbitrary size, for any known calibration measure, predicting the true probabilities does not guarantee minimum expected error, and a lower expected error does not guarantee that the predictor is more calibrated.
1.1 Non-truthfulness of Known Calibration Measures
We illustrate the non-truthfulness of ECE in Example 1.1 below, with a discussion afterwards showing that the same non-truthfulness issue occurs in a broad family of calibration measures. We consider the batch setting: a sample of $T$ individuals whose binary states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ are independently drawn from the Bernoulli distributions with means $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ (denoted by $\bm{y}\sim\bm{p}$). We say $\bm{p}=(p_{1},\ldots,p_{T})$ are the ground-truth probabilities. In the example below, a predictor strictly benefits from reporting less informative, pooled predictions than the ground-truth probabilities.
Example 1.1 (ECE and binned ECE are not truthful, c.f. Qiao and Valiant, 2021).
Suppose the ground-truth probabilities are $\bm{p}=(p_{1},\dots,p_{T})$ where each $p_{t}$ is distributed independently and uniformly over $[1/3,2/3]$. An uninformative predictor that always predicts $r_{1}=\cdots=r_{T}=0.5$ achieves an expected empirical $\textsc{ECE}=O\left(\sqrt{1/T}\right)$, the sampling error. However, a truthful predictor who reports $r_{t}=p_{t}$ results in a higher empirical $\textsc{ECE}\geq 1/3$. This is because the predictions $r_{1},\ldots,r_{T}\in[1/3,2/3]$ are almost surely distinct, so the empirical conditional expectation $\widehat{r}_{t}:=\operatorname{\mathbb{E}}_{(r,y)\sim\mathrm{Unif}((r_{t},y_{t})_{t\in[T]})}[y|r=r_{t}]$ is simply $y_{t}\in\{0,1\}$, giving $|r_{t}-\widehat{r}_{t}|=|r_{t}-y_{t}|\geq 1/3$.
The same observation holds for binned ECE as well. Suppose the prediction space of $[0,1]$ is discretized into $k$ equal-length intervals (i.e., bins). When the true probabilities are predicted, on average, each interval has roughly $\frac{T}{k}$ predictions, inducing a sampling error of roughly $\sqrt{k/T}$, higher than the binned ECE $O\left(\sqrt{1/T}\right)$ achieved by the uninformative predictor $r_{1}=\cdots=r_{T}=0.5$.
Example 1.1 shows that the truthful predictor $\bm{r}=\bm{p}$ does not achieve a minimum ECE or binned ECE on a sample of size $T$. Even worse, there are miscalibrated predictors (e.g. the predictor that always predicts $0.5+\epsilon$ for a small $\epsilon>0$) achieving smaller ECE and binned ECE than the calibrated truthful predictor. Thus ECE and binned ECE do not rank predictors correctly based on how calibrated they are.
The observation in Example 1.1 that predicting the truth incurs higher (i.e. worse) calibration error than (strategically) assigning the same prediction (i.e. the overall average) to every individual is, unfortunately, very general. It happens not just for ECE and its variants [^1] Variants of ECE include $\ell_{\alpha}$-ECE, where we replace the absolute bias |rr^||r-$\widehat${r}| with $|r-\widehat{r}|^{\alpha}$ for an arbitrary $\alpha\geq 1$, as well as binned versions of $\ell_{\alpha}$-ECE., but also for continuous calibration measures [^2] These are calibration measures that are continuous as a function of the predictions. Note that ECE and binned ECE are not continuous. such as the smooth calibration error (Kakade and Foster, 2008), the distance to calibration (Błasiok et al., 2023) and its variants, etc, irrespective of the sample size $T$ used for estimating the calibration error. Moreover, it happens consistently accross every realization of the states $\bm{y}$, not just in expectation. Specifically, for every realization of states $\bm{y}\in\{0,1\}^{T}$ and every prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$, all these calibration measures Cal satisfy
$\textsc{Cal}(\bm{r},\bm{y})\geq\textsc{Cal}(\bar{\bm{r}},\bm{y}),$ (1)
where $\bar{\bm{r}}=(\bar{r},\ldots,\bar{r})\in[0,1]^{T}$ is the constant predictor that always predicts $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. This means that regardless of the (unknown) realized states $\bm{y}$, the constant predictor $\bar{\bm{r}}$ consistently achieves smaller or equal calibration error than $\bm{r}$, even when $\bm{r}$ is the true underlying probabilities $\bm{p}$. We formally prove this observation in Theorem A.2.
It is very easy to find (many) examples of $\bm{y}$ where the inequality (1) becomes strict, in which case reporting $\bar{\bm{r}}$ instead of the truth $\bm{r}=\bm{p}$ is strictly better. In particular, we can find many examples of $\bm{y}$ where $\textsc{Cal}(\bm{r},\bm{y})>0$ and $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$. This is because for all of the calibration measures Cal mentioned above (including the continuous ones), we always have $\textsc{Cal}(\bm{r},\bm{y})>0$ as long as $r_{1},\ldots,r_{T}$ are distinct values in, say, $[1/3,2/3]$, and we have $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$ as long as the average outcome $\frac{1}{T}\sum_{t=1}^{T}y_{t}$ is equal to $\bar{r}$.
1.2 Problem Statement
Measuring and optimizing for calibration non-truthfully only makes the predictions less trustworthy, going in the very opposite direction of the goal of calibration. Let $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be the ground-truth probabilities. We say a calibration measure Cal is truthful if outputting $\bm{p}$ minimizes the expected empirical calibration error, i.e., the minimizer to $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ is $\bm{r}=\bm{p}$, for every $\bm{p}\in[0,1]^{T}$. From a machine learning perspective, a truthful measure helps identify the Bayesian optimal predictor (Gneiting and Raftery, 2007) because it correctly ranks ground-truth predictions with the lowest expected error. From a game-theoretic perspective, a truthful measure incentivizes an optimizing predictor to output their true beliefs, where we view $\bm{p}$ as the predictor’s subjective belief about the probabilities, which might differ from the true probabilities. For the tumor risk prediction task, if assessed by a non-truthful calibration measure, a doctor is incentivized to report a prediction different from their true probabilistic assessment of the tumor risk, to make the predictions look more calibrated. In this case, an incentivized misreport can hardly be trusted.
In the sequential prediction setting, where the states $y_{1},\ldots,y_{T}$ are revealed sequentially after each prediction $r_{t}$ is made, Haghtalab et al. (2024) and Qiao and Zhao (2025) design approximately truthful calibration errors, where predicting the true probabilities approximately minimizes the expected error. The simpler batch setting, where all predictions are made before the states are revealed, is left unexplored. We observe that, in the batch setting, some existing measures, such as the smooth calibration error (Kakade and Foster, 2008; Błasiok et al., 2023) and the calibration measures proposed by Haghtalab et al. (2024) and Qiao and Zhao (2025), are approximately truthful. Yet, no known calibration measure is perfectly truthful.
Our goal is to design perfectly truthful calibration measures in the batch setting. Such measures correctly rank the truthful predictor with the lowest expected error and incentivize a predictor to output their true beliefs about the state distributions.
Applying a truthful error metric alone does not solve our problem. A calibration measure should, at a minimum, be able to distinguish calibrated predictors from miscalibrated ones given a sufficiently large sample. That is, it should satisfy basic properties, termed completeness and soundness (Definitions 2.8 and 2.9), requiring vanishing error when a predictor is calibrated (completeness) and non-vanishing error when a predictor is miscalibrated (soundness) as the sample size increases. Observed by Haghtalab et al. (2024), some error metrics, such as the well-known squared error $\frac{1}{T}\sum_{t}(r_{t}-y_{t})^{2}$, are truthful but far from being a complete and sound calibration measure.
As our main result, we construct a perfectly truthful, complete and sound calibration measure in the batch setting. Notably, our calibration measure provides a stronger rank-preserving guarantee than perfect truthfulness: the expected error over samples is the same as the idealized calibration error on the unknown true underlying distribution, up to a constant term independent of the predictor. In particular, it correctly ranks all calibrated predictors (not just the truthful one) with the same lowest expected error.
Moreover, we show that our calibration measure maintains and even improves other desiderata in the recent research on constructing good calibration measures beyond completeness, soundness, and truthfulness: consistency (Błasiok et al., 2023), continuity, low sample complexity, and high computational efficiency. Our calibration measure satisfies consistency in Błasiok et al., 2023, requiring it to be polynomially related to the Distance to Calibration, defined as the $\ell_{1}$ distance to the closest calibrated predictor. Our measure is continuous as a function of the predictions. It has the same low sample complexity $O(1/\varepsilon^{2})$ for estimation as the smooth calibration error. It has a very simple definition, making it more efficient to compute than existing continuous measures such as the smooth calibration error and the Distance to Calibration, achieving improved running time and simplicity for the calibraition testing problem studied by Hu et al. (2024). In summary, our construction achieves the best of all worlds while strictly improving many desired properties (e.g. truthfulness, computational efficiency, and simplicity).
1.3 Our Contributions
As our main contribution, we construct a perfectly truthful, complete and sound calibration measure: Averaged Two-Bin Calibration Error ($\mathrm{ATB}$). In addition, ATB satisfies all the other desired properties discussed earlier: continuity, low sample complexity, high computational efficiency, and the rank-preserving property. Moreover, it is quadratically related to two existing calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the (lower) Distance to Calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). This relationship implies that ATB is a consistent calibration measure in the framework of Błasiok et al. (2023).
As a bonus benefit of its simplicity and computational efficiency, ATB gives a faster algorithm for testing calibration compared to the previous algorithms of Hu et al. (2024). The $\ell_{1}$ version of ATB gives a simple constant-factor approximation to smCal and $\underline{\textsc{distCal}}$ while being more efficiently computable.
We also introduce a general framework for constructing truthful error metrics that encompasses ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
Perfectly Truthful Calibration Measure.
Our construction of $\mathrm{ATB}$ is very simple. We first divide the predictions into just two bins $[0,q)$ and $[q,1]$, with the bin boundary $q$ chosen uniformly at random from $[0,1]$. Within each bin, we compute the squared error between the sum of the predictions and the sum of the states, divided by $T^{2}$. After summing up the errors in the two bins, we define $\mathrm{ATB}$ to be the expectation over the random choice of the bin boundary $q$:
Definition 1.2 (Averaged two-bin calibration error).
Given predictions $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$, we define
$\mathrm{ATB}(\bm{r},\bm{y})$ $:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T^{2}}\left(\left(\sum_{t:r_{t}<q}(r_{t}-y_{t})\right)^{2}+\left(\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right)^{2}\right)\right].$
We show that ATB is truthful and continuous (in fact, Lipschitz continuous) as a function of the predictions $\bm{r}$:
Theorem 1.3 (Informal, Theorems 4.3 and 4.4).
$\mathrm{ATB}$ is a truthful and continuous calibration error. Here, truthful means predicting the ground truth $\bm{p}$ achieves lower or equal expected error than any other report vector $\bm{r}$:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]\quad\text{for every $\bm{p},\bm{r}\in[0,1]^{T}$}.$
Continuity means that for every two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$,
$|\mathrm{ATB}(\bm{r}_{1},\bm{y})-\mathrm{ATB}(\bm{r}_{2},\bm{y})|\leq\frac{6}{T}\|\bm{r}_{1}-\bm{r}_{2}\|_{1}.$
We show that ATB is quadratically related to two known continuous, complete and sound calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the lower distance to calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). We include formal definitions of smCal and $\underline{\textsc{distCal}}$ later in Definitions 2.5 and 2.4. Roughly speaking, the lower distance to calibration $\underline{\textsc{distCal}}(\bm{r},\bm{y})$ is the $\ell_{1}$ earthmover distance from the reported predictions $\bm{r}$ to the closest perfectly calibrated predictions. The smooth calibration error smCal is the following supremum over $1$-Lipschitz functions $w:[0,1]\to[-1,1]$:
$\textsc{smCal}(\bm{r},\bm{y})=\sup_{\text{$1$-Lipschitz }w:[0,1]\to[-1,1]}\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t}).$
Błasiok et al. (2023) showed that smCal and $\underline{\textsc{distCal}}$ are both continuous, complete and sound calibration measures, and they are constant-factor approximations of each other.
Theorem 1.4 (see Corollary 5.2).
For any sequences $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$,
$\frac{2}{9}\textsc{smCal}(\bm{r},\bm{y})^{2}$ $\leq\mathrm{ATB}(\bm{r},\bm{y})\leq 6\textsc{smCal}(\bm{r},\bm{y}),\text{\quad and\quad}$
$\frac{1}{18}\underline{\textsc{distCal}}(\bm{r},\bm{y})^{2}$ $\leq\mathrm{ATB}(\bm{r},\bm{y})\leq 3\underline{\textsc{distCal}}(\bm{r},\bm{y}).$
Theorem 1.4 implies that ATB is complete, sound, and is a consistent calibration measure in the framework of Błasiok et al. (2023).
We also show that ATB has desirable sample complexity and supreme computational complexity:
Theorem 1.5 (Informal, see Theorem 4.5 and Section 4.4).
ATB can be estimated within $\varepsilon$ error using $T=O(\varepsilon^{-2})$ examples, and can be computed in $O(T\log T)$ time.
The $O(T\log T)$-time algorithm for computing ATB follows directly from the simplicity of its definition. We need $O(T\log T)$ time to sort the predictions once, after which it only takes $O(T)$ time (see Section 4.4). For comparison, the best-known algorithms for computing smCal and $\underline{\textsc{distCal}}$ take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
More Than Truthfulness: the Rank-Preserving Property of ATB.
We show that ATB has the following stronger property that implies truthfulness. It allows us to compare predictors based on calibration on the sampled states and produce the same ranking as if calibration is evaluated on the unobserved underlying distribution:
Theorem 1.6 (Informal, see Theorem 4.3).
Let $\bm{p}\in[0,1]^{T}$ denote the true probabilities. For two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$, we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r}_{1},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r}_{2},\bm{y})]=\mathrm{ATB}(\bm{r}_{1},\bm{p})-\mathrm{ATB}(\bm{r}_{2},\bm{p}).$
Here, on the left-hand side, we compare the expected empirical error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]$ on the sampled states $\bm{y}$, whereas on the right-hand side, we compare the error $\mathrm{ATB}(\bm{r},\bm{p})$ w.r.t. the unobserved true probabilities $\bm{p}$. The theorem guarantees consistent ordering between these two ways of evaluating calibration. The truthfulness property in Theorem 1.3 then follows immediately as a consequence of Theorem 1.6 (and the basic fact that $\mathrm{ATB}(\bm{p},\bm{p})=0$):
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})-\mathrm{ATB}(\bm{p},\bm{p})\geq 0.$
In particular, Theorem 1.6 implies that any calibrated predictions $\bm{r}$ w.r.t. the true probabilities $\bm{p}$ (i.e. predictions with zero Distance to Calibration $\underline{\textsc{distCal}}(\bm{r},\bm{p})=0$, or equivalently $\mathrm{ATB}(\bm{r},\bm{p})=0$ due to their quadratic relationship) achieve the same minimum expected empirical error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]$ as predicting the truth, because
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{p},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})-\mathrm{ATB}(\bm{p},\bm{p})=0-0=0.$
The converse is also true and similarly implied by Theorem 1.6.
General Framework for Constructing Truthful Calibration Measures.
We establish the truthfulness of ATB by showing that it is a member of a general family of truthful error metrics, which we term the Unnormalized Binned Squared Errors (UBSEs) (Section 3).
We define UBSEs as binning-based calibration measures where the binning scheme can be randomized and can depend arbitrarily on the predictions $\bm{r}$. The key to its truthfulness lies in how the biases in each bin are combined to calculate the final error. We compute the squared biases in each bin similarly to the standard $\ell_{2}$ ECE, but there is a subtle but crucial difference in how these squared biases are weighted and combined. We discuss this family in more detail in Section 1.4.1 and Section 3. We prove that each UBSE has the following key property of error decomposition, which implies truthfulness as well as the rank-preserving property in Theorem 1.6:
Lemma 1.7 (Informal, Lemma 3.3).
[…]
Our construction of $\mathrm{ATB}$ is a special case of UBSE, and thus the truthfulness of $\mathrm{ATB}$ (Theorem 1.3) and the rank-preserving property (Theorem 1.6) follow immediately from Lemma 1.7. We remark that while every UBSE is truthful, it may not have the other desired properties of ATB, such as soundness, continuity, and the relationship to smCal and $\underline{\textsc{distCal}}$.
To illustrate the idea behind our definition of the UBSEs, we provide a technical overview in Section 1.4.1 and demonstrate, as an example, how to make binned ECE truthful using a quantile-based binning scheme, giving another truthful calibration measure Quantile $\ell_{2}$-BinECE.
Faster Optimal Calibration Test via Validity.
We introduce a quantitative framework for evaluating the validity of a calibration error using calibration tests. As the sample size increases, the notions of completeness and soundness require a calibration error to vanish given a calibrated predictor, and not vanish given a mis-calibrated predictor. We define the validity of a calibration error using its ability to distinguish calibration from mis-calibration, which can be viewed as a quantitative combination of completeness and soundness (see Definition 2.13). Any valid calibration measure implies an algorithm for solving the calibration testing problem in Hu et al. (2024).
We show that our $\mathrm{ATB}$ is optimally valid for the smooth calibration error and the lower distance to calibration.
Theorem 1.8 (Informal, see Theorem 6.1).
Given $T$ samples, $\mathrm{ATB}$ can distinguish a calibrated predictor from a predictor with $\textsc{smCal}\geq C/\sqrt{T}$ for an absolute constant $C>0$. Moreover, this rate is information-theoretically optimal (up to constant factors). This also holds when we replace smCal with $\underline{\textsc{distCal}}$ because the two are constant-factor approximations of each other as shown by Błasiok et al. (2023).
As mentioned earlier, ATB can be easily computed in $O(T\log T)$ time. Thus, Theorem 1.8 implies a faster calibration tester than the previous $O(T\log^{2}T)$-time calibration tester for smCal and $\underline{\textsc{distCal}}$ from Hu et al. (2024).
Simple and Efficient Constant-factor Approximation of smCal and $\underline{\textsc{distCal}}$.
We establish the quadratic relationship between $\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$ using the following $\ell_{1}$ variant of $\mathrm{ATB}$. We show that this variant gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$.
Definition 1.9.
We define the $\ell_{1}$ variant of $\mathrm{ATB}$:
$\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})$ $:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-y_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right|\right)\right].$
Theorem 1.10 (Informal, Corollary 5.2).
$\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation to smCal and $\underline{\textsc{distCal}}$:
$\frac{1}{3}\,\underline{\textsc{distCal}}(\bm{r},\bm{y})$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 3\,\underline{\textsc{distCal}}(\bm{r},\bm{y}),$
$\frac{2}{3}\,\textsc{smCal}(\bm{r},\bm{y})$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 6\,\textsc{smCal}(\bm{r},\bm{y}).$
Based on Theorem 1.10, the quadratic relationship between $\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$ follows from the relatively easy observation that $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are themselves quadratically related.
Similar to $\mathrm{ATB}$, $\ell_{1}\textup{-}\mathrm{ATB}$ can also be easily computed in $O(T\log T)$ time (see Section 4.4), and thus gives a faster algorithm for approximating smCal and $\underline{\textsc{distCal}}$ than direct computation, for which the best-known algorithms take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
Previously, Błasiok et al. (2023) also introduced a binning-based approximation to the distance to calibration, termed the interval calibration error. However, the definition of interval calibration error is more involved than $\ell_{1}\textup{-}\mathrm{ATB}$. It requires optimizing the binning scheme (including the number of bins) to minimize the sum of the average bin width and the binned calibration error. Also, the interval calibration error only gives a quadratic approximation for smCal and $\underline{\textsc{distCal}}$. Our Theorem 1.10 shows, perhaps surprisingly, that using only $2$ bins suffices to give a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$ (see Section 1.4.2 for a technical overview).
1.4 Technical Overview
We give a high-level explanation for the two major technical ingredients we use to establish our results. The first is a general recipe for constructing truthful error metrics leveraging the variance additivity of independent random variables. The second is our analysis establishing the constant-factor approximation between $\ell_{1}\textup{-}\mathrm{ATB}$ and the smooth calibration error.
1.4.1 Truthfulness from Variance Additivity
We discuss the idea behind our construction of a general family of truthful measures, i.e., Unnormalized Binned Squared Errors (UBSEs). As mentioned earlier, ATB is a member of this family, so its truthfulness follows as a consequence.
Our goal is to measure the calibration error of a sequence of predictions $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ on the states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$. Here, each state $y_{t}$ is sampled independently from the Bernoulli distribution with mean $p_{t}\in[0,1]$, where $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ are the true probabilities (denoted by $\bm{y}\sim\bm{p}$).
Our first observation is that, if we divide the predictions $r_{t}$ into bins based on the indices $t\in[T]$ rather than the values $r_{t}\in[0,1]$, then truthfulness can be easily achieved by the $\ell_{2}$ version of binned ECE, $\ell_{2}\textup{-}\textsc{BinECE}$. Concretely, consider a fixed partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ of the index space $[T]$ into bins: $T]=B_{1}\cup\cdots\cup B_{k}$. The $\ell_{2}\textup{-}\textsc{BinECE}$ follows the standard computation of ECE but replacing the $\ell_{1}$ error with squared error:
$\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{r},\bm{y})$ $=\sum_{i\in[k]}\underbrace{\frac{|B_{i}|}{T}}_{\text{weigh by fraction}}\cdot\left(\underbrace{\frac{1}{|B_{i}|}}_{\text{normalize by size}}\underbrace{\sum_{t\in B_{i}}(r_{t}-y_{t})}_{\text{the bias in $B_{i}$}}\right)^{2}$
$=\sum_{i\in[k]}\frac{1}{T|B_{i}|}\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}.$ (2)
Assuming the index partition $\mathcal{B}$ is fixed, the truthfulness of $\ell_{2}\textup{-}\textsc{BinECE}$ comes from the truthfulness of squared error: within each bin $B_{i}$, the expected squared bias over $\bm{y}\sim\bm{p}$
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}\right]$ (3)
is minimized if and only if $\sum_{t\in B_{i}}r_{t}=\sum_{t\in B_{i}}p_{t}$, implying minimized by predicting the truth $r_{t}=p_{t}$.
However, $\ell_{2}\textup{-}\textsc{BinECE}$ is not truthful because it does not bin based on indices $t$, but rather the prediction values $r_{t}$. Specifically, $\ell_{2}\textup{-}\textsc{BinECE}$ combines adjacent predictions into the same bin. It works by first partitioning the prediction space $[0,1]$ into intervals, with one bin corresponding to each interval. Each bin is decided by the predictions $v_{t}$ that fall within the corresponding interval. A strategic predictor will try to reduce the expected error by manipulating the partition $\mathcal{B}$ via making untruthful predictions. For the same Example 1.1, always predicting the same value $r_{1}=\cdots=r_{T}$ puts all the indices in the same bin, resulting in a different index partition $\mathcal{B}^{\prime}$ than predicting truthfully. This different partition can significantly reduce the expected $\ell_{2}\textup{-}\textsc{BinECE}$ due to the bin-size-based normalization $1/|B_{i}|$ in (2).
The example above hints that we can achieve truthfulness if the expected error of truthful predictions is invaiant to the index partition $\mathcal{B}$. To see this, our analysis above shows that for any partition $\mathcal{B}^{\prime}$ that could potentially be manipulated from a strategic report $\bm{r}$, predicting the truth achieves smaller or equal expected error on the same parition $\mathcal{B}^{\prime}$:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{r},\bm{y})].$ (4)
To establish truthfulness, we need to show that the truthful predictions $\bm{p}$ achieve smaller or equal expected error on the un-manipulated partition $\mathcal{B}$ from truthful predictions:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{r},\bm{y})].$ (5)
To bridge the gap between what we have (4) and what we need (5), it suffices if the expected truthful error is invariant of the binned partition for truthful predictions:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{p},\bm{y})]=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{p},\bm{y})]\quad\text{for any partitions $\mathcal{B},\mathcal{B}^{\prime}$}.$ (6)
We obtain Equation 6 with our second key observation: we can modify the $\ell_{2}\textup{-}\textsc{BinECE}$ in (2) to achieve partition-invariance by removing the bin-size-based normalization $1/|B_{i}|$. To see this, let us compute the expected squared bias in each bin $B_{i}$ for truthful predictions ($r_{t}=p_{t}$):
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\left(\sum_{t\in B_{i}}(p_{t}-y_{t})\right)^{2}\right]={\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\sum_{t\in B_{i}}y_{t}\right]=\sum_{t\in B_{i}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[y_{t}\right]=\sum_{t\in B_{i}}p_{t}(1-p_{t}),$ (7)
Here we crucially use the variance additivity of independent random variables: since the $y_{t}$’s are independent, the variance of their sum is equal to the sum of their variances. If we directly add up (7) over the bins $i=1,\ldots,k$ without any bin-size-based normalization, we get
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\left(\sum_{t\in B_{i}}(p_{t}-y_{t})\right)^{2}\right]=\sum_{t\in[T]}p_{t}(1-p_{t}),$
which is indeed invariant to the partition $\mathcal{B}$, satisfying (6). Therefore, our analysis shows that the following unnormalized calibration error is truthful, where now the partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ (including the choice of $k$) can (arbitrarily!) depend on the predictions $r_{1},\ldots,r_{T}$:
$\textsc{Cal}(\bm{r},\bm{y}):=\frac{1}{T^{2}}\sum_{i=1}^{k}\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}.$
Moreover, one can show that any calibrated predictions (not just truthful predictions) achieve the same expected error $\frac{1}{T^{2}}\sum_{t\in[T]}p_{t}(1-p_{t})=O\left(\frac{1}{T}\right)$ (see Lemma 3.3), which vanishes as $T\to\infty$, implying the completeness of the error (Definition 2.8) in addition to truthfulness. Note that the global normalization by $1/T^{2}$ ensures completeness while not affecting the truthfulness property, unlike the per-bin normalization in (2) depending on individual bin sizes $|B_{i}|$.
The final observation that completes our construction of UBSEs is that, by the linearity of expectation, we can even allow the partition $\mathcal{B}$ to be randomized and use the expected error over the random partition $\mathcal{B}$, without breaking the truthfulness analysis above. This allows us to construct a general family of truthful calibration errors, each using a different, possibility randomized, partition $\mathcal{B}$ that can depend on the predictions $r_{1},\ldots,r_{T}$. We term these calibration errors Unnormalized Binned Squared Errors (UBSEs) and present the formal definition and analysis in Section 3.
Example: Quantile $\ell_{2}\textup{-}\textsc{BinECE}$ is truthful.
As an example of UBSE, simple modifications make $\ell_{2}\textup{-}\textsc{BinECE}$ truthful by binning predictions according to quantiles. With $k$ bins, the following UBSE is truthful.
  • Sort the samples by reported predictions with $r_{1}\leq\dots\leq r_{T}$. Break ties uniformly at random.
  • Divide predictions into $k$ bins, with $\frac{T}{k}$ numbers of samples in each bin.
  • Calculate UBSE.
Binning according to quantiles ensures that each bin contains the same number of predictions and thus, the normalization factors based on bin sizes $|B_{i}|$ in (2) no longer break truthfulness.
1.4.2 Two-Bin Approximation of the Smooth Calibration Error
Our UBSE framework is flexible with regard to how the bins should be chosen (including how many bins should be chosen). However, it is not obvious to find an appropriate binning scheme and show that the corresponding UBSE is polynomially related to existing calibration error metrics such as smCal and $\underline{\textsc{distCal}}$.
Our construction of $\mathrm{ATB}$ is quadratically related to smCal and $\underline{\textsc{distCal}}$. As mentioned earlier, we prove this result by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ (Definition 1.9) gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$. Here we explain the intuition behind this analysis.
Our analysis is divided into the following two results, showing the upper and lower bounds on $\ell_{1}\textup{-}\mathrm{ATB}$ separately:
Lemma 5.3: $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 3\,\underline{\textsc{distCal}}(\bm{r},\bm{y})$ (8)
Lemma 5.4: $\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\,\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (9)
The desired constant-factor approximation (Theorem 1.10) then follows from the previous result that smCal and $\underline{\textsc{distCal}}$ are themselves constant-factor approximations of each other (Proposition 2.6) (Błasiok et al., 2023).
While neither inequality is straightforward to prove, the relatively more technically involved and, perhaps, more surprising direction is the latter inequality (9) showing that smCal can be upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ up to a constant factor. Indeed, the intuition behind the previous notion of interval calibration error intCal (Błasiok et al., 2023) is that having too few bins tends to underestimate smCal, and if the calibration error is much smaller than the average bin width, we should increase the number of bins to faithfully capture smCal. [^3] Consequently, the number of bins used to define $\textsc{intCal}(\bm{r},\bm{y})$ depends on both $\bm{r}$ and $\bm{y}$. In UBSE, the binning scheme can only depend on $\bm{r}$ in order for our truthfulness analysis to hold. The reasoning is that having fewer bins makes more predictions fall into the same bin, among which the biases $r_{t}-y_{t}$ in positive and negative directions cancel out, thus more likely to cause underestimation. For example, having only one bin gives the following UBSE:
$\textsc{Cal}(\bm{r},\bm{y})=\left(\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right)^{2},$
which clearly underestimates smCal (it can be zero even when $\bm{r}$ is mis-calibrated, in which case $\textsc{smCal}(\bm{r},\bm{y})$ is always positive). Therefore, based on this previous intuition, it is somewhat surprising that having just two bins suffices to establish (9).
Proving (9) is equivalent to showing that for any $1$-Lipschitz weight function $w:[0,1]\to[-1,1]$,
$\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t})\leq\frac{3}{2}\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (10)
This equivalence follows from the definition of smCal: it is the supremum of the left-hand side over all $1$-Lipschitz $w:[0,1]\to[-1,1]$ (Definition 2.5).
Figure 1: Writing $w$ as a convex combination of threshold functions.
To illustrate our proof idea, let us first assume that the weight function $w$ is not only Lipschitz, but also monotonically increasing and differentiable (represented by the curve in Figure 1). The key observation is that we can write $w$ as a convex combination of threshold functions as follows. Take a random threshold $\lambda$ uniformly distributed from [1,1][-1,1] and consider the threshold function $w_{\lambda}(r):=\mathsf{sign}(w(r)-\lambda)$ (represented by the bold step function in Figure 1). That is, $w_{\lambda}(r)=1$ if $w(r)\geq\lambda$, and $w_{\lambda}(r)=-1$ if $w(r)<\lambda$. The following key identity expresses $w$ as a convex combination of the threshold functions $w_{\lambda}$:
$w(r)=\operatorname{\mathbb{E}}_{\lambda\sim\mathrm{Unif}([-1,1])}[w_{\lambda}(r)]\quad\text{for every $r\in[0,1]$}.$ (11)
Now for a fixed threshold $\lambda\in[-1,1]$, let $q:=w^{(-1)}(\lambda)\in[0,1]$ be the corresponding threshold on the $r$-axis, where $w^{(-1)}$ is the inverse of $w$ (see Figure 1). In the boundary cases when $\lambda>w(1)$, we choose $q=1$, and similarly, when $\lambda<w(0)$ we choose $q=0$. This ensures [^4] One tiny caveat which we ignore here is that when $\lambda>w(1)$ and thus $q=1$, this identity does not hold at one point: $r=1$.
$w_{\lambda}(r)=\mathsf{sign}(r-q)\quad\text{for every $r\in[0,1]$}.$ (12)
Let $Q$ be the distribution of the resulting $q$ from $\lambda\sim\mathrm{Unif}([-1,1])$. By (11) and (12), we can rewrite the left-hand side of (10) as
$\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t})=\operatorname{\mathbb{E}}_{q\sim Q}\left[\frac{1}{T}\sum_{t\in[T]}\mathsf{sign}(r_{t}-q)\cdot(r_{t}-y_{t})\right].$ (13)
For each fixed choice of $q$, it is straightforward to show that the quantity inside the expectation in (13) is upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ at the same fixed bin threshold $q$ (Definition 1.9). However, the random variable $q$ is distributed differently in the two cases. It is drawn from the distribution $Q$ in (13), whereas it is uniformly distributed over $[0,1]$ in the definition of $\ell_{1}\textup{-}\mathrm{ATB}$.
What remains is to relate the two distributions: $Q$ and $\mathrm{Unif}([0,1])$. Recall that $q\sim Q$ is obtained as $q=w^{(-1)}(\lambda)$ for uniformly distributed $\lambda\in[-1,1]$. It follows that the probability density function (PDF) of $q\sim Q$ is exactly the PDF of $\lambda$ (which is $1/2$ everywhere in [1,1][-1,1]) times the derivative $\nabla w(q)$, except at the boundaries $q=0,1$. Since $w$ is $1$-Lipschitz, we have $\nabla w(q)\leq 1$, and thus the PDF of $q\sim Q$ is at most $1/2$ everywhere in the open interval $(0,1)$. This is sufficient to bound the expectation over $q\sim Q$ in (13) by the expectation over $\mathrm{Unif}([0,1])$ in the definition of $\ell_{1}\textup{-}\mathrm{ATB}$ (Definition 4.1). The boundary cases of $q=0,1$ need to be handled separately, but that turns out to be relatively straightforward.
To fully prove (9), we need to remove the monotonicity and differentiability assumptions on $w$, which is achieved by our formal proof in Section 5. Roughly speaking, without monotonicity, the convex combination of the threshold functions that expresses $w$ might have negative coefficients (so it is a linear combination rather than a convex combination), but the absolute values of the coefficients can still be controlled using the Lipschitzness of $w$. The differentiability assumption can be removed by focusing on the finite set $\{r_{1},\ldots,r_{T}\}$ rather than the full domain $[0,1]$ of $w$.
1.5 Truthfulness and Monotonicity
In this subsection, we include an informal discussion about why it may appear challenging to construct a perfectly truthful calibration measure. This may give a partial explanation for why truthful calibration measures have not appeared before our work, despite the variety of calibration measures introduced in the literature.
At first glance, the existence of a perfectly truthful calibration measure may appear counterintuitive, which arises from the (seemingly) conflicting monotonicity of a truthful error metric and a sound and complete calibration measure. Previous theories on truthful error metrics (a.k.a. proper losses) show that truthfulness is closely tied to convexity and monotonicity. In contrast, the set of calibrated predictors is not convex, and a complete and sound calibration measure should not monotonically decrease as we move the predictions closer to the truth.
More concretely, Lambert (2011) characterizes a truthful error metric to be monotonically decreasing when a prediction moves closer to the truth. Formally, a truthful error of a report must be higher than any convex combination of the report with the state, shown in Figure 2. Calibration, however, specifies a conflicting monotonicity where all calibrated predictors are ranked lowest. As an example, suppose the $T$ realized states are $(0,1,0,1,\dots)$ alternating between 0 and $1$, with an empirical frequency of $50\%$. An uninformative predictor who always predicts $50\%$ should achieve a vanishing calibration error as well as a perfect predictor whose outputs deterministically match the states. Now consider the uniform interpolation between $50\%$ and the ground truth, an alternating prediction sequence of $(25\%,75\%,25\%,\dots)$. Intuitively, this interpolation seems to move “closer” to the perfect predictor. By monotonicity, the error of the interpolation should be upper-bounded by the vanishing error of reporting $50\%$. Yet this interpolation is very miscalibrated, and soundness requires it to receive a high calibration error.
increasing lossrealized statereport
Figure 2: The order sensitivity of a truthful error metric. The large circle is an abstraction of the probabilistic space, with a realized state on a corner of the space. The reported prediction lies in the interior of the space. Fixing the realized state, the truthful error, as a function of the prediction, is increasing along the convex combination from the realized state to the reported prediction. For one binary state prediction, fixing the realized state, a truthful error is monotone in the distance between the reported prediction and the state.
Perhaps due to the reasons above, previous (Haghtalab et al., 2024) and concurrent (Qiao and Zhao, 2025) work focuses on achieving the weaker approximate truthfulness in the more challenging online setting. We show that there exists a perfectly truthful calibration error in the offline setting. The reason behind the counterintuitive possibility of perfect truthfulness is that when the states are drawn independently across individuals (i.e., a product distribution), the monotonicity of a truthful calibration measure is not violated, as the interpolation of two non-trivial product distributions is no longer a product distribution. Thus, the intuition of “moving closer” in the example above does not hold rigorously true. Our truthful construction based on variance additivity indeed crucially uses this independence/product distribution structure.
1.6 Related Work
Truthful Calibration Errors.
Previous work (Haghtalab et al., 2024; Qiao and Zhao, 2025) on approximate truthful calibration errors are closest to our paper. They design multiplicatively truthful calibration error in the sequential prediction problem. In the sequential prediction setting, a sequence of $T$ potentially correlated states is drawn from a distribution. At each period, the predictor predicts, and one state is revealed. Our work studies a different batch setting where all $T$ states are independently drawn and revealed simultaneously after all predictions. An error metric is approximately truthful if predicting the true conditional probability of the next state is a constant approximation of the optimal strategy. Haghtalab et al. (2024) shows that subsampled smooth calibration error is multiplicatively truthful for the sequential prediction setting, implying the smooth calibration error is multiplicatively truthful for the batch setting. Qiao and Zhao (2025) shows that, in the sequential setting, there does not exist a perfectly truthful calibration error that upperbounds the worst-case external regret for decision-makers. The impossibility in sequential setting does not apply to our problem. It also remains open whether there exists a perfectly truthful calibration error metric for the sequential setting while satisfying other completeness and soundness properties.
Calibration Measures.
Foster and Vohra (1997) first proposes the Expected Calibration Error (ECE). The binned ECE serves as a widely-used empirical proxy of ECE (Guo et al., 2017; Minderer et al., 2021). Kleinberg et al. (2023) observes that, if predictions are used for downstream decision-making, ECE upperbounds the swap regret of any downstream decision-maker. Following the decision-making purpose of calibration, Hu and Wu (2024) proposes Calibration Decision Loss (CDL), the worst-case swap regret of any normalized downstream decision task, and shows CDL is quantitatively different from ECE. Okoroafor et al. (2025) introduce the notion of proper calibration as a key ingredient for designing improved algorithms for omniprediction (Gopalan et al., 2022, 2023).
Błasiok et al. (2023) introduced the distance to calibration. In their framework, a calibration error is consistent if it is polynomially related to the distance to calibration. They showed that the smooth calibration error (Kakade and Foster, 2008) and the Laplace kernel calibration error (Kumar et al., 2018) are both consistent, and introduced a binning-based consistent calibration error: the interval calibration error.
Proper Scoring Rules (a.k.a. truthful losses).
Initiated by McCarthy (1956); Savage (1971), extensive work focused on the characterization of proper scoring rules, the class of truthful loss functions. Lambert (2011) characterizes elicitable statistics of a distribution, for example, the mean of a distribution, not the variance of a distribution. Winkler et al. (1996) provides proper scoring rules for the confidence interval, and Frongillo and Kash (2014) provides a characterization of proper scoring rules for eliciting linear properties. Li ⓡ al. (2022) gives computational results of proper scoring rules.
1.7 Paper Organization
The rest of the paper is organized as follows. Section 2 establishes the basic setup, including the definitions of existing calibration errors (Section 2.1), completeness and soundness (Section 2.2), the validity of calibration errors via calibration tests (Section 2.3), and the truthfulness of calibration errors (Section 2.4). In Section 3, we introduce the Unnormalized Binned Squared Errors (UBSEs), a general family of truthful binning-based error metrics. In Section 4, we introduce our proposed calibration error, the Averaged Two-Bin calibration error ($\mathrm{ATB}$) as a special case of UBSE and prove its truthfulness, continuity, sample efficiency, and computationally efficiency. In Section 5, we prove the quadratic relationship between $\mathrm{ATB}$ and the existing calibration errors $\textsc{smCal},\underline{\textsc{distCal}}$ by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of smCal and $\underline{\textsc{distCal}}$. In Section 6, we show that $\mathrm{ATB}$ is optimally valid for smCal and $\underline{\textsc{distCal}}$, implying a simple and efficient calibration tester for smCal and $\underline{\textsc{distCal}}$ with improved running time over previous work.
2 Preliminaries
Throughout the paper, we use $D$ to denote a joint distribution of $(x,y)$ pairs, where $x\in X$ represents an individual in a domain $X$, and $y\in\{0,1\}$ is the corresponding state (a.k.a. outcome or label). A predictor $r:X\to[0,1]$ reports a prediction $r(x)\in[0,1]$ for each individual $x\in X$.
2.1 Calibration
We present the formal definitions of a few important calibration error metrics in the literature. We start with the definition of calibration:
Definition 2.1 (Calibration).
A predictor $r:X\to[0,1]$ is calibrated on an underlying distribution $D$ of $(x,y)\in X\times\{0,1\}$ if $\operatorname{\mathbb{E}}_{D}[y|r(x)]=r(x)$ holds almost surely.
An important property of the definition of calibration is that it only depends on the distribution of the prediction-state pair $(r(x),y)\in[0,1]\times\{0,1\}$. That is, we can determine whether a predictor $r$ is calibrated on a distribution $D$ just based on the distribution of $(r(x),y)$, without having to know the full joint distribution of $(x,r(x),y)$. Thus, using a random variable $v$ to represent the prediction value $r(x)$, we can define calibration simply given a distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$:
Definition 2.2 (Calibration of prediction-state distributions).
We say a distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$ is calibrated if $\operatorname{\mathbb{E}}_{J}[y|v]=v$ holds almost surely.
For a distribution $D$ of $(x,y)\in X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we use $J_{D,r}$ to denote the joint distribution of $(r(x),y)$. With that, $r$ is calibrated on $D$ if and only if $J_{D,r}$ is calibrated as in Definition 2.2.
A calibration measure $\textsc{Cal}_{D}(r)\in\mathbb{R}_{\geq 0}$ evaluates the deviation of a predictor $r$ from being perfectly calibrated on a distribution $D$. Naturally, we define a calibration measure $\textsc{Cal}(J)$ first for general prediction-state distributions $J$ of $(v,y)\in[0,1]\times\{0,1\}$, and then define
$\textsc{Cal}_{D}(r):=\textsc{Cal}(J_{D,r}).$
The most well-known calibration measure is the expected calibration error (ECE):
Definition 2.3 (Expected Calibration Error (ECE) (Foster and Vohra, 1997)).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$, and let random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ be the conditional expectation of the state $y$ given the prediction value $v$. The expected calibration error (ECE) is defined as
$v^|.\textsc{ECE}(J):=\operatorname{\mathbb{E}}$|v-$\widehat${v}|.
Correspondingly, for a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, defining $\widehat{r}(x):=\operatorname{\mathbb{E}}_{D}[y|r(x)]$, we have
$\textsc{ECE}_{D}(r):=\textsc{ECE}(J_{D,r})=\operatorname{\mathbb{E}}_{D}|r(x)-\widehat{r}(x)|.$
More generally, for every $\alpha\geq 1$, we define $\ell_{\alpha}$-ECE as follows:
$\ell_{\alpha}\textup{-}\textsc{ECE}(J):=\operatorname{\mathbb{E}}[|v-\widehat{v}|^{\alpha}],\quad\ell_{\alpha}\textup{-}\textsc{ECE}_{D}(r):=\operatorname{\mathbb{E}}_{D}[|r(x)-\widehat{r}(x)|^{\alpha}].$
A downside of the ECE is its discontinuity: slight changes in the predictions $r(x)$ can cause significant changes to the ECE value. This motivated Błasiok et al. (2023) to introduce a continuous calibration error metric, termed the distance to calibration. It measures the earthmover distance from the prediction-state distribution $(v,y)$ to a calibrated distribution $(u,y)$.
Definition 2.4 ((Lower) Distance to Calibration (Błasiok et al., 2023)).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. Consider a joint distribution (i.e. coupling) $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed according to $J$, and the distribution of $(u,y)$ is calibrated as in Definition 2.2. The (lower) distance to calibration is defined as the following infimum over all such couplings $\Pi$:
$\underline{\textsc{distCal}}(J):=\inf_{\Pi}\operatorname{\mathbb{E}}_{\Pi}|u-v|.$
Correspondingly, given a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we define $\underline{\textsc{distCal}}_{D}(r):=\underline{\textsc{distCal}}(J_{D,r})$.
One might imagine a different definition of the distance to calibration as the minimum $L_{1}$ distance $\operatorname{\mathbb{E}}_{D}|r(x)-r^{\prime}(x)|$ from the given predictor $r$ to a calibrated predictor $r^{\prime}$. Indeed, this notion (denoted distCal) is the first definition of the distance to calibration introduced by Błasiok et al. (2023). However, as shown by Błasiok et al. (2023), this definition is different from the $\underline{\textsc{distCal}}$ in Definition 2.4 and has the disadvantage of depending on the full joint distribution of $(x,r(x),y)$, not just the prediction-state distribution of $(r(x),y)$. To address this disadvantage, Błasiok et al. (2023) introduced the $\underline{\textsc{distCal}}$ in Definition 2.4 and termed it the lower distance to calibration. They also showed that the two definitions are quadratically related:
$\frac{1}{16}\,\textsc{distCal}_{D}(r)^{2}\leq\underline{\textsc{distCal}}_{D}(r)\leq\textsc{distCal}_{D}(r).$
We will focus on the lower distance to calibration in Definition 2.4 throughout the paper and will often drop the word “lower” for brevity.
Another important continuous calibration measure is the smooth calibration error introduced by Kakade and Foster (2008) (originally termed weak calibration). As shown by Błasiok et al. (2023), the smooth calibration error smCal gives a constant factor approximation to $\underline{\textsc{distCal}}$ (see Proposition 2.6 below).
Definition 2.5 (Smooth Calibration Error (Kakade and Foster, 2008)).
Let $W_{1}$ be the family of $1$-Lipschitz functions $w:[0,1]\to[-1,1]$. For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$, the smooth calibration error is defined as
$\textsc{smCal}(J):=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}_{J}[(v-y)w(v)].$ (14)
Correspondingly, for a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we have
$\textsc{smCal}_{D}(r):=\textsc{smCal}(J_{D,r})=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}_{J}[(r(x)-y)w(r(x))].$
Without the Lipschitzness constraint on $w$, the smooth calibration error would become the ECE (Definition 2.3), where the supremum in (14) is achieved by
$w(v)=\begin{cases}1,&\text{if }\widehat{v}>v;\\ -1,&\text{otherwise.}\end{cases}$
The following proposition shows that $\underline{\textsc{distCal}}$ and smCal are constant factor approximations of each other:
Proposition 2.6 (Błasiok et al. 2023).
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
$\frac{1}{2}\,\underline{\textsc{distCal}}(J)\leq\textsc{smCal}(J)\leq 2\,\underline{\textsc{distCal}}(J).$
2.2 Completeness and Soundness over Samples
A basic property shared by all the calibration measures in Section 2.1 is that they are all minimized when the predictor is calibrated, with the minimum value being zero:
Claim 2.7.
For $\textsc{Cal}\in\{\textsc{ECE},\underline{\textsc{distCal}},\textsc{smCal}\}$, we have $\textsc{Cal}(J)\geq 0$ for any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$. Moreover,
[…]
The claim above tells us that we can determine whether a predictor $r$ is calibrated on a distribution $D$ by checking whether the calibration error $\textsc{Cal}_{D}(r)$ is zero. In practice, however, we rarely get access to the full distribution $D$ and can only compute the calibration error on an i.i.d. sample drawn from $D$. As we shall see, the property in 2.7 does not necessarily generalize to this sample-access scenario, even in the limit of infinitely large sample sizes. To formalize this intuition, we define completeness and soundness below.
Completeness requires that if a predictor is calibrated, the calibration error should vanish as the sample size increases:
Definition 2.8 (Completeness).
We say a calibration measure Cal is complete if the following holds. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is calibrated (see Definition 2.2). Let $S=\{(v_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $J$ (denoted $S\sim J^{T}$), and let $J_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\lim_{T\to\infty}\operatorname{\mathbb{E}}_{S\sim J^{T}}[\textsc{Cal}(J_{S})]=0.$
Conversely, soundness requires that if a predictor is miscalibrated, the calibration error should not vanish as the sample size increases.
Definition 2.9 (Soundness).
We say a calibration measure Cal is sound if the following holds. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is mis-calibrated. (That is, $J$ does not satisfy Definition 2.2.) Let $S=\{(v_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $J$, and let $J_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\liminf_{T\to\infty}\operatorname{\mathbb{E}}_{S\sim J^{T}}[\textsc{Cal}(J_{S}))]>0.$
It can be inferred from the work of Błasiok et al. (2023) that smCal and $\underline{\textsc{distCal}}$ are both complete and sound. However, while ECE satisfies 2.7, it is not complete. To see this, consider the distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, where $v$ is drawn uniformly from $[1/3,2/3]$, and conditioned on $v$, $y$ is drawn from the Bernoulli distribution with mean $v$. Clearly, $J$ is calibrated and $\textsc{ECE}(J)=0$. However, on a finite sample $S=\{(v_{1},y_{1}),\ldots,(v_{T},y_{T})\}$ drawn i.i.d. from $T$, it holds almost surely that all the $v_{t}$’s are distinct, in which case $\textsc{ECE}(J_{S})\geq 1/3$ (see Example 1.1).
Due to ECE’s lack of completeness, in machine learning practice, the binned ECE (BinECE) is widely adopted as an empirical method for estimating ECE (Guo et al., 2017; Minderer et al., 2021). While we do not need this notion to state our main results, we include its definition here for completeness:
Definition 2.10 (Binned ECE).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. Given $\alpha\geq 1$ and a partition $\mathcal{I}=\{I_{i}\}_{i\in[k]}$ of the prediction space $[0,1]$, the $\ell_{\alpha}$-binned ECE is defined as
[…]
We can estimate $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ using a sample $S=\{(v_{t},y_{t})\}_{t\in T}$ of $T$ i.i.d. points drawn from $J$. Specifically, letting $J_{S}$ denote the (empirical) uniform distribution over $S$, we can use $\ell_{\alpha}\textup{-}\textsc{BinECE}(J_{S})$ as a good estimate for $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ when the sample size $T$ is sufficiently large relative to $k$ (the number of bins). In practice, the number $k$ of bins can be selected according to the sample size $T$, e.g. $k=T^{\frac{1}{3}}$.
Remark 2.11 (Comparison to Haghtalab et al. 2024).
Our definition of completeness follows the same idea as Haghtalab et al. (2024), while our soundness is strictly stronger. There exists an error metric that is not reasonably sound but satisfies the completeness and soundness in Haghtalab et al. (2024).
The soundness definition in Haghtalab et al. (2024) requires that for any empirical distribution $D_{T}$ over $T$ samples,
  • if $r_{t}=1-y_{t}$ for all $t$, then $\lim_{T\to\infty}\textsc{Cal}_{T}(r)=\Omega(1)$;
  • if each state $y\sim\text{Ber}(\alpha)$ is independent and identical Bernoulli, then $\lim_{T\to\infty}\textsc{Cal}_{T}(r)=\Omega(1)$ for $r\neq\beta$ being a non-truthful constant.
We see that the error $\textsc{Cal}=({\mathbf{E}}\left[r\right]-{\mathbf{E}}\left[y\right])^{2}+{\mathbf{E}}\left[\mathbb{I}\left[r\in\{0,1\},y\neq r\right]\right]$ satisfies the requirements above. However, for predictions not in $\{0,1\}$, the error metric only evaluates the unconditional bias in predictions, which is far from a calibration error metric.
2.3 Calibration Test and Validity
Completeness and soundness (Definitions 2.8 and 2.9) ensure that a calibration measure Cal is able to distinguish calibrated predictors from mis-calibrated ones, when the sample size $T$ is large enough. Intuitively, we should expect the distinguishing power to grow as a function of $T$. We characterize this quantitative dependence on $T$ below. We first define calibration tests that aim at accepting calibrated predictors while rejecting mis-calibrated ones, based on a sample of size $T$.
Definition 2.12 (Calibration Test).
Consider the following calibration test using a calibration measure Cal. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$. The test first draws $T$ i.i.d. points from $J$ to form a sample $S=\{(v_{t},y_{t})\}_{t\in[T]}$, and then computes the calibration error $\textsc{Cal}(J_{S})$ on the uniform distribution $J_{S}$ over $S$. The test outputs “accept” if the calibration error does not exceed a threshold $\beta$. That is, the acceptance probability of this test is
$\mathsf{accP}^{\textsc{Cal}}(J;T,\beta):=\Pr\nolimits_{S\sim J^{T}}[\textsc{Cal}(J_{S})\leq\beta].$
We define the validity of a calibration measure Cal given a reference calibration measure Ref that is often chosen to be complete and sound.
Definition 2.13 (Validity).
Let $\{\gamma_{T}\}$ be an infinite sequence of real numbers indexed by $T=1,2,\ldots$. We say a calibration measure Cal is $\{\gamma_{T}\}$-valid w.r.t. a reference calibration measure Ref if there exist thresholds $\beta_{1},\beta_{2},\ldots\in\mathbb{R}$ such that
$\liminf_{T\to\infty}\left(\inf_{J\,:\,\text{calibrated}}\mathsf{accP}^{\textsc{Cal}}(J;T,\beta_{T})-\sup_{J\,:\,\textsc{Ref}(J)\geq\gamma_{T}}\mathsf{accP}^{\textsc{Cal}}(J;T,\beta_{T})\right)>0.$
That is, there is a non-vanishing gap between the acceptance probability when $J$ is calibrated, and the acceptance probability when $J$ is mis-calibrated with error at least $\gamma_{T}$ in the reference measure Ref.
In the definition above, one should typically think of $\gamma_{T}$ as a decreasing function of $T$, which indicates the stronger distinguishing power as $T$ grows. Moreover, the faster $\gamma_{T}$ decreases, the stronger is the distinguishing power of a $\{\gamma_{T}\}$-valid calibration error for large $T$.
2.4 Truthfulness
A truthful error metric incentivizes a strategizing predictor to report the true distribution to minimize expected error on a finite sample.
Definition 2.14 (Truthfulness on Samples).
We say a calibration measure Cal is truthful on samples if the following holds. Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$ and let $p:X\to[0,1]$ be the ground-truth predictor $p(x)=\operatorname{\mathbb{E}}_{D}[y|x]$. Let $S=\{(x_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\quad\text{for any predictor $r:X\to[0,1]$}.$
In this paper, we focus on a stronger notion of truthfulness, where the predictor makes predictions for an arbitrary set of $T$ individuals that may not be i.i.d. sampled from any underlying distribution. We first extend our definition of calibration errors to this setting, where we evaluate the calibration error of a reported prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})$ for the $T$ individuals w.r.t. a ground-truth probability sequence $\bm{p}=(p_{1},\ldots,p_{T})$.
Definition 2.15 (Induced calibration error on prediction sequences).
Given a calibration measure $\textsc{Cal}(J)$ defined on prediction-state distributions $J$ over $0,1]\times\{0,1\}$, we define an induced calibration measure $\textsc{Cal}(\bm{r},\bm{p})$ as follows, where $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ is a sequence of predictions and $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ is a sequence of ground-truth probabilities. Let $J_{\bm{r},\bm{p}}$ be the distribution of $(r_{t},y)\in[0,1]\times\{0,1\}$ where $t$ is drawn uniformly from $[T]$, and $y\in\{0,1\}$ is drawn from the Bernoulli distribution with mean $p_{t}$. We define
$\textsc{Cal}(\bm{r},\bm{p}):=\textsc{Cal}(J_{\bm{r},\bm{p}}).$
For example, according to Definition 2.15, we can explicitly calculate $\textsc{ECE}(\bm{r},\bm{p})$ and $\textsc{smCal}(\bm{r},\bm{p})$ as follows. Recall that for $v\in\{r_{1},\ldots,r_{T}\}$, we define
$\widehat{v}:=\operatorname{\mathbb{E}}_{(v,y)\sim J_{\bm{r},\bm{p}}}[y|v]=\frac{\sum_{t\in[T]}p_{t}\mathbb{I}\left[r_{t}=v\right]}{\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]}.$ (15)
We have
$\textsc{ECE}(\bm{r},\bm{p})=\textsc{ECE}(J_{\bm{r},\bm{p}})$ $={\mathbf{E}}_{(v,y)\sim J_{\bm{r},\bm{p}}}\left[\big{|}v-\widehat{v}\big{|}\right$]
$=\frac{1}{T}\sum_{v}\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]\big{|}v-\widehat{v}\big{|}$ ($v$ ranges over all values that appear at least once in the set $\{r_{1},\ldots r_{t}\}$)
$=\frac{1}{T}\sum_{v}\left|(v-\widehat{v})\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]\right|$
$=\frac{1}{T}\sum_{v}\left|\sum_{t\in[T]}(r_{t}-p_{t})\mathbb{I}\left[r_{t}=v\right]\right|.$ (by (15) and $v\mathbb{I}\left[r_{t}=v\right]=r_{t}\mathbb{I}\left[r_{t}=v\right]$)
Similarly for smCal:
$\textsc{smCal}(\bm{r},\bm{p})$ $=\sup_{w\in W_{1}}\frac{1}{T}\sum_{t=1}^{T}(r_{t}-p_{t})w(r_{t}).$ ($W_{1}$ is the same as in Definition 2.5)
We now define the notion of truthfulness for the calibration errors from Definition 2.15 on length-$T$ sequences. We note that this definition is akin to the definition of properness in the literature of proper scoring rules (McCarthy, 1956; Savage, 1971).
Definition 2.16 (Truthfulness on Sequences).
We say a calibration measure Cal is truthful on sequences if the following holds for any $T\in\mathbb{Z}_{>0}$. Let $\bm{p}:=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be an arbitrary sequence of ground-truth predictions. Let $\bm{y}=(y_{1},\ldots,y_{t})$ denote the randomly realized states, where each $y_{t}\in\{0,1\}$ is drawn independently from the Bernoulli distribution with mean $p_{i}$ (denoted $\bm{y}\sim\bm{p}$). Then
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]\quad\text{for any $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$}.$
Claim 2.17.
Truthfulness on sequences implies truthfulness on samples.
Proof.
As in Definition 2.14, consider a sample $S=\{(x_{1},y_{1}),\ldots,\allowbreak(x_{T},y_{T})\}$ of i.i.d. points from a distribution $D$ over $X\times\{0,1\}$, and let $r:X\to[0,1]$ be a predictor. Define $\bm{r}:=(r(x_{1}),\ldots,r(x_{T}))$ and $\bm{y}:=(y_{1},\ldots,y_{T})$. Now $J_{D_{S},r}$ and $J_{\bm{r},\bm{y}}$ are both equal to the distribution of $(r(x_{t}),y_{t})$ for uniform $t\in[T]$. Therefore,
$\textsc{Cal}_{D_{S}}(r)=\textsc{Cal}(J_{D_{S},r})=\textsc{Cal}(J_{\bm{r},\bm{y}})=\textsc{Cal}(\bm{r},\bm{y}).$ (16)
As in Definition 2.14, define $p(x_{t}):=\operatorname{\mathbb{E}}_{D}[y|x=x_{t}]\in[0,1]$ for $t=1,\ldots,T$. Conditioned on $x_{1},\ldots,x_{T}$, each $y_{t}$ is distributed independently from the Bernoulli distribution with mean $p(x_{t})$. That is, we have $\bm{y}\sim\bm{p}$ as in Definition 2.16, where $\bm{p}:=(p(x_{1}),\ldots,p(x_{T}))$. Therefore, by (16),
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)|x_{1},\ldots,x_{T}]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})],$ (17)
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)|x_{1},\ldots,x_{T}]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})].$ (18)
Assuming truthfulness on sequences, we know that the quantity in (17) is no smaller than the quantity in (18). Taking the expectation over $x_{1},\ldots,x_{T}$ proves the desired truthfulness on samples. ∎
3 Truthfulness of Unnormalized Binned Squared Errors
In this section, we present a general family of truthful error metrics, which we term unnormalized binned squared errors (UBSEs). As it will become clear, the error $\mathrm{ATB}$ is a special case of UBSEs, so its truthfulness is an immediate consequence of the truthfulness of UBSEs.
Definition 3.1 (Unnormalized Binned Squared Errors).
Consider an error metric $\textsc{Cal}(\bm{r},\bm{p})$ taking as input a report vector $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and a ground-truth vector $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$. We say Cal is an unnormalized binned squared error (UBSE) if it can be calculated as follows:
  1. 1.
    Partition the indices $[T]$ into $k$ disjoint bins: $T]=B_{1}\cup\cdots\cup B_{k}$. Importantly, we allow the partition (including the choice of $k$) to be randomized, and we allow it to depend on the report vector $\bm{r}$ (but not on $\bm{p}$).
  2. 2.
    Compute the bias $\Delta_{i}$ in each bin $B_{i}$:
    $\Delta_{i}:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-p_{t}).$ (19)
  3. 3.
    Output the sum of the squared biases: $\textsc{Cal}(\bm{r},\bm{p}):=\operatorname{\mathbb{E}}_{\mathcal{B}}[\sum_{i=1}^{k}\Delta_{i}^{2}]$, where the expectation is over the randomness of the partition $\mathcal{B}=(B_{1},\ldots,B_{k})$.
The above definition is very similar to the definition of binned $\ell_{2}\text{-}\textsc{ECE}$, but there is a crucial difference. When defining binned $\ell_{2}\text{-}\textsc{ECE}$ for a fixed partition $\mathcal{B}=(B_{1},\ldots,B_{k})$, the bias in each bin is first normalized by the bin size $|B_{i}|$:
$\widetilde{\Delta}_{i}=\frac{1}{|B_{i}|}\sum_{t\in B_{i}}(r_{t}-p_{t}),$
and then squared and summed with weights $|B_{i}|/T$:
$\ell_{2}\text{-}\textsc{ECE}(\bm{r},\bm{p})=\sum_{i=1}^{k}\frac{|B_{i}|}{T}\widetilde{\Delta}_{i}^{2}=\sum_{i=1}^{k}\frac{1}{|B_{i}|T}\left(\sum_{t\in B_{i}}(r_{t}-p_{t})\right)^{2}.$
In contrast, Definition 3.1 takes the unweighted sum of the unnormalized squared biases $\Delta_{i}^{2}$:
$\textsc{Cal}(\bm{r},\bm{p})=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\frac{1}{T^{2}}\left(\sum_{t\in B_{i}}(r_{t}-p_{t})\right)^{2}\right].$
This small difference makes UBSEs truthful (whereas the binned $\ell_{2}\text{-}\textsc{ECE}$ is not):
Theorem 3.2.
Any UBSE error metric Cal is truthful on sequences (Definition 2.16).
In fact, we prove the a stronger result in Lemma 3.3, showing that the expected empirical UBSE decomposes into the UBSE on the true probabilities $\bm{p}$ plus a variance term independent of $\bm{r}$.
Lemma 3.3 (Error Decomposition).
Let Cal be an arbitrary UBSE. For any report sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and any ground-truth vector $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]=\textsc{Cal}(\bm{r},\bm{p})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$
Here $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ is drawn such that each $y_{t}$ independently follows the Bernoulli distribution with mean $p_{t}$ (as in Definition 2.16).
We first prove Theorem 3.2 using Lemma 3.3, and then prove Lemma 3.3.
Proof of Theorem 3.2.
For any $\bm{r},\bm{p}\in[0,1]^{T}$, by Lemma 3.3,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})]=\textsc{Cal}(\bm{r},\bm{p})-\textsc{Cal}(\bm{p},\bm{p}).$ (20)
Clearly, we have $\textsc{Cal}(\bm{r},\bm{p})\geq 0$ and $\textsc{Cal}(\bm{p},\bm{p})=0$. Therefore, the quantity in (20) is non-negative, which means that Cal is truthful on sequences. ∎
Proof of Lemma 3.3.
For a partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ of $[T]$ as in Definition 3.1, we define
$\Delta_{i}$ $:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-y_{t}).$
$\widehat{\Delta}_{i}$ $:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-p_{t}).$
We have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]\right]=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]\right],$ (21)
$\textsc{Cal}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\widehat{\Delta}_{i}^{2}\right].$ (22)
[…]
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]=\sum_{i=1}^{k}\widehat{\Delta}_{i}^{2}+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$ (23)
For every $i=1,\ldots,k$, we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}^{2}]=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}]^{2}+{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\Delta_{i}\right],$ (24)
where
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}]$ $=\widehat{\Delta}_{i},$
${\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\Delta_{i}\right]$ $={\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-y_{t})\right]$
$=\frac{1}{T^{2}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\sum_{t\in B_{i}}y_{t}\right]$
$=\frac{1}{T^{2}}\sum_{t\in B_{i}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[y_{t}\right]$ (the $y_{t}$’s are distributed independently)
$=\frac{1}{T^{2}}\sum_{t\in B_{i}}p_{t}(1-p_{t}).$
Plugging these into (24), we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}^{2}]=\widehat{\Delta}_{i}^{2}+\frac{1}{T^{2}}\sum_{t\in B_{i}}p_{t}(1-p_{t}).$
Summing up over $i=1,\ldots,k$ proves (23). ∎
We remark that in addition to being truthful, UBSEs are also complete. This is because, by Lemma 3.3, the expected error of calibrated predictions is
$\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})\leq\frac{1}{4T}=O(1/T),$
which vanishes as $T\to\infty$.
Example 3.4 (Quantile-Binned $\ell_{2}\text{-}\textsc{ECE}$ is truthful).
As a special case of UBSE, the quantile-binned $\ell_{2}\text{-}\textsc{ECE}$ is truthful and complete. Choosing the number of bins properly as a growing function of $T$, it is also a sound calibration error. It is defined as follows:
For any report sequence $\bm{r}=(r_{1},\dots,r_{T})$ and any vector of realized state $\bm{y}=(y_{1},\dots,y_{T})$,
  • sort the predictions in increasing order with $\bm{r}_{1}\leq\dots\leq\bm{r}_{T}$, break ties uniformly at random.
  • Partition predictions into $k=T^{\nicefrac{{1}}{{3}}}$ bins by quantile. Each bin has $\frac{T}{k}$ predictions.
  • Given the partition above, output the Unnormalized Binned Squared Error $\textsc{Cal}(\bm{r},\bm{y})$.
4 Calibration Errors with Two Bins
In this section, we formally define our new calibration measure: the averaged two-bin calibration error (ATB) as well as its $\ell_{1}$ variant ($\ell_{1}\textup{-}\mathrm{ATB}$). We show that these two errors are quadratically related (Lemma 4.2). We also discuss their various properties: truthfulness (for ATB only), continuity, sample complexity for estimation, and computational efficiency. In the next section (Section 5), we establish their quantitative relationships to smCal and $\underline{\textsc{distCal}}$.
Definition 4.1.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we define the averaged two-bin calibration error (ATB) and its $\ell_{1}$ variant as follows:
$\mathrm{ATB}(J)$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\Big{(}\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v<q\right]\Big{]}\Big{)}^{2}+\Big{(}\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v\geq q\right]\Big{]}\Big{)}^{2}\Big{]},$
$\ell_{1}\textup{-}\mathrm{ATB}(J)$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\Big{|}\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v<q\right]\Big{]}\Big{|}+\Big{|}\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v\geq q\right]\Big{]}\Big{|}\Big{]}.$
Correspondingly, for any prediction sequence $\bm{r}\in[0,1]^{T}$ and ground-truth sequence $\bm{p}\in[0,1]^{T}$,
$\mathrm{ATB}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T^{2}}\left(\left(\sum_{t:r_{t}<q}(r_{t}-p_{t})\right)^{2}+\left(\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right)^{2}\right)\right],$ (25)
$\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-p_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right|\right)\right].$
As a simple consequence of Jensen’s inequality, $\ell_{1}\textup{-}\mathrm{ATB}$ is quadratically related to $\mathrm{ATB}$.
Lemma 4.2.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$,
$\frac{1}{2}\,\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}\leq\mathrm{ATB}(J)\leq\ell_{1}\textup{-}\mathrm{ATB}(J).$
Proof.
Fix a threshold $q$, we write $\Delta_{1}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]$ and $\Delta_{2}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v\geq q\right]]$. The right inequality follows from the fact that $\Delta_{1},\Delta_{2}\in[0,1]$.
Using Jensen’s inequality, we get the left inequality:
$\frac{1}{2}\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}=2\left({\mathbf{E}}_{q}\left[\frac{1}{2}\Delta_{1}(q)+\frac{1}{2}\Delta_{2}(q)\right]\right)^{2}\leq 2{\mathbf{E}}_{q}\left[\frac{1}{2}\Delta_{1}(q)^{2}+\frac{1}{2}\Delta_{2}(q)^{2}\right]=\mathrm{ATB}(J).$
4.1 Truthfulness
From its definition (25), $\mathrm{ATB}$ is clearly a special case of UBSE (Definition 3.1), so its truthfulness follows immediately from Theorem 3.2.
Theorem 4.3 (Truthfulness).
The calibration measure $\mathrm{ATB}$ is truthful on sequences (Definition 2.16). Moreover, $\mathrm{ATB}$ inherits the error decomposition:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$
4.2 Continuity
The following theorem establishes the continuity of $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ with a general formalization. Both errors change continuously as the predictions change.
Theorem 4.4 (Continuity).
Let $\Pi$ be a joint distribution of $(v_{1},v_{2},y)\in[0,1]\times[0,1]\times\{0,1\}$. Let $J_{1}$ denote the distribution of $(v_{1},y)$, and let $J_{2}$ denote the distribution of $(v_{2},y)$. We have
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{1})-\ell_{1}\textup{-}\mathrm{ATB}(J_{2})|$ $\leq 3\,\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (26)
$|\mathrm{ATB}(J_{1})-\mathrm{ATB}(J_{2})|$ $\leq 6\,\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (27)
Proof.
By Definition 4.1, we have
$\ell_{1}\textup{-}\mathrm{ATB}(J_{1})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{1}-y\big{)}\mathbb{I}\left[v_{1}<q\right]\Big{]}\Big{|}+\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{1}-y\big{)}\mathbb{I}\left[v_{1}\geq q\right]\Big{]}\Big{|}\Big{]},$
$\ell_{1}\textup{-}\mathrm{ATB}(J_{2})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{2}-y\big{)}\mathbb{I}\left[v_{2}<q\right]\Big{]}\Big{|}+\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{2}-y\big{)}\mathbb{I}\left[v_{2}\geq q\right]\Big{]}\Big{|}\Big{]}.$
We define an intermediate quantity
$\kappa:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{2}-y\big{)}\mathbb{I}\left[v_{1}<q\right]\Big{]}\Big{|}+\Big{|}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{(}v_{2}-y\big{)}\mathbb{I}\left[v_{1}\geq q\right]\Big{]}\Big{|}\Big{]}.$
By the triangle inequality,
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{1})-\kappa|$ $\leq\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{|}v_{1}-v_{2}\big{|}\mathbb{I}\left[v_{1}<q\right]\Big{]}+\operatorname{\mathbb{E}}_{\Pi}\Big{[}\big{|}v_{1}-v_{2}\big{|}\mathbb{I}\left[v_{1}\geq q\right]\Big{]}\Big{]}$
$=\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (28)
Similarly, noting that $|v_{2}-y|\leq 1$, we have
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{2})-\kappa|$ $\leq\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{[}2\,\operatorname{\mathbb{E}}_{\Pi}\Big{|}\mathbb{I}\left[v_{1}<q\right]-\mathbb{I}\left[v_{2}<q\right]\Big{|}\Big{]}$
$=2\operatorname{\mathbb{E}}_{\Pi}\Big{[}\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big{|}\mathbb{I}\left[v_{1}<q\right]-\mathbb{I}\left[v_{2}<q\right]\Big{|}\Big{]}$
$=2\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (29)
Summing up (28) and (29) proves (26). A similar strategy proves (27), using one extra observation: the function $u^{2}$ is $2$-Lipshitz for $u\in[-1,1]$. We omit the details. ∎
4.3 Sample Complexity
Both $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ can be estimated within error $\varepsilon$ using $O(1/\varepsilon^{2})$ i.i.d. examples:
Theorem 4.5 (Sample complexity).
Let $J$ be any distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, and let $S$ be a sample of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. For $\varepsilon,\delta\in(0,1/3)$, assume $T>C\varepsilon^{-2}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$. With probability at least $1-\delta$ (over the randomness in the sample $S$),
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{S})-\ell_{1}\textup{-}\mathrm{ATB}(J))|\leq\varepsilon,$
$|\mathrm{ATB}(J_{S})-\mathrm{ATB}(J)|\leq\varepsilon.$
Proof.
It suffices to show that with probability at least $1-\delta$, for every $q\in[0,1]$,
$\Big{|}\operatorname{\mathbb{E}}_{J_{S}}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v<q\right]\Big{]}-\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v<q\right]\Big{]}\Big{|}$ $\leq\varepsilon/4,\quad\text{and}$
$\Big{|}\operatorname{\mathbb{E}}_{J_{S}}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v\geq q\right]\Big{]}-\operatorname{\mathbb{E}}_{J}\Big{[}\big{(}v-y\big{)}\mathbb{I}\left[v\geq q\right]\Big{]}\Big{|}$ $\leq\varepsilon/4.$
By Proposition B.2, it suffices to prove the following Rademacher complexity bounds for the function families $F=\{f_{q}\}_{q\in[0,1]}$ and $G=\{g_{q}\}_{q\in[0,1]}$ where $f_{q}(v,y)=(v-y)\mathbb{I}\left[v<q\right]$ and $g_{q}(v,y)=(v-y)\mathbb{I}\left[v\geq q\right]$: for every $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$,
$\mathcal{R}(F;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))$ $\leq O\left(\sqrt{\frac{1}{T}}\right),\quad\text{and}$ (30)
$\quad\mathcal{R}(G;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))$ $\leq O\left(\sqrt{\frac{1}{T}}\right).$ (31)
Now consider the family $H=\{h_{q}\}_{q\in[0,1]}$ where $h_{q}(v,y)=\mathbb{I}\left[v<q\right]$. Clearly, $H$ has VC dimension at most $1$. By Proposition B.5, we have
$\mathcal{R}(H;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))\leq O\left(\sqrt{\frac{1}{T}}\right).$ (32)
Observe that $f_{q}(v_{i},y_{i})=\eta_{i}(h_{q}(v_{i},y_{i}))$ for every $i=1,\ldots,T$ and $q\in[0,1]$, where $\eta_{i}$ is the uni-variate function $\eta_{i}(u)=(v_{i}-y_{i})u$ for $u\in\mathbb{R}$. Since $|v_{i}-y_{i}|\leq 1$, the function $\eta_{i}$ is $1$-Lipschitz. Therefore, by Proposition B.3, inequality (30) follows from (32). Inequality (31) can be proved similarly. ∎
We remark that an analogous sample complexity bound for smCal has been shown by Błasiok et al. (2023) using a similar analysis:
Proposition 4.6 (Błasiok et al. (2023)).
Let $J$ be any distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, and let $S$ be a sample of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. For $\varepsilon,\delta\in(0,1/3)$, assume $T>C\varepsilon^{-2}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$. With probability at least $1-\delta$ (over the randomness in the sample $S$),
$|\textsc{smCal}(J_{S})-\textsc{smCal}(J)|\leq\varepsilon.$
4.4 Computational Efficiency
Given $\bm{r},\bm{p}\in[0,1]^{T}$, we can compute $\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time. Moreover, the algorithm is extremely easy to describe and implement. Define
$\Delta_{1}(q)=\frac{1}{T}\left|\sum_{t:r_{t}<q}(r_{t}-p_{t})\right|\quad\text{and}\quad\Delta_{2}(q)=\frac{1}{T}\left|\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right|.$
The following algorithm computes $\mathrm{ATB}$:
  • $O(T\log T)$ time: sort predictions in increasing order such that $r_{1}\leq r_{2}\leq\dots\leq r_{T}$.
  • $O(T)$ time: calculate $\Delta_{1}(q)$ for each $q$ in the space of reported predictions by scanning predictions in increasing order.
    Similarly, calculate $\Delta_{2}(q)$ by scanning predictions in decreasing order.
  • $O(T)$ time: Write $r_{0}=0$ and $r_{T+1}=1$. Calculate the expectation over threshold $q$: for each $i=0\dots T$, sum $\Delta_{1}(r_{i})^{2}+\Delta_{2}(r_{i})^{2}$ with weight $|r_{i+1}-r_{i}|$.
A similar algorithm computes $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time. We note that currently known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated, with the best known running time being $O(T\log^{2}T)$ and $O(T^{2}\log T)$, respectively (Hu et al., 2024).
5 Approximating the Distance to Calibration Using Two Bins
In this section, we show that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of both smCal and $\underline{\textsc{distCal}}$ (recall Proposition 2.6 that smCal and $\underline{\textsc{distCal}}$ are constant-factor approximations to each other):
Theorem 5.1.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we have
$\frac{2}{3}\,\textsc{smCal}(J)\leq\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J).$
Combining Theorem 5.1 with Proposition 2.6 and Lemma 4.2, we have the following corollary about the relationship between $\mathrm{ATB},\ell_{1}\textup{-}\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$:
Corollary 5.2.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we have
$\frac{1}{3}\,\underline{\textsc{distCal}}(J)\leq\frac{2}{3}\,\textsc{smCal}(J)$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J)\leq 6\,\textsc{smCal}(J),$
$\frac{1}{18}\,\underline{\textsc{distCal}}(J)^{2}\leq\frac{2}{9}\,\textsc{smCal}(J)^{2}$ $\leq\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J)\leq 6\,\textsc{smCal}(J).$
We prove the two inequalities in Theorem 5.1 in two separate lemmas below. We start with the easier one showing the upper bound on $\ell_{1}\textup{-}\mathrm{ATB}$:
Lemma 5.3.
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
$\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J).$
Proof.
Let $\Pi$ be an arbitrary distribution of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where the distribution of $(v,y)$ is $J$, and the distribution of $(u,y)$ (denoted by $\widehat{J}$) is calibrated. Since $\widehat{J}$ is calibrated, we have
$\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})=0.$
By Theorem 4.4,
$\ell_{1}\textup{-}\mathrm{ATB}(J)=\ell_{1}\textup{-}\mathrm{ATB}(J)-\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})\leq 3\operatorname{\mathbb{E}}_{\Pi}|u-v|.$
The lemma is proved by taking the infimum over $\Pi$. ∎
Now we prove the other inequality in Theorem 5.1 showing the lower bound on $\ell_{1}\textup{-}\mathrm{ATB}$. It turns out to be convenient to first focus on the setting with $T$ fixed individuals:
Lemma 5.4.
For any prediction sequence $\bm{r}\in[0,1]^{T}$ and any state sequence $\bm{y}\in\{0,1\}^{T}$, we have
$\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$
Proof of Lemma 5.4.
It suffices to prove that for any $1$-Lipschitz function $w:[0,1]\to[-1,1]$,
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})\leq\frac{3}{2}\cdot\mathrm{ATB}(\bm{r},\bm{y}).$ (33)
Assume without loss of generality that the predictions are sorted: $r_{1}\leq\cdots\leq r_{T}$. Define $w(r_{0})=0,w(r_{T+1})=0$. For $t=0,\ldots,T$, define $\Delta_{t}:=w(r_{t+1})-w(r_{t})$. We have
$w(r_{t})=\frac{1}{2}((w(r_{t})-w(r_{0}))-(w(r_{T+1})-w(r_{t})))$ $=\frac{1}{2}\left(\sum_{s<t}\Delta_{s}-\sum_{s\geq t}\Delta_{s}\right)$
$=\frac{1}{2}\sum_{s=0}^{T}\Delta_{s}\mathsf{sign}(t-s),$
where $\mathsf{sign}(u)=1$ if $u>0$, and $\mathsf{sign}(u)=-1$ if $u\leq 0$. Therefore,
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})=\frac{1}{2T}\sum_{s=0}^{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s).$ (34)
For $s=1,\ldots,T-1$, by the Lipschitzness of $w$, we have $|\Delta_{s}|\leq r_{s+1}-r_{s}$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|$ $\leq(r_{s+1}-r_{s})\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\mathsf{sign}(t-s)\right|$
$\leq(r_{s+1}-r_{s})\cdot\frac{1}{T}\left(\left|\sum_{t\leq s}(r_{t}-y_{t})\right|+\left|\sum_{t>s}(r_{t}-y_{t})\right|\right)$
$=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\mathbb{I}_{q\in[r_{s},r_{s+1}]}\cdot\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-y_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right|\right)\right].$
Summing up over $s=1,\ldots T-1$, we have
$\sum_{s=1}^{T-1}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq\mathrm{ATB}(\bm{r},\bm{y}).$ (35)
Moreover, since $w_{1},w_{T}\in[-1,1]$, we have $|\Delta_{0}|,|\Delta_{T}|\leq 1$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{0}\mathsf{sign}(t-0)\right|$ $=|\Delta_{0}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\mathrm{ATB}(\bm{r},\bm{y}),$ (36)
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{T}\mathsf{sign}(t-T)\right|$ $=|\Delta_{T}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\mathrm{ATB}(\bm{r},\bm{y}).$ (37)
Adding up the three inequalities (35) (36) (37) above, we get
$\sum_{s=0}^{T}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq 3\mathrm{ATB}(\bm{r},\bm{y}).$
Combining this with (34) using the triangle inequality, we get (33), as desired. ∎
Proof of Theorem 5.1.
The upper bound on $\ell_{1}\textup{-}\mathrm{ATB}$ has been proved in Lemma 5.3. It remains to establish the lower bound on $\ell_{1}\textup{-}\mathrm{ATB}$:
$\textsc{smCal}(J)\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(J).$ (38)
Consider a sample $S$ of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. Defining $\bm{r}:=(v_{1},\ldots,v_{T})$ and $\bm{y}:=(y_{1},\ldots,y_{T})$, we have
$\textsc{smCal}(J_{S})$ $=\textsc{smCal}(J_{\bm{r},\bm{y}})=\textsc{smCal}(\bm{r},\bm{y}),$
$\ell_{1}\textup{-}\mathrm{ATB}(J_{S})$ $=\ell_{1}\textup{-}\mathrm{ATB}(J_{\bm{r},\bm{y}})=\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}),$
where we use the $J_{\bm{r},\bm{y}}$ notation from Definition 2.15. By Lemma 5.4,
$\textsc{smCal}(J_{S})=\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})=\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(J_{S}).$ (39)
Taking $T\to\infty$, by Theorem 4.5 and Proposition 4.6, we know that $\textsc{smCal}(J_{S})$ converges in probability to $\textsc{smCal}(J)$, and that $\ell_{1}\textup{-}\mathrm{ATB}(J_{S})$ converges in probability to $\ell_{1}\textup{-}\mathrm{ATB}(J)$. Therefore, our goal (38) follows from (39). ∎
6 Extremely Efficient Optimal Calibration Tester
In this section, we show that our $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are both optimally valid for smCal and $\underline{\textsc{distCal}}$, which implies a $O(T\log T)$ time calibration test for smCal and $\underline{\textsc{distCal}}$. It is fairly straightforward to show that $\ell_{1}\textup{-}\mathrm{ATB}$ is $O(1/\sqrt{T})$ valid using its constant approximation to smCal (Theorem 5.1) and its sample complexity bound (Theorem 4.5). In Theorem 6.1 below, we show that $\mathrm{ATB}$ is $O(1/\sqrt{T})$ valid as well, and that this is optimal up to constant.
Theorem 6.1.
The calibration measure $\mathrm{ATB}$ is $O(\frac{1}{\sqrt{T}})$-valid w.r.t. the reference calibration error $\underline{\textsc{distCal}}$. That is, $\mathrm{ATB}$ is $\{\gamma_{T}\}$-valid for some sequence $\gamma_{1},\gamma_{2},\ldots$ with $\gamma_{T}=O(1/\sqrt{T})$. Moreover, this is optimal up to constant factors: if there exists a $\{\gamma_{T}\}$-valid calibration error w.r.t. $\underline{\textsc{distCal}}$, then $\gamma_{T}=\Omega(1/\sqrt{T})$.
Theorem 6.1 is an immediate corollary of the following Lemmas 6.2, 6.3, and 6.4.
Lemma 6.2.
Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is calibrated. For any $T\in\mathbb{Z}_{>0}$, consider a sample $S$ of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$ from $J$, and let $J_{S}$ be the uniform distribution over $S$. We have
$\Pr_{S\sim J^{T}}[\mathrm{ATB}(J_{S})\leq 1/T]\geq 3/4.$
Proof.
Define $\bm{r}=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$. It is clear that the distribution $J_{\bm{r},\bm{y}}$ (see Definition 2.15) is equal to the distribution $J_{S}$. Therefore,
$\mathrm{ATB}(J_{S})=\mathrm{ATB}(\bm{r},\bm{y}).$
Since $J$ is calibrated, we have $\operatorname{\mathbb{E}}_{J}[y|v=v_{t}]=v_{t}$ for every $t=1,\ldots,T$. Conditioned on $\bm{r}=(v_{1},\ldots,v_{T})$, each $y_{t}$ is independently distributed as the Bernoulli distribution with mean $v_{t}$. Thus, we have $\bm{y}\sim\bm{r}$ as in Definition 2.16. Therefore,
$\Pr_{S}[\mathrm{ATB}(J_{S})\leq 1/T\ |\ v_{1},\ldots,v_{T}]=\Pr_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})\leq 1/T].$ (40)
By Lemma 3.3,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{r})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})=\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})\leq\frac{1}{4T}.$
By Markov’s inequality,
$\Pr_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})\leq 1/T]\geq 3/4.$
Plugging this into (40) and taking the expectation over $v_{1},\ldots,v_{T}$ completes the proof. ∎
Lemma 6.3.
There exists an absolute constant $C>0$ such that the following holds. For any $T\in\mathbb{Z}_{>0}$ and any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$ with $\underline{\textsc{distCal}}(J)\geq C/\sqrt{T}$, let $S$ be a sample of $T$ i.i.d. points from $J$. Then
$\Pr_{S\sim J^{T}}[\mathrm{ATB}(J_{S})\leq 1/T]\leq 1/4.$
Proof.
By Theorem 4.5, there exists an absolute constant $C^{\prime}>0$ such that with probability at least $3/4$ over $S\sim J^{T}$,
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{S})-\ell_{1}\textup{-}\mathrm{ATB}(J)|\leq C^{\prime}/\sqrt{T}.$ (41)
It remains to show that whenever (41) holds, we have
$\mathrm{ATB}(J_{S})>1/T.$
By Corollary 5.2 and our assumption that $\underline{\textsc{distCal}}(J)>C/\sqrt{T}$, we have $\ell_{1}\textup{-}\mathrm{ATB}(J)\geq(C/3)/\sqrt{T}$. Therefore, whenever (41) holds, we have
$\ell_{1}\textup{-}\mathrm{ATB}(J_{S})\geq(C/3-C^{\prime})/\sqrt{T}.$
Assuming $C/3-C^{\prime}>0$ which is guaranteed by a sufficiently large $C$, by Lemma 4.2, we have
$\mathrm{ATB}(J_{S})\geq(1/2)(C/3-C^{\prime})^{2}/T.$
The proof is completed by choosing $C$ large enough so that $(1/2)(C/3-C^{\prime})^{2}>1$. ∎
Lemma 6.4.
Let $\{\gamma_{T}\}_{T=1,2,\ldots}$ be a sequence of nonnegative real numbers such that there exists a $\{\gamma_{T}\}$-valid calibration error Cal w.r.t. $\underline{\textsc{distCal}}$. Then $\gamma_{T}=\Omega(1/\sqrt{T})$.
Proof.
Let us focus on the choices of $T$ such that $\gamma_{T}<1/2$. We define $J_{1}$ to be the uniform distribution over $\{(1/2,0),(1/2,1)\}\subseteq[0,1]\times\{0,1\}$. We define $J_{2}$ to be the distribution with probability mass $1/2-\gamma_{T}$ on $(1/2,0)$, and the remaining probability mass $1/2+\gamma_{T}$ on $(1/2,1)$.
Clearly, $J_{1}$ is calibrated. We claim that $\underline{\textsc{distCal}}(J_{2})\geq\gamma_{T}$. Indeed, consider any coupling distribution $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed as $J_{2}$, and the distribution of $(u,y)$ is calibrated. By calibration, $\operatorname{\mathbb{E}}[u]=\operatorname{\mathbb{E}}[y]=1/2+\gamma_{T}$. Therefore, $\operatorname{\mathbb{E}}|u-v|\geq\operatorname{\mathbb{E}}[u]-\operatorname{\mathbb{E}}[v]=\gamma_{T}$, implying that $\underline{\textsc{distCal}}(J_{2})\geq\gamma_{T}$.
Let $\beta_{T}\in\mathbb{R}$ be the threshold satisfying the requirement of validity (Definition 2.13). Define
$\delta_{T}:=\mathsf{accP}^{\textsc{Cal}}(J_{1};T,\beta_{T})-\mathsf{accP}^{\textsc{Cal}}(J_{2};T,\beta_{T}).$
Note that the two acceptance probabilities above are w.r.t. the randomness in the samples $S_{1}\sim J_{1}^{T}$ and $S_{2}\sim J_{2}^{T}$, respectively, where $J_{1}^{T}$ (resp. $J_{2}^{T}$) is the joint distribution of $T$ i.i.d. points from $J_{1}$ (resp. $J_{2}$). A standard argument (e.g. via Pinsker’s inequality) shows that the total variation distance between $J_{1}^{T}$ and $J_{2}^{T}$ is $O(\gamma_{T}\sqrt{T})$. Therefore,
$\delta_{T}\leq O(\gamma_{T}\sqrt{T}).$
Validity requires $\liminf_{T\to\infty}\delta_{T}>0$. Therefore,
$\liminf_{T\to\infty}\gamma_{T}\sqrt{T}>0.$
This implies $\gamma_{T}=\Omega(1/\sqrt{T})$. ∎
Appendix A Non-Truthfulness of Known Calibration Measures
In this section, we prove Theorem A.2 showing that condition (1) holds for a broad family of calibration measures: $\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal}$ and $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$ (Definition A.1), where $\alpha\geq 1$ is arbitrary. Condition (1) states that, for any realization of the states, reporting the average over predictions is weakly better than reporting truthfully for known calibration measures. By our discussion in Section 1.1 after (1), this proves that these calibration measures are not truthful.
Definition A.1 ($\ell_{\alpha}$-Distance to Calibration).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. We define its $\ell_{\alpha}$-distance to calibration (denoted by $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}(J)$) similarly to the definition of $\underline{\textsc{distCal}}(J)$ in Definition 2.4. The only difference is that we change the $\ell_{1}$ distance |uv||u-v| to $|u-v|^{\alpha}$:
$\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}(J):=\inf_{\Pi}\operatorname{\mathbb{E}}_{\Pi}[|u-v|^{\alpha}],$
where the infimum is over joint distributions $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed according to $J$, and the distribution of $(u,y)$ is calibrated.
Theorem A.2.
Let Cal be a calibration measure from $\{\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal},\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}\}$, where $\alpha\geq 1$ is arbitrary. For every $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and every $\bm{y}\in\{0,1\}^{T}$, it holds that
$\textsc{Cal}(\bar{\bm{r}},\bm{y})\leq\textsc{Cal}(\bm{r},\bm{y}),$
where $\bar{\bm{r}}:=(\bar{r},\ldots,\bar{r})\in[0,1]^{T}$ for $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$.
Theorem A.2 is a corollary of the following theorem.
Theorem A.3.
Let Cal be a calibration measure from $\{\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal},\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}\}$, where $\alpha\geq 1$ is arbitrary. Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$. We have
$\textsc{Cal}(\bar{J})\leq\textsc{Cal}(J).$ (42)
Here $\bar{J}$ is the distribution of $(\bar{v},y)\in[0,1]\times\{0,1\}$, where we draw $(v,y)\sim J$ and replace $v$ with the deterministic quantity $\bar{v}:=\operatorname{\mathbb{E}}_{J}[v]$.
Proof.
We prove the theorem separately for each choice of Cal. Similarly to the definition of $\bar{v}$, we define $\bar{y}:=\operatorname{\mathbb{E}}_{J}[y]$.
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\textsc{ECE}$, defining $\widehat{v}:=\operatorname{\mathbb{E}}[y|v]$, by Jensen’s Inequality we have
$\textsc{ECE}(J)=\operatorname{\mathbb{E}}[|v-\widehat{v}|^{\alpha}]\geq|\operatorname{\mathbb{E}}[v-\widehat{v}]|^{\alpha}=|\bar{v}-\bar{y}|^{\alpha},$ (43)
where we used the fact that $\operatorname{\mathbb{E}}[\widehat{v}]=\operatorname{\mathbb{E}}[y]=\bar{y}$. Also,
[…]
where we used the fact that $\bar{v}$ is a deterministic quantity, so $\operatorname{\mathbb{E}}[y|\bar{v}]=\operatorname{\mathbb{E}}[y]=\bar{y}$. Combining (43) and (44) proves (42).
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\textsc{BinECE}$, we can prove (42) as follows. Let $\mathcal{I}=\{I_{i}\}_{i\in[k]}$ be the partition of the prediction space $[0,1]$ in the definition of $\ell_{\alpha}\textup{-}\textsc{BinECE}$ (Definition 2.10). By Jensen’s Inequality,
[…]
$=|\operatorname{\mathbb{E}}_{J}[v-y]|^{\alpha}$
$=|\bar{v}-\bar{y}|^{\alpha}.$ (45)
Also, since $\bar{v}$ is deterministic, we have
[…]
Combining (45) and (46) proves (42).
When $\textsc{Cal}=\textsc{smCal}$, (42) follows from the following calculation:
$\textsc{smCal}(\bar{J})$ $=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}[(\bar{v}-y)w(\bar{v})]=\sup_{w\in W_{1}}(\bar{v}-\bar{y})w(\bar{v})=|\bar{v}-\bar{y}|,$
$\textsc{smCal}(J)$ $=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}[(v-y)w(v)]\geq\sup_{\sigma\in\{\pm 1\}}\operatorname{\mathbb{E}}[(v-y)\sigma]=\sup_{\sigma\in\{\pm 1\}}(\bar{v}-\bar{y})\sigma=|\bar{v}-\bar{y}|.$
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$, (42) holds because of the following argument. Consider any joint distribution $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where the marginal distribution of $(v,y)$ is $J$, and the marginal distribution of $(u,y)$ is calibrated. By Jensen’s Inequality,
$\operatorname{\mathbb{E}}_{\Pi}[|u-v|^{\alpha}]\geq|\operatorname{\mathbb{E}}_{\Pi}[u-v]|^{\alpha}=|\operatorname{\mathbb{E}}_{\Pi}[u]-\operatorname{\mathbb{E}}_{\Pi}[v]|^{\alpha}=|\operatorname{\mathbb{E}}_{\Pi}[y]-\operatorname{\mathbb{E}}_{\Pi}[v]]|^{\alpha}=|\bar{y}-\bar{v}|^{\alpha}.$
We used the fact that $\operatorname{\mathbb{E}}_{\Pi}[u]=\operatorname{\mathbb{E}}_{\Pi}[y]$, which holds because the distribution of $(u,y)$ is calibrated. Taking the infimum over $\Pi$, we get
$\underline{\textsc{distCal}}(J)\geq|\bar{y}-\bar{v}|^{\alpha}.$ (47)
Moreover, since always predicting $\bar{y}$ yields a calibrated predictor, we can choose $\Pi$ to be the joint distribution of $(\bar{y},\bar{v},y)$ and get
$\underline{\textsc{distCal}}(\bar{J})\leq\operatorname{\mathbb{E}}_{\Pi}[|\bar{y}-\bar{v}|^{\alpha}]=|\bar{v}-\bar{y}|^{\alpha}.$ (48)
Combining (47) and (48) proves (42). ∎
Table 1 provides an example illustrating the non-truthfulness of known calibration measures and the truthfulness of our $\mathrm{ATB}$. The table compares two strategies: predicting the average over ground truths and predicting truthfully.
Prob. States smCal $\underline{\textsc{distCal}}$ $\ell_{2}\text{-}\underline{\textsc{distCal}}$ $\mathrm{ATB}$ (ours)
avg truth avg truth avg truth avg truth
$\frac{3}{16}$ $(0,0)$ 0.5 0.5 0.5 0.5 0.25 0.3125 0.25 0.203125
$\frac{3}{16}$ $(1,1)$ 0.5 0.5 0.5 0.5 0.25 0.3125 0.25 0.203125
$\frac{9}{16}$ $(0,1)$ 0 $0.0625$ 0 >0>0 0 >0>0 0 0.015625
$\frac{1}{16}$ $(1,0)$ 0 $0.1875$ 0 >0>0 0 >0>0 0 0.140625
Expected Error 0.1875 0.234375 0.1875 >0.1875>0.1875 0.09375 >0.11>0.11 0.09375 0.09375
Table 1: The calibration errors of predictors with two samples. The ground truth distribution of the two samples are $25\%$ and $75\%$, respectively. In the table, avg stands for the uninformative predictor that always outputs $50\%$ and truth stands for the truthful predictor that outputs $25\%$ and $75\%$. We calculate the error of the predictors given each realization of the state and the total expected error. For non-truthful error metrics, the expected error of a truthful predictor is strictly higher than the expected error of an uninformative predictor. For $\mathrm{ATB}$, the expected errors are the same.
Appendix B Standard Uniform Convergence Bounds
We include some standard notions and results on concentration inequalities and sample complexity bounds for uniform convergence. They are used when we prove the sample complexity bounds for estimating $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ in Theorem 4.5.
We start with the definition of the Rademacher complexity.
Definition B.1 (Rademacher complexity).
Let $F$ be a family of real-valued functions $f:Z\to\mathbb{R}$ on some domain $Z$. Given $z_{1},\ldots,z_{n}\in Z$, we define the Rademacher complexity as follows:
$\mathcal{R}(F;z_{1,\ldots,n}):=\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right],$
where the expectation is over $s_{1},\ldots,s_{n}$ drawn uniformly at random from $\{-1,1\}^{n}$.
The following theorem is a standard application of the Rademacher complexity for proving uniform convergence bounds.
Proposition B.2 (Uniform convergence from Rademacher complexity).
Let $F$ be a family of functions $f:Z\to[a,b]$ on some domain $Z$ and with range bounded in $[a,b]$. Let $\Gamma$ be an arbitrary distribution over $Z$. Then for $n$ i.i.d. examples $z_{1},\ldots,z_{n}$ from $\Gamma$,
$\operatorname{\mathbb{E}}_{z_{1,\ldots,n}}\left[\sup_{f\in F}\left|\frac{1}{n}\sum_{i=1}^{n}f(z_{i})-\operatorname{\mathbb{E}}_{z\sim\Gamma}[f(z)]\right|\right]\leq 2\operatorname{\mathbb{E}}_{z_{1,\ldots,n}}[\mathcal{R}(F;z_{1,\ldots,n})].$
Moreover, for any $\delta\in(0,\frac{1}{3})$ and $n\in\mathbb{N}$, with probability at least $1-\delta$ over the random draw of $n$ i.i.d. examples $z_{1},\ldots,z_{n}$ from $\Gamma$, it holds that
$\sup_{f\in F}\left|\frac{1}{n}\sum_{i=1}^{n}f(z_{i})-\operatorname{\mathbb{E}}_{z\sim\Gamma}[f(z)]\right|\leq 2\mathcal{R}(F;z_{1,\ldots,n})+O\left((b-a)\sqrt{\frac{\log(1/\delta)}{n}}\right).$
Proposition B.3 (Rademacher Complexity after Lipschitz Postprocessing).
Let $F$ be a family of functions $f:Z\to\mathbb{R}$. For $i=1,\ldots,n$, let $z_{i}\in Z$ be an element of the domain $Z$ and let $\eta_{i}:\mathbb{R}\to\mathbb{R}$ be any $1$-Lipschitz function. It holds that
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}\eta_{i}(f(z_{i}))\right]\leq\mathcal{R}(F;z_{1,\ldots,n})=\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right].$
Proof.
By induction, it suffices to consider the case where all the $\eta_{i}$’s are the identity function except $\eta_{1}$.
Now we have
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}\eta_{i}(f(z_{i}))\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f\in F}\left(s_{1}\eta_{1}(f(z_{1}))+\sum_{i=2}^{n}s_{i}f(z_{i})\right)+\sup_{f\in F}\left(-s_{1}\eta_{1}(f(z_{1}))+\sum_{i=2}^{n}s_{i}f(z_{i})\right)\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(|\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))|+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right].$ (49)
Similarly,
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(|f_{+}(z_{1})-f_{-}(z_{1})|+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right].$ (50)
By the 11-Lipschitz property of $\eta_{1}$, we have
$|\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))|\leq|f_{+}(z_{1})-f_{-}(z_{1})|.$
This implies that (49) is a lower bound of (50), completing the proof. ∎
The following is the standard definition of the VC dimention for binary function families:
Definition B.4 (VC Dimension (Vapnik and Chervonenkis, 1971)).
The VC dimension of a family $F$ of binary functions $f:Z\to\{0,1\}$ is the largest size $d$ of a subset $Z^{\prime}=\{z_{1},\ldots,z_{d}\}\subseteq Z$ such that for each of the $2^{d}$ choices of $\bm{s}:=(s_{1},\ldots,s_{d})\in\{0,1\}^{n}$, there exists $f_{\bm{s}}\in F$ such that $f_{\bm{s}}(z_{i})=s_{i}$ for every $i=1,\ldots,d$.
The following standard result can be proved using Dudley’s chaining argument (see e.g. Theorem 8.3.23 of Vershynin (2018)):
Proposition B.5 (Rademacher Complexity from VC Dimension).
Let $F$ be a family of binary functions $f:Z\to\{0,1\}$ with VC dimension at most $d$. Then for any $n\in\mathbb{Z}_{>0}$ and any $z_{1},\ldots,z_{n}\in Z$, we have
$\mathcal{R}(F;z_{1,\ldots,n})\leq O\left(\sqrt{\frac{d}{n}}\right).$