### v2

# A Perfectly Truthful Calibration Measure

Abstract
Calibration requires that predictions are conditionally unbiased and, therefore, reliably interpretable as probabilities. A calibration measure quantifies how far a predictor is from perfect calibration. As introduced by Haghtalab et al. (2024), a calibration measure is truthful if it is minimized in expectation when a predictor outputs the ground-truth probabilities. Predicting the true probabilities guarantees perfect calibration, but in reality, when calibration is evaluated on a random sample, all known calibration measures incentivize predictors to lie in order to appear more calibrated. Such lack of truthfulness motivated Haghtalab et al. (2024) and Qiao and Zhao (2025) to construct approximately truthful calibration measures in the sequential prediction setting, but no perfectly truthful calibration measure was known to exist even in the more basic batch setting.
We design a simple, perfectly and strictly truthful, sound and complete calibration measure in the batch setting: averaged two-bin calibration error (ATB). ATB is quadratically related to two existing calibration measures: the smooth calibration error smCal and the lower distance to calibration $\underline{\textsc{distCal}}$. The simplicity in our definition of ATB makes it efficient and straightforward to compute, allowing us to give the first linear-time calibration testing algorithm, improving a result of Hu et al. (2024). We also introduce a general recipe for constructing truthful measures based on the variance additivity of independent random variables, which proves the truthfulness of ATB as a special case and allows us to construct other truthful calibration measures such as quantile-binned $\ell_{2}$-ECE.
1 Introduction
Probabilistic forecasting has become increasingly important in modern AI-assisted decision-making. Unlike deterministic classification, probabilistic forecasts provide uncertainty quantification, allowing assessment of risks. One desired property of probabilistic predictions is calibration, which requires predictions to be conditionally unbiased and, therefore, reliably interpretable as probabilities. For example, neural networks for tumor diagnosis are trained to output a prediction $r\in[0,1]$, ideally interpretable as the expectation of a binary state $y\in\{0,1\}$: the tumor segment being malignant or not. The neural network is calibrated if, conditioned on the output $r$ being, say $40\%$, the probability that the tumor is malignant $\Pr[y=1|r=40\%]$ is also $40\%$.
[…]
A recent line of work studies the truthfulness of calibration measures (Haghtalab et al., 2024; Qiao and Zhao, 2025). An error measure is truthful if it incentivizes a predictor to output the truth, i.e., the expected error is minimized when the predictor reports the true probabilities. However, no known calibration measure is truthful. Even a miscalibrated predictor can have lower expected error than the truthful predictor when evaluated by known calibration measures. We explain this non-truthfulness in Section 1.1.
1.1 Non-truthfulness of Known Calibration Measures
For existing non-truthful calibration measures, there exists obvious and uninformative prediction strategies that induce a lower calibration error than truth-telling. Concretely, the “obvious” strategy is to ignore all features and always predict the base rate, i.e., output the same constant for every input. In this section, we explain why this obvious strategy has a strictly lower error in expectation, and a lower loss even for every realization of samples. Following this observation, it is well-known that calibration is a poor measure of accuracy (DeGroot and Fienberg, 1983).
Example 1.1 illustrates the non-truthfulness of ECE. We consider the batch setting: a sample of $T$ individuals whose binary states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ are independently drawn from the Bernoulli distributions with means $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ (denoted by $\bm{y}\sim\bm{p}$). We say $\bm{p}=(p_{1},\ldots,p_{T})$ are the ground-truth probabilities. In the example below, a predictor strictly benefits from reporting the base rate of ground truths.
Example 1.1 (ECE is not truthful, c.f. Qiao and Valiant, 2021).
Suppose the ground-truth probabilities are $\bm{p}=(p_{1},\dots,p_{T})$ where each $p_{t}$ is distributed independently and uniformly from $[1/3,2/3]$. An uninformative predictor that always predicts $r_{1}=\cdots=r_{T}=0.5$ achieves an expected empirical $\textsc{ECE}=O\left(\sqrt{1/T}\right)$, the sampling error. However, a truthful predictor who reports $r_{t}=p_{t}$ results in a higher empirical $\textsc{ECE}\geq 1/3$. This is because the predictions $r_{1},\ldots,r_{T}\in[1/3,2/3]$ are almost surely distinct, so the empirical conditional expectation $\widehat{r}_{t}:=\operatorname{\mathbb{E}}_{(r,y)\sim\mathrm{Unif}((r_{t},y_{t})_{t\in[T]})}[y|r=r_{t}]$ is simply $y_{t}\in\{0,1\}$, giving $|r_{t}-\widehat{r}_{t}|=|r_{t}-y_{t}|\geq 1/3$.
Example 1.1 shows that the obvious uninformative prediction achieves a lower ECE. Even worse, there are miscalibrated predictors (e.g. the predictor that always predicts $0.5+\epsilon$ for a small $\epsilon>0$) achieving smaller ECE than the calibrated truthful predictor. Thus ECE do not rank predictors correctly based on how calibrated they are.
Predicting the uninformative base rate incurs a lower calibration error for all known calibration measures. It happens not just for ECE in Example 1.1 and its variants [^1] Variants of ECE include $\ell_{\alpha}$-ECE, where we replace the absolute bias |rr^||r-$\widehat${r}| with $|r-\widehat{r}|^{\alpha}$ for an arbitrary $\alpha\geq 1$, as well as binned versions of $\ell_{\alpha}$-ECE., but also for continuous calibration measures [^2] These are calibration measures that are continuous as a function of the predictions. Note that ECE and binned ECE are not continuous. such as the smooth calibration error (Kakade and Foster, 2008), the distance to calibration (Błasiok et al., 2023) and its variants, etc, irrespective of the sample size $T$. Moreover, it happens consistently accross every realization of the states $\bm{y}$, not just in expectation. Specifically, for every realization of $\bm{y}\in\{0,1\}^{T}$ and every prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$, all these calibration measures Cal satisfy
$\textsc{Cal}(\bm{r},\bm{y})\geq\textsc{Cal}(\bar{\bm{r}},\bm{y}),$ (1)
where $\bar{\bm{r}}=(\bar{r},\ldots,\bar{r})$ is the constant predictor that always predicts the average $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. This obvious and uninformative strategy always achieves (weakly) lower calibration error. For many realizations of states $\bm{y}$, the error is strictly lower. We formally prove this observation in Theorem A.2.
1.2 Our Goal: Perfectly Truthful Calibration Measures
Measuring and optimizing for calibration non-truthfully only makes the predictions less trustworthy, going in the very opposite direction of the goal of calibration. Recall the definition of truthfulness: a calibration measure Cal is truthful if for every ground-truth probabilities $\bm{p}\in[0,1]^{T}$, the expected empirical calibration error $\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ of predictions $\bm{r}\in[0,1]^{T}$ on a random sample $\bm{y}\sim\bm{p}$ is minimized when $\bm{r}=\bm{p}$. From a machine learning perspective, a truthful measure helps identify the Bayes optimal predictor (Gneiting, 2011) because it correctly ranks ground-truth predictions with the lowest expected error. From a game-theoretic perspective, a truthful measure incentivizes an optimizing predictor to output their true beliefs, where we view $\bm{p}$ as the predictor’s subjective belief about the probabilities, which might differ from the true probabilities. For the tumor risk prediction task, if assessed by a non-truthful calibration measure, a doctor is incentivized to report a prediction different from their true probabilistic assessment of the tumor risk, to make the predictions look more calibrated. Such an incentivized misreport can hardly be trusted.
We focus on the perfect truthfulness in the batch setting. Previous work (Haghtalab et al., 2024; Qiao and Zhao, 2025) design approximately truthful calibration measures in the sequential prediction setting, where the states $y_{1},\ldots,y_{T}$ are revealed sequentially after each prediction $r_{t}$ is made. We observe that, in the simpler batch setting, some existing measures are approximately truthful, such as the smooth calibration error (Kakade and Foster, 2008; Błasiok et al., 2023) and the calibration measures proposed by Haghtalab et al. (2024) and Qiao and Zhao (2025). Yet, no known calibration measure is perfectly truthful.
The two minimum requirements of a calibration measure are completeness and soundness. A calibration measure should be able to distinguish calibrated predictors from miscalibrated ones given a sufficiently large sample. Completeness requires vanishing error when a predictor is calibrated (Definition 2.8), and soundness requires non-vanishing error when a predictor is miscalibrated as the sample size increases (Definition 2.9). Haghtalab et al. (2024) point out that some error metrics, such as the well-known squared error $\frac{1}{T}\sum_{t}(r_{t}-y_{t})^{2}$, are truthful but far from being a complete and sound calibration measure. The squared error of a calibrated predictor may not vanish as the sample size $T$ increases.
Our main result shows that truthfulness can be achieved via surprisingly simple constructions in the batch setting, while preserving the completeness and soundness of existing calibration measures.
1.3 A Roadmap of Our Contributions
As our main contribution, we construct a perfectly truthful, complete and sound calibration measure: Averaged Two-Bin Calibration Error ($\mathrm{ATB}$). In addition, ATB maintains and even significantly improves other desired properties considered in the literature: continuity and consistency (Błasiok et al., 2023), low sample complexity, and high computational efficiency.
We propose a general framework for constructing truthful error metrics, where ATB is a special case. This general framework allows us to construct other truthful calibration measures while preserving existing properties. For example, we construct the truthful quantile-binned $\ell_{2}$-ECE with surprisingly simple changes to binned ECE.
We identify key properties that must be satisfied by every truthful measure: it must preserve the Blackwell ordering of calibrated predictors and must (weakly) decrease in expectation when we recalibrate a miscalibrated predictor. Since our $\mathrm{ATB}$ and quantile-binned $\ell_{2}$-ECE are truthful, they inherit these properties and can be used as meaningful and robust evaluation metrics for predictors in practice.
As bonus side products of our simple truthful construction, we give the first linear-time algorithm for the calibration testing problem studied by Hu et al. (2024) as well as a simple, faster constant-factor approximation to the smooth calibration error and the Lower Distance to Calibration.
Perfectly Truthful Calibration Measure.
Section 4 introduces ATB, constructed with two bins and a randomized binning boundary:
  • The prediction space of $[0,1]$ is divided into two bins $[0,q)$ and $[q,1]$, with the bin boundary $q$ chosen uniformly at random from $[0,1]$.
  • Within each bin, we compute the squared error between the sum of the predictions and the sum of the states.
  • After summing up the errors in the two bins and dividing the result by $T^{2}$, we define $\mathrm{ATB}$ to be the expectation over the random choice of the bin boundary $q$.
Definition 1.2 (Averaged two-bin calibration error).
Given predictions $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$, we define
$\mathrm{ATB}(\bm{r},\bm{y})$ $:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T^{2}}\left(\left(\sum_{t:r_{t}<q}(r_{t}-y_{t})\right)^{2}+\left(\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right)^{2}\right)\right].$
We prove the following properties of ATB.
  • Perfect truthfulness (Theorem 4.8). We additionally show that ATB is strictly truthful, meaning that truth-telling is the only minimizer of the expected error, in the ex-ante stage where a predictor is evaluated on a random sample of $T$ i.i.d. individuals.
  • Lipschitz continuity in predictions. For every two prediction sequences $\bm{r}_{1},\bm{r}_{2}\in[0,1]^{T}$,
    $|\mathrm{ATB}(\bm{r}_{1},\bm{y})-\mathrm{ATB}(\bm{r}_{2},\bm{y})|\leq\frac{6}{T}\|\bm{r}_{1}-\bm{r}_{2}\|_{1}.$
  • Completeness and soundness. The completeness and soundness of ATB follow from the quadratic relationship to two known continuous, complete and sound calibration measures: the smooth calibration error smCal (Kakade and Foster, 2008) and the lower distance to calibration $\underline{\textsc{distCal}}$ (Błasiok et al., 2023). See Theorem 4.3.
  • Sample and computational complexity. ATB can be estimated within $\varepsilon$ error using $T=O(\varepsilon^{-2})$ examples in $O(T)$ time. For exact computation of ATB, Theorem 4.12 shows an $O(T\log T)$-time algorithm on $T$ examples. For approximate computation, Theorem 4.12 shows an $O(T+1/\varepsilon)$-time algorithm for approximating up to additive error $\varepsilon>0$. For $\varepsilon\approx 1/\sqrt{T}$, the running time is $O(T+1/\varepsilon)=O(T)$.
    These algorithms are simple and easy to implement, following the simplicity of the definition of ATB. For exact computation, our algorithm sorts the predictions in $O(T\log T)$ time, and does the rest of the computation in $O(T)$ time. For $\varepsilon$-additive approximation, the sorting step can be implemented in $O(T+1/\varepsilon)$ time by discretizing the predictions before sorting. For comparison, the best-known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated and take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively, even when $O(1/\sqrt{T})$ additive error is allowed (Hu et al., 2024).
General Framework for Constructing Truthful Calibration Measures.
We establish the truthfulness of ATB by showing that it is a member of a general family of truthful error metrics, which we term the Unnormalized Binned Squared Errors (UBSEs) (Section 3).
We define UBSEs as binning-based calibration measures where the binning scheme can be randomized and can depend arbitrarily on the predictions $\bm{r}$. The key to its truthfulness lies in how the biases in each bin are combined to calculate the final error. We compute the squared biases in each bin similarly to the standard $\ell_{2}$ ECE, but there is a subtle but crucial difference in how these squared biases are weighted and combined. We discuss this family in more detail in Section 1.4.1 and Section 3. We prove that each UBSE has the following key property of error decomposition, which implies truthfulness.
Lemma 1.3 (Informal, Lemma 3.3).
[…]
To see the truthfulness, when $\bm{r}=\bm{p}$, the middle term is 0, while the third term is invariant of the predictions. The expected error is thus minimized at $\bm{r}=\bm{p}$.
Our construction of $\mathrm{ATB}$ is a special case of UBSE, and thus the truthfulness of $\mathrm{ATB}$ follows immediately from Lemma 1.3. We remark that while every UBSE is truthful, it may not have the other desired properties of ATB, such as soundness, continuity, and the relationship to smCal and $\underline{\textsc{distCal}}$.
To illustrate the idea behind our definition of the UBSEs, we provide a technical overview in Section 1.4.1 and demonstrate, as an example, how to make binned ECE truthful using a quantile-based binning scheme, giving another truthful calibration measure Quantile $\ell_{2}$-BinECE.
Implications of Truthfulness.
Section 5 discusses the implications of a truthful error measure beyond the definition of truthfulness. Following the literature on proper (a.k.a., truthful) loss functions (Gneiting, 2011), we characterize the ranking of predictors by the truthful expected error.
  • Theorem 5.2 shows, among calibrated predictors, a truthful error measure consistently ranks a Blackwell dominating predictor with a (weakly) lower expected error. A Blackwell dominating predictor is more informative and helps any decision maker achieve a lower loss regardless of the loss function. For example, the truth-telling predictor dominates all other predictors.
  • Theorem 5.3 shows, for every miscalibrated predictor, calibrating the predictor always (weakly) reduces the expected truthful error.
Results in Section 5 hold generally for every truthful error measure, not just for sound and complete calibration measures.
Linear-Time Calibration Test via Validity.
We introduce a quantitative framework for evaluating the validity of a calibration error using calibration tests. As the sample size increases, the notions of completeness and soundness require a calibration error to vanish given a calibrated predictor, and not vanish given a mis-calibrated predictor. We define the validity of a calibration error using its ability to distinguish calibration from mis-calibration, which can be viewed as a quantitative combination of completeness and soundness (see Definition 2.17). Any valid calibration measure implies an algorithm for solving the calibration testing problem in Hu et al. (2024).
We show that our $\mathrm{ATB}$ is optimally valid for the smooth calibration error and the lower distance to calibration.
Theorem 1.4 (Informal, see Theorem 6.1).
Given $T$ examples, $\mathrm{ATB}$ can distinguish a calibrated predictor from a predictor with $\textsc{smCal}\geq C/\sqrt{T}$ for an absolute constant $C>0$. Moreover, this rate is information-theoretically optimal (up to constant factors). This also holds when we replace smCal with $\underline{\textsc{distCal}}$ because the two are constant-factor approximations of each other as shown by Błasiok et al. (2023).
As mentioned earlier, ATB can be computed in $O(T\log T)$ time. Thus, Theorem 1.4 implies a faster calibration tester than the previous $O(T\log^{2}T)$-time calibration tester for smCal and $\underline{\textsc{distCal}}$ from Hu et al. (2024). Moreover, the running time can be further improved to $O(T)$ because we do not need to compute ATB exactly and just need to approximate it up to error $O(1/T)$ (see Section 6). This gives the first linear-time algorithm for calibration testing.
Simple and Efficient Constant-factor Approximation of smCal and $\underline{\textsc{distCal}}$.
As mentioned above, the completeness and soundness are proven by the quadratic relationship of ATB to $\textsc{smCal},\underline{\textsc{distCal}}$. This relationship is established via the following non-truthful $\ell_{1}$ variant of $\mathrm{ATB}$, which quadratically approximates $\mathrm{ATB}$. We show that this variant gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$.
Definition 1.5.
We define the $\ell_{1}$ variant of $\mathrm{ATB}$:
$\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})$ $:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-y_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right|\right)\right].$
Theorem 1.6 (Informal, Corollary 4.5).
$\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation to smCal and $\underline{\textsc{distCal}}$:
$\frac{1}{3}\,\underline{\textsc{distCal}}(\bm{r},\bm{y})$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 3\,\underline{\textsc{distCal}}(\bm{r},\bm{y}),$
$\frac{2}{3}\,\textsc{smCal}(\bm{r},\bm{y})$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 6\,\textsc{smCal}(\bm{r},\bm{y}).$
Based on Theorem 1.6, the quadratic relationship between $\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$ follows from the relatively easy observation that $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are themselves quadratically related.
Similar to $\mathrm{ATB}$, $\ell_{1}\textup{-}\mathrm{ATB}$ can also be easily computed in $O(T\log T)$ time and approximated up to $\varepsilon>0$ additive error in $O(T+1/\varepsilon)$ time (see Section 4.5). Therefore, Theorem 1.6 implies a faster constant-factor approximation algorithm for smCal and $\underline{\textsc{distCal}}$ than direct computation, for which the best-known algorithms take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
Previously, Błasiok et al. (2023) also introduced a binning-based approximation to the distance to calibration, termed the interval calibration error. However, the definition of interval calibration error is more involved than $\ell_{1}\textup{-}\mathrm{ATB}$. It requires optimizing the binning scheme (including the number of bins) to minimize the sum of the average bin width and the binned calibration error. Also, the interval calibration error only gives a quadratic approximation for smCal and $\underline{\textsc{distCal}}$. Our Theorem 1.6 shows, perhaps surprisingly, that using only $2$ bins suffices to give a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$ (see Section 1.4.2 for a technical overview).
1.4 Technical Overview
We give a high-level explanation for the two major technical ingredients we use to establish our results. The first is a general recipe for constructing truthful error metrics leveraging the variance additivity of independent random variables. The second is our analysis establishing the constant-factor approximation between $\ell_{1}\textup{-}\mathrm{ATB}$ and the smooth calibration error.
1.4.1 Truthfulness from Variance Additivity
We discuss the idea behind our construction of a general family of truthful measures, i.e., Unnormalized Binned Squared Errors (UBSEs). As mentioned earlier, ATB is a member of this family, so its truthfulness follows as a consequence.
Our goal is to measure the calibration error of a sequence of predictions $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ on the states $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$. Here, each state $y_{t}$ is sampled independently from the Bernoulli distribution with mean $p_{t}\in[0,1]$, where $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ are the true probabilities (denoted by $\bm{y}\sim\bm{p}$).
Our first observation is that, if we divide the predictions $r_{t}$ into bins based on the indices $t\in[T]$ rather than the values $r_{t}\in[0,1]$, then truthfulness can be easily achieved by the $\ell_{2}$ version of binned ECE, $\ell_{2}\textup{-}\textsc{BinECE}$. Concretely, consider a fixed partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ of the index space $[T]$ into bins: $T]=B_{1}\cup\cdots\cup B_{k}$. The $\ell_{2}\textup{-}\textsc{BinECE}$ follows the standard computation of ECE but replacing the $\ell_{1}$ error with squared error:
$\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{r},\bm{y})$ $=\sum_{i\in[k]}\underbrace{\frac{|B_{i}|}{T}}_{\text{weigh by fraction}}\cdot\left(\underbrace{\frac{1}{|B_{i}|}}_{\text{normalize by size}}\underbrace{\sum_{t\in B_{i}}(r_{t}-y_{t})}_{\text{the bias in $B_{i}$}}\right)^{2}$
$=\sum_{i\in[k]}\frac{1}{T|B_{i}|}\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}.$ (2)
Assuming the index partition $\mathcal{B}$ is fixed, the truthfulness of $\ell_{2}\textup{-}\textsc{BinECE}$ comes from the truthfulness of squared error: within each bin $B_{i}$, the expected squared bias over $\bm{y}\sim\bm{p}$
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}\right]$ (3)
is minimized if and only if $\sum_{t\in B_{i}}r_{t}=\sum_{t\in B_{i}}p_{t}$, implying minimized by predicting the truth $r_{t}=p_{t}$.
However, $\ell_{2}\textup{-}\textsc{BinECE}$ in its standard form is not truthful because it does not bin based on indices $t$, but rather the prediction values $r_{t}$. Specifically, $\ell_{2}\textup{-}\textsc{BinECE}$ combines adjacent predictions into the same bin. It works by first partitioning the prediction space $[0,1]$ into intervals, with one bin corresponding to each interval. Each bin consists of the predictions $v_{t}$ that fall within the corresponding interval. A strategic predictor will try to reduce the expected error by manipulating the partition $\mathcal{B}$ via making untruthful predictions. For the same Example 1.1, always predicting the same value $r_{1}=\cdots=r_{T}$ puts all the indices in the same bin, resulting in a different index partition $\mathcal{B}^{\prime}$ than predicting truthfully. This different partition can significantly reduce the expected $\ell_{2}\textup{-}\textsc{BinECE}$ due to the bin-size-based normalization $1/|B_{i}|$ in (2).
The example above hints that we can achieve truthfulness if the expected error of truthful predictions is invaiant to the index partition $\mathcal{B}$. To see this, our analysis above shows that for any index partition $\mathcal{B}^{\prime}$ that could potentially be manipulated from a strategic report $\bm{r}$, predicting the truth achieves smaller or equal expected error on the same index partition $\mathcal{B}^{\prime}$:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{r},\bm{y})].$ (4)
To establish truthfulness, we need to show that the truthful predictions $\bm{p}$ achieve smaller or equal expected error on the un-manipulated partition $\mathcal{B}$ from truthful predictions:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{r},\bm{y})].$ (5)
To bridge the gap between what we have (4) and what we need (5), it suffices if the expected truthful error is invariant of the binned partition for truthful predictions:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}^{\prime}}(\bm{p},\bm{y})]=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\ell_{2}\textup{-}\textsc{BinECE}_{\mathcal{B}}(\bm{p},\bm{y})]\quad\text{for any partitions $\mathcal{B},\mathcal{B}^{\prime}$}.$ (6)
We obtain Equation 6 with our second key observation: we can modify the $\ell_{2}\textup{-}\textsc{BinECE}$ in (2) to achieve partition-invariance by removing the bin-size-based normalization $1/|B_{i}|$. To see this, let us compute the expected squared bias in each bin $B_{i}$ for truthful predictions ($r_{t}=p_{t}$):
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\left(\sum_{t\in B_{i}}(p_{t}-y_{t})\right)^{2}\right]={\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\sum_{t\in B_{i}}y_{t}\right]=\sum_{t\in B_{i}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[y_{t}\right]=\sum_{t\in B_{i}}p_{t}(1-p_{t}),$ (7)
Here we crucially use the variance additivity of independent random variables: since the $y_{t}$’s are independent, the variance of their sum is equal to the sum of their variances. If we directly add up (7) over the bins $i=1,\ldots,k$ without any bin-size-based normalization, we get
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\left(\sum_{t\in B_{i}}(p_{t}-y_{t})\right)^{2}\right]=\sum_{t\in[T]}p_{t}(1-p_{t}),$
which is indeed invariant to the partition $\mathcal{B}$, satisfying (6). Therefore, our analysis shows that the following unnormalized calibration error is truthful, where now the partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ (including the choice of $k$) can (arbitrarily!) depend on the predictions $r_{1},\ldots,r_{T}$:
$\textsc{Cal}(\bm{r},\bm{y}):=\frac{1}{T^{2}}\sum_{i=1}^{k}\left(\sum_{t\in B_{i}}(r_{t}-y_{t})\right)^{2}.$
Moreover, one can show that any calibrated predictions (not just truthful predictions) achieve the same expected error $\frac{1}{T^{2}}\sum_{t\in[T]}p_{t}(1-p_{t})=O\left(\frac{1}{T}\right)$ (see Lemma 3.3), which vanishes as $T\to\infty$, implying the completeness of the error (Definition 2.8) in addition to truthfulness. Note that the global normalization by $1/T^{2}$ ensures completeness while not affecting the truthfulness property, unlike the per-bin normalization in (2) depending on individual bin sizes $|B_{i}|$.
The final observation that completes our construction of UBSEs is that, by linearity of expectation, we can even allow the partition $\mathcal{B}$ to be randomized and use the expected error over the random partition $\mathcal{B}$, without breaking the truthfulness analysis above. This allows us to construct a general family of truthful calibration errors, each using a different, possibility randomized, partition $\mathcal{B}$ that can depend on the predictions $r_{1},\ldots,r_{T}$. We term these calibration errors Unnormalized Binned Squared Errors (UBSEs) and present the formal definition and analysis in Section 3.
Example: Quantile $\ell_{2}\textup{-}\textsc{BinECE}$ is truthful.
As an example of UBSE, simple modifications make $\ell_{2}\textup{-}\textsc{BinECE}$ truthful by binning predictions according to quantiles. With $k$ bins, the following UBSE is truthful.
  • Sort the samples by reported predictions with $r_{1}\leq\dots\leq r_{T}$. Break ties uniformly at random.
  • Divide predictions into $k$ bins, with $\frac{T}{k}$ predictions in each bin.
  • Calculate UBSE.
Binning according to quantiles ensures that each bin contains the same number of predictions and thus, the normalization factors based on bin sizes $|B_{i}|$ in (2) no longer break truthfulness.
1.4.2 Two-Bin Approximation of the Smooth Calibration Error
Our UBSE framework is flexible with regard to how the bins should be chosen (including how many bins should be chosen). However, it is not obvious to find an appropriate binning scheme and show that the corresponding UBSE is polynomially related to existing calibration error metrics such as smCal and $\underline{\textsc{distCal}}$.
Our construction of $\mathrm{ATB}$ is quadratically related to smCal and $\underline{\textsc{distCal}}$. As mentioned earlier, we prove this result by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ (Definition 1.5) gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$. Here we explain the intuition behind this analysis.
Our analysis is divided into the following two results, showing the upper and lower bounds on $\ell_{1}\textup{-}\mathrm{ATB}$ separately:
Lemma 4.6: $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})\leq 3\,\underline{\textsc{distCal}}(\bm{r},\bm{y})$ (8)
Lemma 4.7: $\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\,\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (9)
The desired constant-factor approximation (Theorem 1.6) then follows from the previous result that smCal and $\underline{\textsc{distCal}}$ are themselves constant-factor approximations of each other (Proposition 2.6) (Błasiok et al., 2023).
While neither inequality is straightforward to prove, the relatively more technically involved and, perhaps, more surprising direction is the latter inequality (9) showing that smCal can be upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ up to a constant factor. Indeed, the intuition behind the previous notion of interval calibration error intCal (Błasiok et al., 2023) is that having too few bins tends to underestimate smCal, and if the calibration error is much smaller than the average bin width, we should increase the number of bins to faithfully capture smCal. [^3] Consequently, the number of bins used to define $\textsc{intCal}(\bm{r},\bm{y})$ depends on both $\bm{r}$ and $\bm{y}$. In UBSE, the binning scheme can only depend on $\bm{r}$ in order for our truthfulness analysis to hold. The reasoning is that having fewer bins makes more predictions fall into the same bin, among which the positive and negative biases $r_{t}-y_{t}$ cancel out, thus more likely to cause underestimation. For example, having only one bin gives the following UBSE:
$\textsc{Cal}(\bm{r},\bm{y})=\left(\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right)^{2},$
which clearly underestimates smCal (it can be zero even when $\bm{r}$ is mis-calibrated, in which case $\textsc{smCal}(\bm{r},\bm{y})$ is always positive). Therefore, based on this previous intuition, it is somewhat surprising that having just two bins suffices to establish (9).
Proving (9) is equivalent to showing that for any $1$-Lipschitz weight function $w:[0,1]\to[-1,1]$,
$\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t})\leq\frac{3}{2}\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (10)
This equivalence follows from the definition of smCal: it is the supremum of the left-hand side over all $1$-Lipschitz $w:[0,1]\to[-1,1]$ (Definition 2.5).
Figure 1: Writing $w$ as a convex combination of threshold functions.
To illustrate our proof idea, let us first assume that the weight function $w$ is not only Lipschitz, but also monotonically increasing and differentiable (represented by the curve in Figure 1). The key observation is that we can write $w$ as a convex combination of threshold functions as follows. Take a random threshold $\lambda$ uniformly distributed from [1,1][-1,1] and consider the threshold function $w_{\lambda}(r):=\mathsf{sign}(w(r)-\lambda)$ (represented by the bold step function in Figure 1). That is, $w_{\lambda}(r)=1$ if $w(r)\geq\lambda$, and $w_{\lambda}(r)=-1$ if $w(r)<\lambda$. The following key identity expresses $w$ as a convex combination of the threshold functions $w_{\lambda}$:
$w(r)=\operatorname{\mathbb{E}}_{\lambda\sim\mathrm{Unif}([-1,1])}[w_{\lambda}(r)]\quad\text{for every $r\in[0,1]$}.$ (11)
Now for a fixed threshold $\lambda\in[-1,1]$, let $q:=w^{(-1)}(\lambda)\in[0,1]$ be the corresponding threshold on the $r$-axis, where $w^{(-1)}$ is the inverse of $w$ (see Figure 1). In the boundary cases when $\lambda>w(1)$, we choose $q=1$, and similarly, when $\lambda<w(0)$ we choose $q=0$. This ensures [^4] One tiny caveat which we ignore here is that when $\lambda>w(1)$ and thus $q=1$, this identity does not hold at one point: $r=1$.
$w_{\lambda}(r)=\mathsf{sign}(r-q)\quad\text{for every $r\in[0,1]$}.$ (12)
Let $Q$ be the distribution of the resulting $q$ from $\lambda\sim\mathrm{Unif}([-1,1])$. By (11) and (12), we can rewrite the left-hand side of (10) as
$\frac{1}{T}\sum_{t\in[T]}w(r_{t})\cdot(r_{t}-y_{t})=\operatorname{\mathbb{E}}_{q\sim Q}\left[\frac{1}{T}\sum_{t\in[T]}\mathsf{sign}(r_{t}-q)\cdot(r_{t}-y_{t})\right].$ (13)
For each fixed choice of $q$, it is straightforward to show that the quantity inside the expectation in (13) is upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ at the same fixed bin threshold $q$ (Definition 1.5). However, the random variable $q$ is distributed differently in the two cases. It is drawn from the distribution $Q$ in (13), whereas it is uniformly distributed over $[0,1]$ in the definition of $\ell_{1}\textup{-}\mathrm{ATB}$.
What remains is to relate the two distributions: $Q$ and $\mathrm{Unif}([0,1])$. Recall that $q\sim Q$ is obtained as $q=w^{(-1)}(\lambda)$ for uniformly distributed $\lambda\in[-1,1]$. It follows that the probability density function (PDF) of $q\sim Q$ is exactly the PDF of $\lambda$ (which is $1/2$ everywhere in [1,1][-1,1]) times the derivative $\nabla w(q)$, except at the boundaries $q=0,1$. Since $w$ is $1$-Lipschitz, we have $\nabla w(q)\leq 1$, and thus the PDF of $q\sim Q$ is at most $1/2$ everywhere in the open interval $(0,1)$. This is sufficient to bound the expectation over $q\sim Q$ in (13) by the expectation over $\mathrm{Unif}([0,1])$ in the definition of $\ell_{1}\textup{-}\mathrm{ATB}$ (Definition 4.1). The boundary cases of $q=0,1$ need to be handled separately, but that turns out to be relatively straightforward.
To fully prove (9), we need to remove the monotonicity and differentiability assumptions on $w$, which is achieved by our formal proof in Section 4.1.1. Roughly speaking, without monotonicity, the convex combination of the threshold functions that expresses $w$ might have negative coefficients (so it is a linear combination rather than a convex combination), but the absolute values of the coefficients can still be controlled using the Lipschitzness of $w$. The differentiability assumption can be removed by focusing on the finite set $\{r_{1},\ldots,r_{T}\}$ rather than the full domain $[0,1]$ of $w$.
1.5 Truthfulness and Monotonicity
In this subsection, we include an informal discussion about why it may appear challenging to construct a perfectly truthful calibration measure. This may give a partial explanation for why truthful calibration measures have not appeared before our work, despite the variety of calibration measures introduced in the literature.
At first glance, the existence of a perfectly truthful calibration measure may appear counterintuitive, which arises from the (seemingly) conflicting monotonicity of a truthful error metric and a sound and complete calibration measure. Previous theories on truthful error metrics (a.k.a. proper losses) show that truthfulness is closely tied to convexity and monotonicity. In contrast, the set of calibrated predictors is not convex, and a complete and sound calibration measure should not monotonically decrease as we move the predictions closer to the truth.
More concretely, Lambert (2011) characterizes a truthful error metric to be monotonically decreasing when a prediction moves closer to the truth. Formally, a truthful error of a report must be higher than any convex combination of the report with the state, shown in Figure 2. Calibration, however, specifies a conflicting monotonicity where all calibrated predictors are ranked lowest. As an example, suppose the $T$ realized states are $(0,1,0,1,\dots)$ alternating between 0 and $1$, with an empirical frequency of $50\%$. An uninformative predictor who always predicts $50\%$ should achieve a vanishing calibration error as well as a perfect predictor whose outputs deterministically match the states. Now consider the uniform interpolation between $50\%$ and the ground truth, an alternating prediction sequence of $(25\%,75\%,25\%,\dots)$. Intuitively, this interpolation seems to move “closer” to the perfect predictor. By monotonicity, the error of the interpolation should be upper-bounded by the vanishing error of reporting $50\%$. Yet this interpolation is very miscalibrated, and soundness requires it to receive a high calibration error.
increasing lossrealized statereport
Figure 2: The order sensitivity of a truthful error metric. The large circle is an abstraction of the probabilistic space, with a realized state on a corner of the space. The reported prediction lies in the interior of the space. Fixing the realized state, the truthful error, as a function of the prediction, is increasing along the convex combination from the realized state to the reported prediction. For one binary state prediction, fixing the realized state, a truthful error is monotone in the distance between the reported prediction and the state.
Perhaps due to the reasons above, previous (Haghtalab et al., 2024) and concurrent (Qiao and Zhao, 2025) work focuses on achieving the weaker approximate truthfulness in the more challenging online setting. We show that there exists a perfectly truthful calibration error in the offline setting. The reason behind the counterintuitive possibility of perfect truthfulness is that when the states are drawn independently across individuals (i.e., a product distribution), the monotonicity of a truthful calibration measure is not violated, as the interpolation of two non-trivial product distributions is no longer a product distribution. Thus, the intuition of “moving closer” in the example above does not hold rigorously true. Our truthful construction based on variance additivity indeed crucially uses this independence / product distribution structure.
1.6 Related Work
Truthful Calibration Errors.
Previous work (Haghtalab et al., 2024; Qiao and Zhao, 2025) on approximate truthful calibration errors are closest to our paper. They design multiplicatively truthful calibration error in the sequential prediction problem. In the sequential prediction setting, a sequence of $T$ potentially correlated states is drawn from a distribution. At each period, the predictor predicts, and one state is revealed. Our work studies a different batch setting where all $T$ states are independently drawn and revealed simultaneously after all predictions. An error metric is approximately truthful if predicting the true conditional probability of the next state is a constant approximation of the optimal strategy. Haghtalab et al. (2024) shows that subsampled smooth calibration error is multiplicatively truthful for the sequential prediction setting, implying the smooth calibration error is multiplicatively truthful for the batch setting. Qiao and Zhao (2025) shows that, in the sequential setting, there does not exist a perfectly truthful calibration error that upperbounds the worst-case external regret for decision-makers. The impossibility in sequential setting does not apply to our problem. It also remains open whether there exists a perfectly truthful calibration error metric for the sequential setting while satisfying other completeness and soundness properties.
Calibration Measures.
Foster and Vohra (1997) first proposes the Expected Calibration Error (ECE). The binned ECE serves as a widely-used empirical proxy of ECE (Guo et al., 2017; Minderer et al., 2021). Kleinberg et al. (2023) observes that, if predictions are used for downstream decision-making, ECE upperbounds the swap regret of any downstream decision-maker. Following the decision-making purpose of calibration, Hu and Wu (2024) proposes Calibration Decision Loss (CDL), the worst-case swap regret of any normalized downstream decision task, and shows CDL is quantitatively different from ECE. Okoroafor et al. (2025) introduce the notion of proper calibration as a key ingredient for designing improved algorithms for omniprediction (Gopalan et al., 2022, 2023).
Błasiok et al. (2023) introduced the distance to calibration. In their framework, a calibration error is consistent if it is polynomially related to the distance to calibration. They showed that the smooth calibration error (Kakade and Foster, 2008) and the Laplace kernel calibration error (Kumar et al., 2018) are both consistent, and introduced a binning-based consistent calibration error: the interval calibration error.
Proper Scoring Rules (a.k.a. truthful losses).
Initiated by McCarthy (1956); Savage (1971), extensive work focused on the characterization of proper scoring rules, the class of truthful loss functions. Lambert (2011) characterizes elicitable statistics of a distribution, for example, the mean of a distribution, not the variance of a distribution. Winkler et al. (1996) provides proper scoring rules for the confidence interval, and Frongillo and Kash (2014) provides a characterization of proper scoring rules for eliciting linear properties. Li ⓡ al. (2022) gives computational results of proper scoring rules.
1.7 Paper Organization
The rest of the paper is organized as follows. Section 2 establishes the basic setup, including the definitions of existing calibration errors (Section 2.1), completeness and soundness (Section 2.2), the validity of calibration errors via calibration tests (Section 2.4), and the truthfulness of calibration errors (Section 2.3). In Section 3, we introduce the Unnormalized Binned Squared Errors (UBSEs), a general family of truthful binning-based error metrics. In Section 4, we introduce our proposed calibration error, the Averaged Two-Bin calibration error ($\mathrm{ATB}$) as a special case of UBSE and prove its truthfulness, continuity, sample efficiency, and computationally efficiency. In Section 4.1.1, we prove the quadratic relationship between $\mathrm{ATB}$ and the existing calibration errors $\textsc{smCal},\underline{\textsc{distCal}}$ by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of smCal and $\underline{\textsc{distCal}}$. Section 5 discusses the implications of truthfulness: preservation of Blackwell ordering and monotonicity under recalibration, as well as robustness to hyperparameter selection. In Section 6, we show that $\mathrm{ATB}$ is optimally valid for smCal and $\underline{\textsc{distCal}}$, implying a linear-time calibration tester for smCal and $\underline{\textsc{distCal}}$.
2 Preliminaries
Throughout the paper, we use $D$ to denote a joint distribution of $(x,y)$ pairs, where $x\in X$ represents an individual in a domain $X$, and $y\in\{0,1\}$ is the corresponding state (a.k.a. outcome or label). A predictor $r:X\to[0,1]$ reports a prediction $r(x)\in[0,1]$ for each individual $x\in X$.
We present useful definitions and preliminary theorems for our paper. Section 2.1 introduces existing calibration measures. Section 2.2 defines the completeness and soundness of a calibration measure. Section 2.3 formalizes truthfulness of an error measure. Section 2.4 introduces calibration test, preparing for the result on linear-time calibration tester.
2.1 Calibration
We present the formal definitions of a few important calibration error metrics in the literature. We start with the definition of calibration:
Definition 2.1 (Calibration).
A predictor $r:X\to[0,1]$ is calibrated on an underlying distribution $D$ of $(x,y)\in X\times\{0,1\}$ if $\operatorname{\mathbb{E}}_{D}[y|r(x)]=r(x)$ holds almost surely.
An important property of the definition of calibration is that it only depends on the distribution of the prediction-state pair $(r(x),y)\in[0,1]\times\{0,1\}$. That is, we can determine whether a predictor $r$ is calibrated on a distribution $D$ just based on the distribution of $(r(x),y)$, without having to know the full joint distribution of $(x,r(x),y)$. Thus, using a random variable $v$ to represent the prediction value $r(x)$, we can define calibration simply given a distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$:
Definition 2.2 (Calibration of prediction-state distributions).
We say a distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$ is calibrated if $\operatorname{\mathbb{E}}_{J}[y|v]=v$ holds almost surely.
For a distribution $D$ of $(x,y)\in X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we use $J_{D,r}$ to denote the joint distribution of $(r(x),y)$. With that, $r$ is calibrated on $D$ if and only if $J_{D,r}$ is calibrated as in Definition 2.2.
A calibration measure $\textsc{Cal}_{D}(r)\in\mathbb{R}_{\geq 0}$ evaluates the deviation of a predictor $r$ from being perfectly calibrated on a distribution $D$. Naturally, we define a calibration measure $\textsc{Cal}(J)$ first for general prediction-state distributions $J$ of $(v,y)\in[0,1]\times\{0,1\}$, and then define
$\textsc{Cal}_{D}(r):=\textsc{Cal}(J_{D,r}).$
The most well-known calibration measure is the expected calibration error (ECE):
Definition 2.3 (Expected Calibration Error (ECE) (Foster and Vohra, 1997)).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$, and let random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ be the conditional expectation of the state $y$ given the prediction value $v$. The expected calibration error (ECE) is defined as
$v^|.\textsc{ECE}(J):=\operatorname{\mathbb{E}}$|v-$\widehat${v}|.
Correspondingly, for a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, defining $\widehat{r}(x):=\operatorname{\mathbb{E}}_{D}[y|r(x)]$, we have
$\textsc{ECE}_{D}(r):=\textsc{ECE}(J_{D,r})=\operatorname{\mathbb{E}}_{D}|r(x)-\widehat{r}(x)|.$
More generally, for every $\alpha\geq 1$, we define $\ell_{\alpha}$-ECE as follows:
$\ell_{\alpha}\textup{-}\textsc{ECE}(J):=\operatorname{\mathbb{E}}[|v-\widehat{v}|^{\alpha}],\quad\ell_{\alpha}\textup{-}\textsc{ECE}_{D}(r):=\operatorname{\mathbb{E}}_{D}[|r(x)-\widehat{r}(x)|^{\alpha}].$
A downside of the ECE is its discontinuity: slight changes in the predictions $r(x)$ can cause significant changes to the ECE value. This motivated Błasiok et al. (2023) to introduce a continuous calibration error metric, termed the distance to calibration. It measures the earthmover distance from the prediction-state distribution $(v,y)$ to a calibrated distribution $(u,y)$.
Definition 2.4 ((Lower) Distance to Calibration (Błasiok et al., 2023)).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. Consider a joint distribution (i.e. coupling) $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed according to $J$, and the distribution of $(u,y)$ is calibrated as in Definition 2.2. The (lower) distance to calibration is defined as the following infimum over all such couplings $\Pi$:
$\underline{\textsc{distCal}}(J):=\inf_{\Pi}\operatorname{\mathbb{E}}_{\Pi}|u-v|.$
Correspondingly, given a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we define $\underline{\textsc{distCal}}_{D}(r):=\underline{\textsc{distCal}}(J_{D,r})$.
One might imagine a different definition of the distance to calibration as the minimum $L_{1}$ distance $\operatorname{\mathbb{E}}_{D}|r(x)-r^{\prime}(x)|$ from the given predictor $r$ to a calibrated predictor $r^{\prime}$. Indeed, this notion (denoted distCal) is the first definition of the distance to calibration introduced by Błasiok et al. (2023). However, as shown by Błasiok et al. (2023), this definition is different from the $\underline{\textsc{distCal}}$ in Definition 2.4 and has the disadvantage of depending on the full joint distribution of $(x,r(x),y)$, not just the prediction-state distribution of $(r(x),y)$. To address this disadvantage, Błasiok et al. (2023) introduced the $\underline{\textsc{distCal}}$ in Definition 2.4 and termed it the lower distance to calibration. They also showed that the two definitions are quadratically related:
$\frac{1}{16}\,\textsc{distCal}_{D}(r)^{2}\leq\underline{\textsc{distCal}}_{D}(r)\leq\textsc{distCal}_{D}(r).$
We will focus on the lower distance to calibration in Definition 2.4 throughout the paper and will often drop the word “lower” for brevity.
Another important continuous calibration measure is the smooth calibration error introduced by Kakade and Foster (2008) (originally termed weak calibration). As shown by Błasiok et al. (2023), the smooth calibration error smCal gives a constant factor approximation to $\underline{\textsc{distCal}}$ (see Proposition 2.6 below).
Definition 2.5 (Smooth Calibration Error (Kakade and Foster, 2008)).
Let $W_{1}$ be the family of $1$-Lipschitz functions $w:[0,1]\to[-1,1]$. For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$, the smooth calibration error is defined as
$\textsc{smCal}(J):=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}_{J}[(v-y)w(v)].$ (14)
Correspondingly, for a distribution $D$ of $(x,y)\sim X\times\{0,1\}$ and a predictor $r:X\to[0,1]$, we have
$\textsc{smCal}_{D}(r):=\textsc{smCal}(J_{D,r})=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}_{J}[(r(x)-y)w(r(x))].$
Without the Lipschitzness constraint on $w$, the smooth calibration error would become the ECE (Definition 2.3), where the supremum in (14) is achieved by
$w(v)=\begin{cases}1,&\text{if }\widehat{v}>v;\\ -1,&\text{otherwise.}\end{cases}$
The following proposition shows that $\underline{\textsc{distCal}}$ and smCal are constant factor approximations of each other:
Proposition 2.6 (Błasiok et al. 2023).
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
$\frac{1}{2}\,\underline{\textsc{distCal}}(J)\leq\textsc{smCal}(J)\leq 2\,\underline{\textsc{distCal}}(J).$
2.2 Completeness and Soundness over Samples
A basic property shared by all the calibration measures in Section 2.1 is that they are all minimized when the predictor is calibrated, with the minimum value being zero:
Claim 2.7.
For $\textsc{Cal}\in\{\textsc{ECE},\underline{\textsc{distCal}},\textsc{smCal}\}$, we have $\textsc{Cal}(J)\geq 0$ for any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$. Moreover,
[…]
The claim above tells us that we can determine whether a predictor $r$ is calibrated on a distribution $D$ by checking whether the calibration error $\textsc{Cal}_{D}(r)$ is zero. In practice, however, we rarely get access to the full distribution $D$ and can only compute the calibration error on an i.i.d. sample drawn from $D$. As we shall see, the property in 2.7 does not necessarily generalize to this sample-access scenario, even in the limit of infinitely large sample sizes. To formalize this intuition, we define completeness and soundness below.
Completeness requires that if a predictor is calibrated, the calibration error should vanish as the sample size increases:
Definition 2.8 (Completeness).
We say a calibration measure Cal is complete if the following holds. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is calibrated (see Definition 2.2). Let $S=\{(v_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $J$ (denoted $S\sim J^{T}$), and let $J_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\lim_{T\to\infty}\operatorname{\mathbb{E}}_{S\sim J^{T}}[\textsc{Cal}(J_{S})]=0.$
Conversely, soundness requires that if a predictor is miscalibrated, the calibration error should not vanish as the sample size increases.
Definition 2.9 (Soundness).
We say a calibration measure Cal is sound if the following holds. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is mis-calibrated. (That is, $J$ does not satisfy Definition 2.2.) Let $S=\{(v_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $J$, and let $J_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\liminf_{T\to\infty}\operatorname{\mathbb{E}}_{S\sim J^{T}}[\textsc{Cal}(J_{S}))]>0.$
It can be inferred from the work of Błasiok et al. (2023) that smCal and $\underline{\textsc{distCal}}$ are both complete and sound. However, while ECE satisfies 2.7, it is not complete. To see this, consider the distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, where $v$ is drawn uniformly from $[1/3,2/3]$, and conditioned on $v$, $y$ is drawn from the Bernoulli distribution with mean $v$. Clearly, $J$ is calibrated and $\textsc{ECE}(J)=0$. However, on a finite sample $S=\{(v_{1},y_{1}),\ldots,(v_{T},y_{T})\}$ drawn i.i.d. from $T$, it holds almost surely that all the $v_{t}$’s are distinct, in which case $\textsc{ECE}(J_{S})\geq 1/3$ (see Example 1.1).
Due to ECE’s lack of completeness, in machine learning practice, the binned ECE (BinECE) is widely adopted as an empirical method for estimating ECE (Guo et al., 2017; Minderer et al., 2021). While we do not need this notion to state our main results, we include its definition here for completeness:
Definition 2.10 (Binned ECE).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. Given $\alpha\geq 1$ and a partition $\mathcal{I}=\{I_{i}\}_{i\in[k]}$ of the prediction space $[0,1]$, the $\ell_{\alpha}$-binned ECE is defined as
[…]
We can estimate $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ using a sample $S=\{(v_{t},y_{t})\}_{t\in T}$ of $T$ i.i.d. points drawn from $J$. Specifically, letting $J_{S}$ denote the (empirical) uniform distribution over $S$, we can use $\ell_{\alpha}\textup{-}\textsc{BinECE}(J_{S})$ as a good estimate for $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ when the sample size $T$ is sufficiently large relative to $k$ (the number of bins). In practice, the number $k$ of bins can be selected according to the sample size $T$, e.g. $k=T^{\frac{1}{3}}$, to ensure soundness and completeness.
Remark 2.11 (Comparison to Haghtalab et al. 2024).
Our definition of completeness follows the same idea as Haghtalab et al. (2024), while our soundness is strictly stronger. There exists an error metric that is not reasonably sound but satisfies the completeness and soundness in Haghtalab et al. (2024).
The soundness definition in Haghtalab et al. (2024) requires that for any empirical distribution $D_{T}$ over $T$ samples,
  • if $r_{t}=1-y_{t}$ for all $t$, then $\lim_{T\to\infty}\textsc{Cal}_{T}(r)=\Omega(1)$;
  • if each state $y\sim\text{Ber}(\alpha)$ is independent and identical Bernoulli, then $\lim_{T\to\infty}\textsc{Cal}_{T}(r)=\Omega(1)$ for $r\neq\beta$ being a non-truthful constant.
We see that the error $\textsc{Cal}=({\operatorname{\mathbb{E}}}\left[r\right]-{\operatorname{\mathbb{E}}}\left[y\right])^{2}+{\operatorname{\mathbb{E}}}\left[\mathbb{I}\left[r\in\{0,1\},y\neq r\right]\right]$ satisfies the requirements above. However, for predictions not in $\{0,1\}$, the error metric only evaluates the unconditional bias in predictions, which is far from a calibration error metric.
2.3 Truthfulness
A truthful error metric incentivizes a strategizing predictor to report the true distribution to minimize expected error on a finite sample. Definition 2.12 defines the ex-ante truthfulness where a predictor output is assumed to be function of the feature space.
Definition 2.12 (Ex-Ante Truthfulness).
We say a calibration measure Cal is ex-ante truthful if the following holds. Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$ and let $p:X\to[0,1]$ be the ground-truth predictor $p(x)=\operatorname{\mathbb{E}}_{D}[y|x]$. Let $S=\{(x_{t},y_{t})\}_{t\in[T]}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\quad\text{for every predictor $r:X\to[0,1]$}.$
In this paper, we study a strictly stronger notion: interim truthfulness. In the interim stage, the true distribution of $T$ samples are realized, and the predictor is allowed to deviate and report any prediction sequence, not necessarily a function of the feature space. We first extend our definition of calibration errors to this setting, where we evaluate the calibration error of a reported prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})$ for the $T$ individuals w.r.t. a ground-truth probability sequence $\bm{p}=(p_{1},\ldots,p_{T})$. We will refer to both definition as truthfulness when it is clear from the context.
Definition 2.13 (Induced calibration error on prediction sequences).
Given a calibration measure $\textsc{Cal}(J)$ defined on prediction-state distributions $J$ over $0,1]\times\{0,1\}$, we define an induced calibration measure $\textsc{Cal}(\bm{r},\bm{p})$ as follows, where $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ is a sequence of predictions and $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ is a sequence of ground-truth probabilities. Let $J_{\bm{r},\bm{p}}$ be the distribution of $(r_{t},y)\in[0,1]\times\{0,1\}$ where $t$ is drawn uniformly from $[T]$, and $y\in\{0,1\}$ is drawn from the Bernoulli distribution with mean $p_{t}$. We define
$\textsc{Cal}(\bm{r},\bm{p}):=\textsc{Cal}(J_{\bm{r},\bm{p}}).$
For example, according to Definition 2.13, we can explicitly calculate $\textsc{ECE}(\bm{r},\bm{p})$ and $\textsc{smCal}(\bm{r},\bm{p})$ as follows. Recall that for $v\in\{r_{1},\ldots,r_{T}\}$, we define
$\widehat{v}:=\operatorname{\mathbb{E}}_{(v,y)\sim J_{\bm{r},\bm{p}}}[y|v]=\frac{\sum_{t\in[T]}p_{t}\mathbb{I}\left[r_{t}=v\right]}{\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]}.$ (15)
We have
[…]
$=\frac{1}{T}\sum_{v}\left|(v-\widehat{v})\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]\right|$
$=\frac{1}{T}\sum_{v}\left|\sum_{t\in[T]}(r_{t}-p_{t})\mathbb{I}\left[r_{t}=v\right]\right|.$ (by (15) and $v\mathbb{I}\left[r_{t}=v\right]=r_{t}\mathbb{I}\left[r_{t}=v\right]$)
Similarly for smCal:
$\textsc{smCal}(\bm{r},\bm{p})$ $=\sup_{w\in W_{1}}\frac{1}{T}\sum_{t=1}^{T}(r_{t}-p_{t})w(r_{t}).$ ($W_{1}$ is the same as in Definition 2.5)
We now define the notion of truthfulness for the calibration errors from Definition 2.13 on length-$T$ sequences. We note that this definition is akin to the definition of properness in the literature of proper scoring rules (McCarthy, 1956; Savage, 1971).
Definition 2.14 (Interim Truthfulness).
We say a calibration measure Cal is interim truthful if the following holds for any $T\in\mathbb{Z}_{>0}$. Let $\bm{p}:=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be an arbitrary sequence of ground-truth predictions. Let $\bm{y}=(y_{1},\ldots,y_{t})$ denote the randomly realized states, where each $y_{t}\in\{0,1\}$ is drawn independently from the Bernoulli distribution with mean $p_{i}$ (denoted $\bm{y}\sim\bm{p}$). Then
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})]\leq\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]\quad\text{for any $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$}.$
Claim 2.15 (Interim truthfulness implies ex-ante truthfulness).
Let $\textsc{Cal}(\bm{r},\bm{p})$ be a calibration measure induced by $\textsc{Cal}(J)$ (Definition 2.13). If $\textsc{Cal}(\bm{r},\bm{p})$ is interim truthful, then $\textsc{Cal}(J)$ is ex-ante truthful.
Proof.
As in Definition 2.12, consider a sample $S=\{(x_{1},y_{1}),\ldots,\allowbreak(x_{T},y_{T})\}$ of i.i.d. points from a distribution $D$ over $X\times\{0,1\}$, and let $r:X\to[0,1]$ be a predictor. Define $\bm{r}:=(r(x_{1}),\ldots,r(x_{T}))$ and $\bm{y}:=(y_{1},\ldots,y_{T})$. Now $J_{D_{S},r}$ and $J_{\bm{r},\bm{y}}$ are both equal to the distribution of $(r(x_{t}),y_{t})$ for uniform $t\in[T]$. Therefore,
$\textsc{Cal}_{D_{S}}(r)=\textsc{Cal}(J_{D_{S},r})=\textsc{Cal}(J_{\bm{r},\bm{y}})=\textsc{Cal}(\bm{r},\bm{y}).$ (16)
As in Definition 2.12, define $p(x_{t}):=\operatorname{\mathbb{E}}_{D}[y|x=x_{t}]\in[0,1]$ for $t=1,\ldots,T$. Conditioned on $x_{1},\ldots,x_{T}$, each $y_{t}$ is distributed independently from the Bernoulli distribution with mean $p(x_{t})$. That is, we have $\bm{y}\sim\bm{p}$ as in Definition 2.14, where $\bm{p}:=(p(x_{1}),\ldots,p(x_{T}))$. Therefore, by (16),
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)|x_{1},\ldots,x_{T}]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})],$ (17)
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)|x_{1},\ldots,x_{T}]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})].$ (18)
Assuming interim truthfulness, we know that the quantity in (17) is no smaller than the quantity in (18). Taking the expectation over $x_{1},\ldots,x_{T}$ proves the desired ex-ante truthfulness. ∎
2.4 Calibration Test and Validity
Completeness and soundness (Definitions 2.8 and 2.9) ensure that a calibration measure Cal is able to distinguish calibrated predictors from mis-calibrated ones, when the sample size $T$ is large enough. Intuitively, we should expect the distinguishing power to grow as a function of $T$. We characterize this quantitative dependence on $T$ below. We first define calibration tests that aim at accepting calibrated predictors while rejecting mis-calibrated ones, based on a sample of size $T$.
Definition 2.16 (Calibration Test).
Consider the following calibration test using a calibration measure Cal. Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$. The test first draws $T$ i.i.d. points from $J$ to form a sample $S=\{(v_{t},y_{t})\}_{t\in[T]}$, and then computes the calibration error $\textsc{Cal}(J_{S})$ on the uniform distribution $J_{S}$ over $S$. The test outputs “accept” if the calibration error does not exceed a threshold $\beta$. That is, the acceptance probability of this test is
$\mathsf{accP}^{\textsc{Cal}}(J;T,\beta):=\Pr\nolimits_{S\sim J^{T}}[\textsc{Cal}(J_{S})\leq\beta].$
We define the validity of a calibration measure Cal given a reference calibration measure Ref that is often chosen to be complete and sound.
Definition 2.17 (Validity).
Let $\{\gamma_{T}\}$ be an infinite sequence of real numbers indexed by $T=1,2,\ldots$. We say a calibration measure Cal is $\{\gamma_{T}\}$-valid w.r.t. a reference calibration measure Ref if there exist thresholds $\beta_{1},\beta_{2},\ldots\in\mathbb{R}$ such that
$\liminf_{T\to\infty}\left(\inf_{J\,:\,\text{calibrated}}\mathsf{accP}^{\textsc{Cal}}(J;T,\beta_{T})-\sup_{J\,:\,\textsc{Ref}(J)\geq\gamma_{T}}\mathsf{accP}^{\textsc{Cal}}(J;T,\beta_{T})\right)>0.$
That is, there is a non-vanishing gap between the acceptance probability when $J$ is calibrated, and the acceptance probability when $J$ is mis-calibrated with error at least $\gamma_{T}$ in the reference measure Ref.
In the definition above, one should typically think of $\gamma_{T}$ as a decreasing function of $T$, which indicates the stronger distinguishing power as $T$ grows. Moreover, the faster $\gamma_{T}$ decreases, the stronger is the distinguishing power of a $\{\gamma_{T}\}$-valid calibration error for large $T$.
3 Truthful Family: Unnormalized Binned Squared Errors
In this section, we present a general family of truthful error metrics, which we term unnormalized binned squared errors (UBSEs). As it will become clear, the error $\mathrm{ATB}$ is a special case of UBSEs, so its truthfulness is an immediate consequence of the truthfulness of UBSEs.
Definition 3.1 (Unnormalized Binned Squared Errors).
Consider an error metric $\textsc{Cal}(\bm{r},\bm{p})$ taking as input a report vector $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and a ground-truth vector $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$. We say Cal is an unnormalized binned squared error (UBSE) if it can be calculated as follows:
  1. 1.
    Partition the indices $[T]$ into $k$ disjoint bins: $T]=B_{1}\cup\cdots\cup B_{k}$. Importantly, we allow the partition (including the choice of $k$) to be randomized, and we allow it to depend on the report vector $\bm{r}$ (but not on $\bm{p}$).
  2. 2.
    Compute the bias $\Delta_{i}$ in each bin $B_{i}$:
    $\Delta_{i}:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-p_{t}).$ (19)
  3. 3.
    Output the sum of the squared biases: $\textsc{Cal}(\bm{r},\bm{p}):=\operatorname{\mathbb{E}}_{\mathcal{B}}[\sum_{i=1}^{k}\Delta_{i}^{2}]$, where the expectation is over the randomness of the partition $\mathcal{B}=(B_{1},\ldots,B_{k})$.
The above definition is very similar to the definition of binned $\ell_{2}\text{-}\textsc{ECE}$, but there is a crucial difference. When defining binned $\ell_{2}\text{-}\textsc{ECE}$ for a fixed partition $\mathcal{B}=(B_{1},\ldots,B_{k})$, the bias in each bin is first normalized by the bin size $|B_{i}|$:
$\widetilde{\Delta}_{i}=\frac{1}{|B_{i}|}\sum_{t\in B_{i}}(r_{t}-p_{t}),$
and then squared and summed with weights $|B_{i}|/T$:
$\ell_{2}\text{-}\textsc{ECE}(\bm{r},\bm{p})=\sum_{i=1}^{k}\frac{|B_{i}|}{T}\widetilde{\Delta}_{i}^{2}=\sum_{i=1}^{k}\frac{1}{|B_{i}|T}\left(\sum_{t\in B_{i}}(r_{t}-p_{t})\right)^{2}.$
In contrast, Definition 3.1 takes the unweighted sum of the unnormalized squared biases $\Delta_{i}^{2}$:
$\textsc{Cal}(\bm{r},\bm{p})=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\frac{1}{T^{2}}\left(\sum_{t\in B_{i}}(r_{t}-p_{t})\right)^{2}\right].$
3.1 Interim Truthfulness
UBSEs is interim truthful (whereas the binned $\ell_{2}\text{-}\textsc{ECE}$ is not, with the small difference above):
Theorem 3.2.
Any UBSE error metric Cal is interim truthful (Definition 2.14).
In fact, we prove the a stronger result in Lemma 3.3, showing that the expected empirical UBSE decomposes into the UBSE on the true probabilities $\bm{p}$ plus a variance term independent of $\bm{r}$.
Lemma 3.3 (Error Decomposition).
Let Cal be an arbitrary UBSE. For any report sequence $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and any ground-truth vector $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]=\textsc{Cal}(\bm{r},\bm{p})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$
Here $\bm{y}=(y_{1},\ldots,y_{T})\in\{0,1\}^{T}$ is drawn such that each $y_{t}$ independently follows the Bernoulli distribution with mean $p_{t}$ (as in Definition 2.14).
We first prove Theorem 3.2 using Lemma 3.3, and then prove Lemma 3.3.
Proof of Theorem 3.2.
For any $\bm{r},\bm{p}\in[0,1]^{T}$, by Lemma 3.3,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]-\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{p},\bm{y})]=\textsc{Cal}(\bm{r},\bm{p})-\textsc{Cal}(\bm{p},\bm{p}).$ (20)
Clearly, we have $\textsc{Cal}(\bm{r},\bm{p})\geq 0$ and $\textsc{Cal}(\bm{p},\bm{p})=0$. Therefore, the quantity in (20) is non-negative, which means that Cal is interim truthful. ∎
Proof of Lemma 3.3.
For a partition $\mathcal{B}=(B_{1},\ldots,B_{k})$ of $[T]$ as in Definition 3.1, we define
$\Delta_{i}$ $:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-y_{t}).$
$\widehat{\Delta}_{i}$ $:=\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-p_{t}).$
We have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\textsc{Cal}(\bm{r},\bm{y})]$ $=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]\right]=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]\right],$ (21)
$\textsc{Cal}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{\mathcal{B}}\left[\sum_{i=1}^{k}\widehat{\Delta}_{i}^{2}\right].$ (22)
In (21), we used the fact that the distribution of $\mathcal{B}$ depends only on $\bm{r}$ and not on $\bm{y}$. For the same reason, the two distributions of $\mathcal{B}$ in (21) and (22) are the same. Therefore, to prove the lemma, it suffices to show that for any fixed partition $\mathcal{B}$,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}\left[\sum_{i=1}^{k}\Delta_{i}^{2}\right]=\sum_{i=1}^{k}\widehat{\Delta}_{i}^{2}+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$ (23)
For every $i=1,\ldots,k$, we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}^{2}]=\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}]^{2}+{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\Delta_{i}\right],$ (24)
where
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}]$ $=\widehat{\Delta}_{i},$
${\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\Delta_{i}\right]$ $={\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\frac{1}{T}\sum_{t\in B_{i}}(r_{t}-y_{t})\right]$
$=\frac{1}{T^{2}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[\sum_{t\in B_{i}}y_{t}\right]$
$=\frac{1}{T^{2}}\sum_{t\in B_{i}}{\textsc{Var}}_{\bm{y}\sim\bm{p}}\left[y_{t}\right]$ (the $y_{t}$’s are distributed independently)
$=\frac{1}{T^{2}}\sum_{t\in B_{i}}p_{t}(1-p_{t}).$
Plugging these into (24), we have
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\Delta_{i}^{2}]=\widehat{\Delta}_{i}^{2}+\frac{1}{T^{2}}\sum_{t\in B_{i}}p_{t}(1-p_{t}).$
Summing up over $i=1,\ldots,k$ proves (23). ∎
We remark that in addition to being truthful, UBSEs are also complete. This is because, by Lemma 3.3, the expected error of calibrated predictions is
$\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})\leq\frac{1}{4T}=O(1/T),$
which vanishes as $T\to\infty$.
Example 3.4 (Quantile-Binned $\ell_{2}\text{-}\textsc{ECE}$ is truthful).
As a special case of UBSE, the quantile-binned $\ell_{2}\text{-}\textsc{ECE}$ is truthful and complete. Choosing the number of bins properly as a growing function of $T$, it is also a sound calibration error. It is defined as follows:
For any report sequence $\bm{r}=(r_{1},\dots,r_{T})$ and any vector of realized state $\bm{y}=(y_{1},\dots,y_{T})$,
  • sort the predictions in increasing order with $\bm{r}_{1}\leq\dots\leq\bm{r}_{T}$, break ties uniformly at random.
  • Partition predictions into $k=T^{\nicefrac{{1}}{{3}}}$ bins by quantile. Each bin has $\frac{T}{k}$ predictions.
  • Given the partition above, output the Unnormalized Binned Squared Error $\textsc{Cal}(\bm{r},\bm{y})$.
3.2 Strict Ex-Ante Truthfulness
In the ex-ante stage before the ground-truth probabilities for each sample are drawn, UBSE is strictly truthful, i.e., the unique minimizer to the expected error is the ground truth predictor $p(x):=\operatorname{\mathbb{E}}[y|x]$.
Theorem 3.5.
Let $X$ be an arbitrary non-empty domain and let $D$ be an arbitrary distribution of $(x,y)\in X\times\{0,1\}$. Let $p:X\to[0,1]$ be the ground-truth predictor $p(x):=\operatorname{\mathbb{E}}[y|x]$ and let $r:X\to[0,1]$ be an arbitrary predictor. Let Cal be an arbitrary UBSE and $T$ be an arbitrary positive integer. For a sample $S=((x_{t},y_{t}))_{t\in[T]}$ of $T$ i.i.d. examples drawn from $D$, let $D_{S}$ denote the uniform distribution on $S$. Suppose
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)].$
Then $r(x)=p(x)$ holds almost surely over $(x,y)\sim D$.
We prove Theorem 3.5 using the following lemma. We are able to prove a stronger version of this lemma for the special case of $\mathrm{ATB}$ in Lemma C.1.
Lemma 3.6.
Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$, and define random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ as a function of $v$. Let Cal be an arbitrary UBSE. For $T$ i.i.d. examples $(\widehat{v}_{1},v_{1},y_{1}),\ldots,\allowbreak(\widehat{v}_{T},v_{T},y_{T})$, defining $\widehat{\bm{v}}:=(\widehat{v}_{1},\ldots,\widehat{v}_{T}),\bm{v}:=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]$ $=\frac{1}{T}\operatorname{\mathbb{E}}_{J}[(\widehat{v}-y)^{2}],$ (25)
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})]$ $\geq\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}_{J}[(\widehat{v}-v)^{2}].$ (26)
Proof.
By the definition of $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$, the distribution of $(\widehat{v},y)$ is calibrated. Therefore, $\operatorname{\mathbb{E}}[y_{t}|\widehat{v}_{t}]=\widehat{v}_{t}$ for every $t=1,\ldots,T$. By Lemma 3.3, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]=\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\widehat{v}_{t}(1-\widehat{v}_{t})\right]=\frac{1}{T}\operatorname{\mathbb{E}}[\widehat{v}(1-\widehat{v})]=\frac{1}{T}\operatorname{\mathbb{E}}[(\widehat{v}-y)^{2}],$
where we use the fact that $\widehat{v}_{1},\ldots,\widehat{v}_{T}$ are i.i.d. random variables. This completes the proof of (25).
By Lemma 3.3 again, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})]=\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\widehat{\bm{v}})]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\widehat{v}_{t}(1-\widehat{v}_{t})\right]=\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\widehat{\bm{v}})]+\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})].$ (27)
Since Cal is a UBSE (Definition 3.1) and there are at most $T$ non-empty bins, by the Cauchy-Schwarz inequality, we have
$\textsc{Cal}(\bm{v},\widehat{\bm{v}})\geq\frac{1}{T^{3}}\left(\sum_{t=1}^{T}(v_{t}-\widehat{v}_{t})\right)^{2}.$
Taking expectation, we get
[…]
$\geq\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(v-\widehat{v})^{2}].$
Plugging this into (27) proves (26). ∎
We are now ready to prove Theorem 3.5.
Proof of Theorem 3.5.
Define predictor $\widehat{r}:X\to[0,1]$ such that $\widehat{r}(x)=\operatorname{\mathbb{E}}_{D}[y|r(x)]$. By Lemma 3.6,
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]$ $=\frac{1}{T}\operatorname{\mathbb{E}}_{D}[(p(x)-y)^{2}],$
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]$ $\geq\frac{1}{T}\operatorname{\mathbb{E}}_{D}[(\widehat{r}(x)-y)^{2}]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}_{D}[(r(x)-\widehat{r}(x))^{2}].$
Since $\operatorname{\mathbb{E}}[y|x]=p(x)$, we have
$\operatorname{\mathbb{E}}_{D}[(\widehat{r}(x)-y)^{2}]$ $=\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\operatorname{\mathbb{E}}[(p(x)-y)^{2}]+2\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))(p(x)-y)]$
$=\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\operatorname{\mathbb{E}}[(p(x)-y)^{2}].$
Therefore,
$0$ $\geq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]-\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]$
$\geq\frac{1}{T}(\operatorname{\mathbb{E}}[(\widehat{r}(x)-y)^{2}]-\operatorname{\mathbb{E}}[(p(x)-y)^{2}])+\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(r(x)-\widehat{r}(x))^{2}]$
$=\frac{1}{T}\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(r(x)-\widehat{r}(x))^{2}].$
This implies that $r(x)=\widehat{r}(x)=p(x)$ almost surely. ∎
4 Calibration Errors with Two Bins
In this section, we formally define our calibration measure: the averaged two-bin calibration error (ATB). We show that $\mathrm{ATB}$ satisfies the following properties in the literature: completeness and soundness, truthfulness, continuity, sample complexity for estimation, and computational efficiency. Our proof of the completeness and soundness relies heavily on the quadratic relationship between $\mathrm{ATB}$ and its $\ell_{1}$ variant ($\ell_{1}\textup{-}\mathrm{ATB}$). We will show that $\ell_{1}\textup{-}\mathrm{ATB}$ linearly approximates existing calibration measures, implying the completeness and soundness of both $\ell_{1}\textup{-}\mathrm{ATB}$ and $\mathrm{ATB}$.
Definition 4.1.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we define the averaged two-bin calibration error (ATB) and its $\ell_{1}$ variant as follows:
$\mathrm{ATB}(J)$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\Big(\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v<q\right]\Big]\Big)^{2}+\Big(\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v\geq q\right]\Big]\Big)^{2}\Big],$
$\ell_{1}\textup{-}\mathrm{ATB}(J)$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\Big|\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v<q\right]\Big]\Big|+\Big|\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v\geq q\right]\Big]\Big|\Big].$
Correspondingly, for any prediction sequence $\bm{r}\in[0,1]^{T}$ and ground-truth sequence $\bm{p}\in[0,1]^{T}$,
$\mathrm{ATB}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T^{2}}\left(\left(\sum_{t:r_{t}<q}(r_{t}-p_{t})\right)^{2}+\left(\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right)^{2}\right)\right],$ (28)
$\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-p_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right|\right)\right].$
To prepare for the proof, $\ell_{1}\textup{-}\mathrm{ATB}$ is quadratically related to $\mathrm{ATB}$ by Jensen’s inequality.
Lemma 4.2.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$,
$\frac{1}{2}\,\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}\leq\mathrm{ATB}(J)\leq\ell_{1}\textup{-}\mathrm{ATB}(J).$
Proof.
Fix a threshold $q$, we write $\Delta_{1}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]$ and $\Delta_{2}(q)=\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v\geq q\right]]$. The right inequality follows from the fact that $\Delta_{1},\Delta_{2}\in[-1,1]$.
Using Jensen’s inequality, we get the left inequality:
$\frac{1}{2}\ell_{1}\textup{-}\mathrm{ATB}(J)^{2}=2\left({\operatorname{\mathbb{E}}}_{q}\left[\frac{1}{2}|\Delta_{1}(q)|+\frac{1}{2}|\Delta_{2}(q)|\right]\right)^{2}\leq 2{\operatorname{\mathbb{E}}}_{q}\left[\frac{1}{2}\Delta_{1}(q)^{2}+\frac{1}{2}\Delta_{2}(q)^{2}\right]=\mathrm{ATB}(J).$
4.1 Completeness and Soundness
The completeness and soundness of $\mathrm{ATB}$ follows from the quadratic approximation of existing calibration measure Distance to Calibration.
Theorem 4.3.
Both $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are complete and sound.
Section 4.1.1 will prove that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant approximation of Distance to Calibration. Combined with Lemma 4.2, $\mathrm{ATB}$ is quadratically related to Distance to Calibration, which implies Theorem 4.3.
4.1.1 Approximating the Distance to Calibration Using Two Bins
In this section, we show that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of both smCal and $\underline{\textsc{distCal}}$ (recall Proposition 2.6 that smCal and $\underline{\textsc{distCal}}$ are constant-factor approximations to each other):
Theorem 4.4.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we have
$\frac{2}{3}\,\textsc{smCal}(J)\leq\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J).$
Combining Theorem 4.4 with Proposition 2.6 and Lemma 4.2, we have the following corollary about the relationship between $\mathrm{ATB},\ell_{1}\textup{-}\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$:
Corollary 4.5.
For any distribution $J$ of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, we have
$\frac{1}{3}\,\underline{\textsc{distCal}}(J)\leq\frac{2}{3}\,\textsc{smCal}(J)$ $\leq\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J)\leq 6\,\textsc{smCal}(J),$
$\frac{1}{18}\,\underline{\textsc{distCal}}(J)^{2}\leq\frac{2}{9}\,\textsc{smCal}(J)^{2}$ $\leq\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J)\leq 6\,\textsc{smCal}(J).$
We prove the two inequalities in Theorem 4.4 in two separate lemmas below. We start with the easier one showing the upper bound on $\ell_{1}\textup{-}\mathrm{ATB}$:
Lemma 4.6.
For any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$,
$\ell_{1}\textup{-}\mathrm{ATB}(J)\leq 3\,\underline{\textsc{distCal}}(J).$
Proof.
Let $\Pi$ be an arbitrary distribution of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where the distribution of $(v,y)$ is $J$, and the distribution of $(u,y)$ (denoted by $\widehat{J}$) is calibrated. Since $\widehat{J}$ is calibrated, we have
$\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})=0.$
By Theorem 4.9,
$\ell_{1}\textup{-}\mathrm{ATB}(J)=\ell_{1}\textup{-}\mathrm{ATB}(J)-\ell_{1}\textup{-}\mathrm{ATB}(\widehat{J})\leq 3\operatorname{\mathbb{E}}_{\Pi}|u-v|.$
The lemma is proved by taking the infimum over $\Pi$. ∎
Now we prove the other inequality in Theorem 4.4 showing the lower bound on $\ell_{1}\textup{-}\mathrm{ATB}$. It turns out to be convenient to first focus on the setting with $T$ fixed individuals:
Lemma 4.7.
For any prediction sequence $\bm{r}\in[0,1]^{T}$ and any state sequence $\bm{y}\in\{0,1\}^{T}$, we have
$\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$
Proof of Lemma 4.7.
It suffices to prove that for any $1$-Lipschitz function $w:[0,1]\to[-1,1]$,
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (29)
Assume without loss of generality that the predictions are sorted: $r_{1}\leq\cdots\leq r_{T}$. Define $w(r_{0})=0,w(r_{T+1})=0$. For $t=0,\ldots,T$, define $\Delta_{t}:=w(r_{t+1})-w(r_{t})$. We have
$w(r_{t})=\frac{1}{2}((w(r_{t})-w(r_{0}))-(w(r_{T+1})-w(r_{t})))$ $=\frac{1}{2}\left(\sum_{s<t}\Delta_{s}-\sum_{s\geq t}\Delta_{s}\right)$
$=\frac{1}{2}\sum_{s=0}^{T}\Delta_{s}\mathsf{sign}(t-s),$
where $\mathsf{sign}(u)=1$ if $u>0$, and $\mathsf{sign}(u)=-1$ if $u\leq 0$. Therefore,
$\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})w(r_{t})=\frac{1}{2T}\sum_{s=0}^{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s).$ (30)
For $s=1,\ldots,T-1$, by the Lipschitzness of $w$, we have $|\Delta_{s}|\leq r_{s+1}-r_{s}$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|$ $\leq(r_{s+1}-r_{s})\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\mathsf{sign}(t-s)\right|$
$\leq(r_{s+1}-r_{s})\cdot\frac{1}{T}\left(\left|\sum_{t\leq s}(r_{t}-y_{t})\right|+\left|\sum_{t>s}(r_{t}-y_{t})\right|\right)$
$=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\left[\mathbb{I}_{q\in[r_{s},r_{s+1}]}\cdot\frac{1}{T}\left(\left|\sum_{t:r_{t}<q}(r_{t}-y_{t})\right|+\left|\sum_{t:r_{t}\geq q}(r_{t}-y_{t})\right|\right)\right].$
Summing up over $s=1,\ldots T-1$, we have
$\sum_{s=1}^{T-1}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (31)
Moreover, since $w(r_{1}),w(r_{T})\in[-1,1]$, we have $|\Delta_{0}|,|\Delta_{T}|\leq 1$. Therefore,
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{0}\mathsf{sign}(t-0)\right|$ $=|\Delta_{0}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}),$ (32)
$\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{T}\mathsf{sign}(t-T)\right|$ $=|\Delta_{T}|\cdot\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\right|\leq\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$ (33)
Adding up the three inequalities (31) (32) (33) above, we get
$\sum_{s=0}^{T}\left|\frac{1}{T}\sum_{t=1}^{T}(r_{t}-y_{t})\Delta_{s}\mathsf{sign}(t-s)\right|\leq 3\,\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}).$
Combining this with (30) using the triangle inequality, we get (29), as desired. ∎
Proof of Theorem 4.4.
The upper bound on $\ell_{1}\textup{-}\mathrm{ATB}$ has been proved in Lemma 4.6. It remains to establish the lower bound on $\ell_{1}\textup{-}\mathrm{ATB}$:
$\textsc{smCal}(J)\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(J).$ (34)
Consider a sample $S$ of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. Defining $\bm{r}:=(v_{1},\ldots,v_{T})$ and $\bm{y}:=(y_{1},\ldots,y_{T})$, we have
$\textsc{smCal}(J_{S})$ $=\textsc{smCal}(J_{\bm{r},\bm{y}})=\textsc{smCal}(\bm{r},\bm{y}),$
$\ell_{1}\textup{-}\mathrm{ATB}(J_{S})$ $=\ell_{1}\textup{-}\mathrm{ATB}(J_{\bm{r},\bm{y}})=\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y}),$
where we use the $J_{\bm{r},\bm{y}}$ notation from Definition 2.13. By Lemma 4.7,
$\textsc{smCal}(J_{S})=\textsc{smCal}(\bm{r},\bm{y})\leq\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{y})=\frac{3}{2}\cdot\ell_{1}\textup{-}\mathrm{ATB}(J_{S}).$ (35)
Taking $T\to\infty$, by Theorem 4.10 and Proposition 4.11, we know that $\textsc{smCal}(J_{S})$ converges in probability to $\textsc{smCal}(J)$, and that $\ell_{1}\textup{-}\mathrm{ATB}(J_{S})$ converges in probability to $\ell_{1}\textup{-}\mathrm{ATB}(J)$. Therefore, our goal (34) follows from (35). ∎
4.2 (Strict) Truthfulness
From its definition (28), $\mathrm{ATB}$ is clearly a special case of UBSE (Definition 3.1), so its truthfulness follows immediately from Theorem 3.2 and Theorem 3.5.
Theorem 4.8 (Truthfulness).
The calibration measure $\mathrm{ATB}$ is interim truthful (Definition 2.14). Moreover, $\mathrm{ATB}$ inherits the error decomposition:
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{p}}[\mathrm{ATB}(\bm{r},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{p})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t}).$
Moreover, $\mathrm{ATB}$ is strictly ex-ante truthful.
4.3 Continuity
The following theorem establishes the continuity of $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ with a general formalization. Both errors change continuously as the predictions change.
Theorem 4.9 (Continuity).
Let $\Pi$ be a joint distribution of $(v_{1},v_{2},y)\in[0,1]\times[0,1]\times\{0,1\}$. Let $J_{1}$ denote the distribution of $(v_{1},y)$, and let $J_{2}$ denote the distribution of $(v_{2},y)$. We have
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{1})-\ell_{1}\textup{-}\mathrm{ATB}(J_{2})|$ $\leq 3\,\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (36)
$|\mathrm{ATB}(J_{1})-\mathrm{ATB}(J_{2})|$ $\leq 6\,\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (37)
Proof.
By Definition 4.1, we have
$\ell_{1}\textup{-}\mathrm{ATB}(J_{1})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{1}-y\big)\mathbb{I}\left[v_{1}<q\right]\Big]\Big|+\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{1}-y\big)\mathbb{I}\left[v_{1}\geq q\right]\Big]\Big|\Big],$
$\ell_{1}\textup{-}\mathrm{ATB}(J_{2})$ $=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{2}-y\big)\mathbb{I}\left[v_{2}<q\right]\Big]\Big|+\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{2}-y\big)\mathbb{I}\left[v_{2}\geq q\right]\Big]\Big|\Big].$
We define an intermediate quantity
$\kappa:=\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{2}-y\big)\mathbb{I}\left[v_{1}<q\right]\Big]\Big|+\Big|\operatorname{\mathbb{E}}_{\Pi}\Big[\big(v_{2}-y\big)\mathbb{I}\left[v_{1}\geq q\right]\Big]\Big|\Big].$
By the triangle inequality,
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{1})-\kappa|$ $\leq\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[\operatorname{\mathbb{E}}_{\Pi}\Big[\big|v_{1}-v_{2}\big|\mathbb{I}\left[v_{1}<q\right]\Big]+\operatorname{\mathbb{E}}_{\Pi}\Big[\big|v_{1}-v_{2}\big|\mathbb{I}\left[v_{1}\geq q\right]\Big]\Big]$
$=\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (38)
Similarly, noting that $|v_{2}-y|\leq 1$, we have
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{2})-\kappa|$ $\leq\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big[2\,\operatorname{\mathbb{E}}_{\Pi}\Big|\mathbb{I}\left[v_{1}<q\right]-\mathbb{I}\left[v_{2}<q\right]\Big|\Big]$
$=2\operatorname{\mathbb{E}}_{\Pi}\Big[\operatorname{\mathbb{E}}_{q\sim\mathrm{Unif}([0,1])}\Big|\mathbb{I}\left[v_{1}<q\right]-\mathbb{I}\left[v_{2}<q\right]\Big|\Big]$
$=2\operatorname{\mathbb{E}}_{\Pi}|v_{1}-v_{2}|.$ (39)
Summing up (38) and (39) proves (36). A similar strategy proves (37), using one extra observation: the function $u^{2}$ is $2$-Lipshitz for $u\in[-1,1]$. We omit the details. ∎
4.4 Sample Complexity
Both $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ can be estimated within error $\varepsilon$ using $O(1/\varepsilon^{2})$ i.i.d. examples:
Theorem 4.10 (Sample complexity).
Let $J$ be any distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, and let $S$ be a sample of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. For $\varepsilon,\delta\in(0,1/3)$, assume $T>C\varepsilon^{-2}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$. With probability at least $1-\delta$ (over the randomness in the sample $S$),
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{S})-\ell_{1}\textup{-}\mathrm{ATB}(J))|\leq\varepsilon,$
$|\mathrm{ATB}(J_{S})-\mathrm{ATB}(J)|\leq\varepsilon.$
Proof.
It suffices to show that with probability at least $1-\delta$, for every $q\in[0,1]$,
$\Big|\operatorname{\mathbb{E}}_{J_{S}}\Big[\big(v-y\big)\mathbb{I}\left[v<q\right]\Big]-\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v<q\right]\Big]\Big|$ $\leq\varepsilon/4,\quad\text{and}$
$\Big|\operatorname{\mathbb{E}}_{J_{S}}\Big[\big(v-y\big)\mathbb{I}\left[v\geq q\right]\Big]-\operatorname{\mathbb{E}}_{J}\Big[\big(v-y\big)\mathbb{I}\left[v\geq q\right]\Big]\Big|$ $\leq\varepsilon/4.$
By Proposition D.2, it suffices to prove the following Rademacher complexity bounds for the function families $F=\{f_{q}\}_{q\in[0,1]}$ and $G=\{g_{q}\}_{q\in[0,1]}$ where $f_{q}(v,y)=(v-y)\mathbb{I}\left[v<q\right]$ and $g_{q}(v,y)=(v-y)\mathbb{I}\left[v\geq q\right]$: for every $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$,
$\mathcal{R}(F;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))$ $\leq O\left(\sqrt{\frac{1}{T}}\right),\quad\text{and}$ (40)
$\quad\mathcal{R}(G;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))$ $\leq O\left(\sqrt{\frac{1}{T}}\right).$ (41)
Now consider the family $H=\{h_{q}\}_{q\in[0,1]}$ where $h_{q}(v,y)=\mathbb{I}\left[v<q\right]$. Clearly, $H$ has VC dimension at most $1$. By Proposition D.5, we have
$\mathcal{R}(H;(v_{1},y_{1}),\ldots,(v_{T},y_{T}))\leq O\left(\sqrt{\frac{1}{T}}\right).$ (42)
Observe that $f_{q}(v_{i},y_{i})=\eta_{i}(h_{q}(v_{i},y_{i}))$ for every $i=1,\ldots,T$ and $q\in[0,1]$, where $\eta_{i}$ is the uni-variate function $\eta_{i}(u)=(v_{i}-y_{i})u$ for $u\in\mathbb{R}$. Since $|v_{i}-y_{i}|\leq 1$, the function $\eta_{i}$ is $1$-Lipschitz. Therefore, by Proposition D.3, inequality (40) follows from (42). Inequality (41) can be proved similarly. ∎
We remark that an analogous sample complexity bound for smCal has been shown by Błasiok et al. (2023) using a similar analysis:
Proposition 4.11 (Błasiok et al. (2023)).
Let $J$ be any distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$, and let $S$ be a sample of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ from $J$. For $\varepsilon,\delta\in(0,1/3)$, assume $T>C\varepsilon^{-2}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$. With probability at least $1-\delta$ (over the randomness in the sample $S$),
$|\textsc{smCal}(J_{S})-\textsc{smCal}(J)|\leq\varepsilon.$
4.5 Computational Efficiency
As we show in the following theorem, $\mathrm{ATB}$ can be computed and approximated efficiently.
Theorem 4.12.
Given $\bm{r},\bm{p}\in[0,1]^{T}$, we can compute $\mathrm{ATB}(\bm{r},\bm{p})$ in time $O(T\log T)$. We can also approximate $\mathrm{ATB}(\bm{r},\bm{p})$ up to arbitrary additive error $\varepsilon>0$ in time $O(T+1/\varepsilon)$.
The algorithm we use to prove Theorem 4.12 is extremely easy to describe and implement. Define
$\Delta_{1}(q)=\frac{1}{T}\left|\sum_{t:r_{t}<q}(r_{t}-p_{t})\right|\quad\text{and}\quad\Delta_{2}(q)=\frac{1}{T}\left|\sum_{t:r_{t}\geq q}(r_{t}-p_{t})\right|.$
The following algorithm computes $\mathrm{ATB}$:
  • $O(T\log T)$ time: sort predictions in increasing order such that $r_{1}\leq r_{2}\leq\dots\leq r_{T}$. Define $r_{0}=0$ and $r_{T+1}=1$.
  • $O(T)$ time: for $q=r_{1},\ldots,r_{T+1}$, calculate $\Delta_{1}(q)$ by scanning predictions in increasing order. Similarly, calculate $\Delta_{2}(q)$ by scanning predictions in decreasing order.
  • $O(T)$ time: Calculate the expectation over threshold $q$: for $t=1,\dots,T+1$, sum up $\Delta_{1}(r_{t})^{2}+\Delta_{2}(r_{t})^{2}$ with weight $|r_{t}-r_{t-1}|$.
If we allow additive error $\varepsilon\in(0,1)$, by Theorem 4.9, we can round the predictions $r_{1},\ldots,r_{T}$ to multiples of $\varepsilon/6$ and then compute $\mathrm{ATB}$ exactly. This makes all predictions $r_{1},\ldots,r_{T}$ lie in a finite set $\{0,\varepsilon/6,2\varepsilon/6,\ldots\}\cap[0,1]$ of size $O(1/\varepsilon)$, so the sorting step can be implemented in time $O(T+1/\varepsilon)$ by bucket sort.
A similar algorithm computes $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ in $O(T\log T)$ time, or approximates $\ell_{1}\textup{-}\mathrm{ATB}(\bm{r},\bm{p})$ up to error $\varepsilon$ in $O(T+1/\varepsilon)$ time. We note that currently known algorithms for computing smCal and $\underline{\textsc{distCal}}$ are much more complicated, with the best known running time being $O(T\log^{2}T)$ and $O(T^{2}\log T)$, respectively, even when $O(1/\sqrt{T})$ additive error is allowed (Hu et al., 2024).
5 Implications of Truthfulness
In this section, we discuss the implications of a calibration measure being truthful. These implications follow from truthfulness alone and do not require soundness or completeness. Aside from the obvious game-theoretic implications by definition, we show that a truthful measure preserves the Blackwell dominance between calibrated predictors. Moreover, Lu et al. (2025), a subsequence work of our paper, show that our construction of a truthful calibration measure mitigates the well-known binning-size selection problem of ECE. Results in this section resonate with the work of Gneiting (2011) which discusses implications of properness (a.k.a., truthfulness) of loss functions.
5.1 Preservation of Blackwell Dominance
One motivation for using a truthful loss function is that a truthful loss correctly preserves the Blackwell dominance among predictors. Consider two calibrated predictors $A$ and $B$. AA Blackwell dominates $B$ if $A$ achieves a lower loss for any proper loss function. We show that, similarly, a truthful calibration measure preserves the Blackwell dominance between calibrated predictors (Theorem 5.2). We also show that recalibrating a miscalibrated predictor must (weakly) reduce its expected error as long as we use a truthful measure (Theorem 5.3). We state both theorems in the binary outcome setting ($y\in\{0,1\}$) for consistency with the rest of the paper, but they generalize naturally to the multi-class setting ($y\in\{1,\ldots,k\}$) with essentially the same proofs.
We say a loss $l:[0,1]\times\{0,1\}\to\mathbb{R}$ is a proper loss if it is interim truthful for eliciting prediction on a size-one sample, i.e., for every true probability $p\in[0,1]$ and every non-truthful report $r\in[0,1]$,
${\operatorname{\mathbb{E}}}_{y\sim\mathsf{Ber}(p)}\left[l(p,y)\right]\leq{\operatorname{\mathbb{E}}}_{y\sim\mathsf{Ber}(p)}\left[l(r,y)\right].$
Throughout this section, we will refer to a proper loss as a truthful error measure evaluating the quality of one prediction, which is different from a calibration measure (or an error measure) evaluating multiple samples.
Blackwell (1951) proposes the Blackwell dominance as the ordering among calibrated (Bayesian) predictors. In statistical decision theory, proper loss functions equivalently evaluate the decision loss when a decision-maker best responds to a prediction, assuming the prediction is correct (see more details in Section B.1). If predictor $A$ Blackwell dominates predictor $B$, AA is uniformly better than $B$ for any decision maker with any arbitrary loss function. This partial order serves as a minimal ranking of trustworthiness among calibrated predictors.
Definition 5.1 (Blackwell Dominance).
Assume predictors $A$ and $B$ are both calibrated. Let $J^{A}$ be the joint distribution of $(r^{A},y)\in[0,1]\times\{0,1\}$ for the prediction of predictor $A$ and the state. Similarly, $J^{B}$ for predictor $B$. AA Blackwell dominates $B$ if for any proper loss function $l$,
${\operatorname{\mathbb{E}}}_{J^{A}}\left[l(r^{A},y)\right]\leq{\operatorname{\mathbb{E}}}_{J^{B}}\left[l(r^{B},y)\right].$
By definition, evaluating by a proper loss preserves the Blackwell dominance among calibrated predictors, which has been one of the main justifications for the use of proper losses (Gneiting, 2011).
We show that this Blackwell preserving property is closed under the composition of samples. Given a truthful error measure on multiple i.i.d. samples from the same predictor, every truthful error measure preserves the same Blackwell dominance.
Theorem 5.2.
Let Cal be a calibration measure that is ex-ante truthful (Definition 2.12). Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$, and let $r^{A},r^{B}:X\to[0,1]$ be two calibrated predictors w.r.t. $D$. Assume that $r^{A}$ Blackwell dominates $r^{B}$. Let $S=\{(x_{t},y_{t})\}_{t\in[T]}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{A})]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{B})].$
The proof is deferred to Section B.2.
Theorem 5.2 shows that, for any two calibrated predictors with Blackwell dominance, a truthful error measure preserves the dominance in expectation. Theorem 5.3 furthur characterizes the ordering between calibrated predictors and miscalibrated predictors. Specifically, calibrating a predictor always reduces the expected error measure for a truthful measure.
Theorem 5.3 (Recalibration reduces expected truthful error).
Let Cal be an error measure that is ex-ante truthful (Definition 2.12). Let $J$ be a joint distribution of $(v,y)\in[0,1]\times\{0,1\}$ and define random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ as a function of $v$ (by calibrating $v$). For $T$ i.i.d. examples $(\widehat{v}_{1},v_{1},y_{1}),\ldots,\allowbreak(\widehat{v}_{T},v_{T},y_{T})$, defining $\widehat{\bm{v}}:=(\widehat{v}_{1},\ldots,\widehat{v}_{T}),\bm{v}:=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]\leq\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})].$
The proof is deferred to Section B.3.
In summary, a truthful error measure preserves the ranking among predictors in the following way.
  • For calibrated predictors, the expected truthful error measure preserves Blackwell ordering.
  • For a miscalibrated predictor, calibrating reduces the truthful error measure in expectation.
Suppose $A$, $B$, and $C$ are predictors such that $B$ is obtained by calibrating $C$, and $A$ Blackwell dominates $B$. With a truthful error measure, the expected error is ordered increasingly by $A$, $B$, and then $C$. As shown in Example 1.1, if evaluated by a non-truthful error measure, both $B$ and $C$ can have a lower (better) error than the dominating predictor $A$, on contrary.
5.2 Robustness to Hyperparameter Selection
One widely observed problem with binned ECE is that the comparison among predictors may flip when the error is estimated with different binning sizes (see, e.g., Minderer et al. 2021). Lu et al. (2025) replicate this non-robustness, which we show in Figure 3(a). This hyperparameter of binning size induces contradicting conclusions about the relationship between the loss (quality) and the calibration error of a predictor.
Subsequent work of ours (Lu et al., 2025) empirically shows that ranking by a truthful calibration measure is robust to hyperparameter selection. A truthful calibration measure preserves the Blackwell ordering among calibrated predictors, which is invariant of the hyperparameter selection. Lu et al. (2025) adopt our truthful Quantile $\ell_{2}\textup{-}\textsc{BinECE}$. They generalize our calibration measure for binary classes to multi-class classification tasks by taking the average over each dimension of prediction.
Lu et al. (2025) explain this robustness as a result of preserving Blackwell dominance, which is invariant of the hyperparameter selection. First, neural network (NN) predictors are approximately calibrated (close to calibration); Second, empirically, there exists a Blackwell dominance between NN predictors across different proper losses. Lastly, for any pairs of (approximately calibrated) predictor with Blackwell dominance, any truthful calibration measure preserves the same ordering.
(a) This plot uses a non-truthful calibration measure, the BinECE. The left estimates ECE with $m=2000$ bins, and more accurate predictors have lower calibration error. The right estimates with $m=20$ bins, and more accurate predictors have higher calibration error.
(b) This plot uses a truthful calibration measure, the Quantile $\ell_{2}\textup{-}\textsc{BinECE}$. The left plot estimates the calibration error with $m=2000$ bins, and the right plot estimates with $m=20$ bins. The ranking between predictors is more consistent for different numbers of bins.
Figure 3: The comparison of calibration measures with different number of bins. Each dot in the plot is a predictor. The $x$-axis plots the log loss (prediction quality), while the $y$-axis plots a calibration error. Figure 3(a) replicates the result in Minderer et al. (2021). The plots are adopted from Lu et al. (2025) with permission.
6 Linear-Time Calibration Tester
In this section, we show that our $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ are both optimally valid (Definition 2.17) for smCal and $\underline{\textsc{distCal}}$. It is fairly straightforward to show that $\ell_{1}\textup{-}\mathrm{ATB}$ is $O(1/\sqrt{T})$-valid using its constant approximation to smCal (Theorem 4.4) and its sample complexity bound (Theorem 4.10). In Theorem 6.1 below, we show that $\mathrm{ATB}$ is $O(1/\sqrt{T})$-valid as well, and that this is optimal up to constant.
These results imply faster algorithms for solving the calibration testing problem studied by Hu et al. (2024), which requires distinguishing, with large constant success probability, whether a distribution $J$ is perfectly calibrated or has $\underline{\textsc{distCal}}(J)>\varepsilon$ given i.i.d. data points drawn from $J$. This can be solved by computing $\mathrm{ATB}$ or $\ell_{1}\textup{-}\mathrm{ATB}$ on $T=O(1/\varepsilon^{2})$ data points and compare the result with the threshold $\beta_{T}$ in the definition of validity (Definition 2.17). By Theorem 4.12, the running time we need is $O(T\log T)$, which already improves the $O(T\log^{2}T)$ time bound of Hu et al. (2024). Moreover, our Lemmas 6.2 and 6.3 show that it suffices to approximate $\mathrm{ATB}$ up to additive error $1/(2T)$, which can be achieved in time $O(T)$ by Theorem 4.12, giving the first linear-time algorithm for calibration testing.
Theorem 6.1.
The calibration measure $\mathrm{ATB}$ is $O(\frac{1}{\sqrt{T}})$-valid w.r.t. the reference calibration error $\underline{\textsc{distCal}}$. That is, $\mathrm{ATB}$ is $\{\gamma_{T}\}$-valid for some sequence $\gamma_{1},\gamma_{2},\ldots$ with $\gamma_{T}=O(1/\sqrt{T})$. Moreover, this is optimal up to constant factors: if there exists a $\{\gamma_{T}\}$-valid calibration error w.r.t. $\underline{\textsc{distCal}}$, then $\gamma_{T}=\Omega(1/\sqrt{T})$.
Theorem 6.1 is an immediate corollary of the following Lemmas 6.2, 6.3, and 6.4.
Lemma 6.2.
Let $J$ be an arbitrary distribution of prediction-state pairs $(v,y)\in[0,1]\times\{0,1\}$ and assume that $J$ is calibrated. For any $T\in\mathbb{Z}_{>0}$, consider a sample $S$ of $T$ i.i.d. points $(v_{1},y_{1}),\ldots,(v_{T},y_{T})\in[0,1]\times\{0,1\}$ from $J$, and let $J_{S}$ be the uniform distribution over $S$. We have
$\Pr_{S\sim J^{T}}[\mathrm{ATB}(J_{S})\leq 1/T]\geq 3/4.$
Proof.
Define $\bm{r}=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$. It is clear that the distribution $J_{\bm{r},\bm{y}}$ (see Definition 2.13) is equal to the distribution $J_{S}$. Therefore,
$\mathrm{ATB}(J_{S})=\mathrm{ATB}(\bm{r},\bm{y}).$
Since $J$ is calibrated, we have $\operatorname{\mathbb{E}}_{J}[y|v=v_{t}]=v_{t}$ for every $t=1,\ldots,T$. Conditioned on $\bm{r}=(v_{1},\ldots,v_{T})$, each $y_{t}$ is independently distributed as the Bernoulli distribution with mean $v_{t}$. Thus, we have $\bm{y}\sim\bm{r}$ as in Definition 2.14. Therefore,
$\Pr_{S}[\mathrm{ATB}(J_{S})\leq 1/T\ |\ v_{1},\ldots,v_{T}]=\Pr_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})\leq 1/T].$ (43)
By Lemma 3.3,
$\operatorname{\mathbb{E}}_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})]=\mathrm{ATB}(\bm{r},\bm{r})+\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})=\frac{1}{T^{2}}\sum_{t=1}^{T}p_{t}(1-p_{t})\leq\frac{1}{4T}.$
By Markov’s inequality,
$\Pr_{\bm{y}\sim\bm{r}}[\mathrm{ATB}(\bm{r},\bm{y})\leq 1/T]\geq 3/4.$
Plugging this into (43) and taking the expectation over $v_{1},\ldots,v_{T}$ completes the proof. ∎
Lemma 6.3.
There exists an absolute constant $C>0$ such that the following holds. For any $T\in\mathbb{Z}_{>0}$ and any distribution $J$ of $(v,y)\in[0,1]\times\{0,1\}$ with $\underline{\textsc{distCal}}(J)\geq C/\sqrt{T}$, let $S$ be a sample of $T$ i.i.d. points from $J$. Then
$\Pr_{S\sim J^{T}}[\mathrm{ATB}(J_{S})\leq 2/T]\leq 1/4.$
Proof.
By Theorem 4.10, there exists an absolute constant $C^{\prime}>0$ such that with probability at least $3/4$ over $S\sim J^{T}$,
$|\ell_{1}\textup{-}\mathrm{ATB}(J_{S})-\ell_{1}\textup{-}\mathrm{ATB}(J)|\leq C^{\prime}/\sqrt{T}.$ (44)
It remains to show that whenever (44) holds, we have
$\mathrm{ATB}(J_{S})>2/T.$
By Corollary 4.5 and our assumption that $\underline{\textsc{distCal}}(J)>C/\sqrt{T}$, we have $\ell_{1}\textup{-}\mathrm{ATB}(J)\geq(C/3)/\sqrt{T}$. Therefore, whenever (44) holds, we have
$\ell_{1}\textup{-}\mathrm{ATB}(J_{S})\geq(C/3-C^{\prime})/\sqrt{T}.$
Assuming $C/3-C^{\prime}>0$ which is guaranteed by a sufficiently large $C$, by Lemma 4.2, we have
$\mathrm{ATB}(J_{S})\geq(1/2)(C/3-C^{\prime})^{2}/T.$
The proof is completed by choosing $C$ large enough so that $(1/2)(C/3-C^{\prime})^{2}>2$. ∎
Lemma 6.4.
Let $\{\gamma_{T}\}_{T=1,2,\ldots}$ be a sequence of nonnegative real numbers such that there exists a $\{\gamma_{T}\}$-valid calibration error Cal w.r.t. $\underline{\textsc{distCal}}$. Then $\gamma_{T}=\Omega(1/\sqrt{T})$.
Proof.
Let us focus on the choices of $T$ such that $\gamma_{T}<1/2$. We define $J_{1}$ to be the uniform distribution over $\{(1/2,0),(1/2,1)\}\subseteq[0,1]\times\{0,1\}$. We define $J_{2}$ to be the distribution with probability mass $1/2-\gamma_{T}$ on $(1/2,0)$, and the remaining probability mass $1/2+\gamma_{T}$ on $(1/2,1)$.
Clearly, $J_{1}$ is calibrated. We claim that $\underline{\textsc{distCal}}(J_{2})\geq\gamma_{T}$. Indeed, consider any coupling distribution $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed as $J_{2}$, and the distribution of $(u,y)$ is calibrated. By calibration, $\operatorname{\mathbb{E}}[u]=\operatorname{\mathbb{E}}[y]=1/2+\gamma_{T}$. Therefore, $\operatorname{\mathbb{E}}|u-v|\geq\operatorname{\mathbb{E}}[u]-\operatorname{\mathbb{E}}[v]=\gamma_{T}$, implying that $\underline{\textsc{distCal}}(J_{2})\geq\gamma_{T}$.
Let $\beta_{T}\in\mathbb{R}$ be the threshold satisfying the requirement of validity (Definition 2.17). Define
$\delta_{T}:=\mathsf{accP}^{\textsc{Cal}}(J_{1};T,\beta_{T})-\mathsf{accP}^{\textsc{Cal}}(J_{2};T,\beta_{T}).$
Note that the two acceptance probabilities above are w.r.t. the randomness in the samples $S_{1}\sim J_{1}^{T}$ and $S_{2}\sim J_{2}^{T}$, respectively, where $J_{1}^{T}$ (resp. $J_{2}^{T}$) is the joint distribution of $T$ i.i.d. points from $J_{1}$ (resp. $J_{2}$). A standard argument (e.g. via Pinsker’s inequality) shows that the total variation distance between $J_{1}^{T}$ and $J_{2}^{T}$ is $O(\gamma_{T}\sqrt{T})$. Therefore,
$\delta_{T}\leq O(\gamma_{T}\sqrt{T}).$
Validity requires $\liminf_{T\to\infty}\delta_{T}>0$. Therefore,
$\liminf_{T\to\infty}\gamma_{T}\sqrt{T}>0.$
This implies $\gamma_{T}=\Omega(1/\sqrt{T})$. ∎
Appendix A Non-Truthfulness of Known Calibration Measures
In this section, we prove Theorem A.2 showing that condition (1) (restated below as condition (45)) holds for a broad family of calibration measures: $\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal}$ and $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$ (Definition A.1), where $\alpha\geq 1$ is arbitrary. Condition (1) states that, for any realization of the states, reporting the average over predictions is weakly better than reporting truthfully for known calibration measures. By Remark A.3, this proves that these calibration measures are not truthful.
Definition A.1 ($\ell_{\alpha}$-Distance to Calibration).
Let $J$ be a distribution of $(v,y)\in[0,1]\times\{0,1\}$. We define its $\ell_{\alpha}$-distance to calibration (denoted by $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}(J)$) similarly to the definition of $\underline{\textsc{distCal}}(J)$ in Definition 2.4. The only difference is that we change the $\ell_{1}$ distance |uv||u-v| to $|u-v|^{\alpha}$:
$\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}(J):=\inf_{\Pi}\operatorname{\mathbb{E}}_{\Pi}[|u-v|^{\alpha}],$
where the infimum is over joint distributions $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where $(v,y)$ is distributed according to $J$, and the distribution of $(u,y)$ is calibrated.
Theorem A.2.
Let Cal be a calibration measure from $\{\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal},\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}\}$, where $\alpha\geq 1$ is arbitrary. For every $\bm{r}=(r_{1},\ldots,r_{T})\in[0,1]^{T}$ and every $\bm{y}\in\{0,1\}^{T}$, it holds that
$\textsc{Cal}(\bar{\bm{r}},\bm{y})\leq\textsc{Cal}(\bm{r},\bm{y}),$ (45)
where $\bar{\bm{r}}:=(\bar{r},\ldots,\bar{r})\in[0,1]^{T}$ for $\bar{r}:=\frac{1}{T}\sum_{t=1}^{T}r_{t}$.
Remark A.3.
It is very easy to find (many) examples of $\bm{y}$ where the inequality (45) becomes strict, in which case reporting $\bar{\bm{r}}$ instead of the $\bm{r}$ is strictly better (i.e. $\bar{\bm{r}}$ dominates $\bm{r}$). In particular, we can find many examples of $\bm{y}$ where $\textsc{Cal}(\bm{r},\bm{y})>0$ and $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$. This is because for all of the calibration measures Cal mentioned above (including the continuous ones), we always have $\textsc{Cal}(\bm{r},\bm{y})>0$ as long as $r_{1},\ldots,r_{T}$ are distinct values [^5] For binned calibration error, we need $\max_{t}r_{t}-\min_{t}r_{t}$ to be sufficiently large so that $r_{1},\ldots,r_{T}$ do not fall in the same bin. in, say, $[1/3,2/3]$, and we have $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$ as long as the average outcome $\frac{1}{T}\sum_{t=1}^{T}y_{t}$ is equal to $\bar{r}$.
Theorem A.2 is a corollary of the following theorem.
Theorem A.4.
Let Cal be a calibration measure from $\{\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal},\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}\}$, where $\alpha\geq 1$ is arbitrary. Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$. We have
$\textsc{Cal}(\bar{J})\leq\textsc{Cal}(J).$ (46)
Here $\bar{J}$ is the distribution of $(\bar{v},y)\in[0,1]\times\{0,1\}$, where we draw $(v,y)\sim J$ and replace $v$ with the deterministic quantity $\bar{v}:=\operatorname{\mathbb{E}}_{J}[v]$.
Proof.
We prove the theorem separately for each choice of Cal. Similarly to the definition of $\bar{v}$, we define $\bar{y}:=\operatorname{\mathbb{E}}_{J}[y]$.
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\textsc{ECE}$, defining $\widehat{v}:=\operatorname{\mathbb{E}}[y|v]$, by Jensen’s Inequality we have
$\textsc{ECE}(J)=\operatorname{\mathbb{E}}[|v-\widehat{v}|^{\alpha}]\geq|\operatorname{\mathbb{E}}[v-\widehat{v}]|^{\alpha}=|\bar{v}-\bar{y}|^{\alpha},$ (47)
where we used the fact that $\operatorname{\mathbb{E}}[\widehat{v}]=\operatorname{\mathbb{E}}[y]=\bar{y}$. Also,
[…]
where we used the fact that $\bar{v}$ is a deterministic quantity, so $\operatorname{\mathbb{E}}[y|\bar{v}]=\operatorname{\mathbb{E}}[y]=\bar{y}$. Combining (47) and (48) proves (46).
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\textsc{BinECE}$, we can prove (46) as follows. Let $\mathcal{I}=\{I_{i}\}_{i\in[k]}$ be the partition of the prediction space $[0,1]$ in the definition of $\ell_{\alpha}\textup{-}\textsc{BinECE}$ (Definition 2.10). By Jensen’s Inequality,
[…]
$=|\operatorname{\mathbb{E}}_{J}[v-y]|^{\alpha}$
$=|\bar{v}-\bar{y}|^{\alpha}.$ (49)
Also, since $\bar{v}$ is deterministic, we have
[…]
Combining (49) and (50) proves (46).
When $\textsc{Cal}=\textsc{smCal}$, (46) follows from the following calculation:
$\textsc{smCal}(\bar{J})$ $=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}[(\bar{v}-y)w(\bar{v})]=\sup_{w\in W_{1}}(\bar{v}-\bar{y})w(\bar{v})=|\bar{v}-\bar{y}|,$
$\textsc{smCal}(J)$ $=\sup_{w\in W_{1}}\operatorname{\mathbb{E}}[(v-y)w(v)]\geq\sup_{\sigma\in\{\pm 1\}}\operatorname{\mathbb{E}}[(v-y)\sigma]=\sup_{\sigma\in\{\pm 1\}}(\bar{v}-\bar{y})\sigma=|\bar{v}-\bar{y}|.$
When $\textsc{Cal}=\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$, (46) holds because of the following argument. Consider any joint distribution $\Pi$ of $(u,v,y)\in[0,1]\times[0,1]\times\{0,1\}$, where the marginal distribution of $(v,y)$ is $J$, and the marginal distribution of $(u,y)$ is calibrated. By Jensen’s Inequality,
$\operatorname{\mathbb{E}}_{\Pi}[|u-v|^{\alpha}]\geq|\operatorname{\mathbb{E}}_{\Pi}[u-v]|^{\alpha}=|\operatorname{\mathbb{E}}_{\Pi}[u]-\operatorname{\mathbb{E}}_{\Pi}[v]|^{\alpha}=|\operatorname{\mathbb{E}}_{\Pi}[y]-\operatorname{\mathbb{E}}_{\Pi}[v]]|^{\alpha}=|\bar{y}-\bar{v}|^{\alpha}.$
We used the fact that $\operatorname{\mathbb{E}}_{\Pi}[u]=\operatorname{\mathbb{E}}_{\Pi}[y]$, which holds because the distribution of $(u,y)$ is calibrated. Taking the infimum over $\Pi$, we get
$\underline{\textsc{distCal}}(J)\geq|\bar{y}-\bar{v}|^{\alpha}.$ (51)
Moreover, since always predicting $\bar{y}$ yields a calibrated predictor, we can choose $\Pi$ to be the joint distribution of $(\bar{y},\bar{v},y)$ and get
$\underline{\textsc{distCal}}(\bar{J})\leq\operatorname{\mathbb{E}}_{\Pi}[|\bar{y}-\bar{v}|^{\alpha}]=|\bar{v}-\bar{y}|^{\alpha}.$ (52)
Combining (51) and (52) proves (46). ∎
Table 1 provides an example illustrating the non-truthfulness of known calibration measures and the truthfulness of our $\mathrm{ATB}$. The table compares two strategies: predicting the overall average and predicting truthfully.
Prob. States smCal $\underline{\textsc{distCal}}$ $\ell_{2}\text{-}\underline{\textsc{distCal}}$ $\mathrm{ATB}$ (ours)
avg truth avg truth avg truth avg truth
$\frac{3}{16}$ $(0,0)$ 0.5 0.5 0.5 0.5 0.25 0.3125 0.25 0.203125
$\frac{3}{16}$ $(1,1)$ 0.5 0.5 0.5 0.5 0.25 0.3125 0.25 0.203125
$\frac{9}{16}$ $(0,1)$ 0 $0.0625$ 0 >0>0 0 >0>0 0 0.015625
$\frac{1}{16}$ $(1,0)$ 0 $0.1875$ 0 >0>0 0 >0>0 0 0.140625
Expected Error 0.1875 0.234375 0.1875 >0.1875>0.1875 0.09375 >0.11>0.11 0.09375 0.09375
Table 1: The calibration errors of predictors with two data points. The ground truth probabilities of the two points are $25\%$ and $75\%$, respectively. In the table, avg stands for the uninformative predictor that always outputs $50\%$ and truth stands for the truthful predictor that outputs $25\%$ and $75\%$. We calculate the error of the predictors given each realization of the state and the total expected error. For non-truthful error metrics, the expected error of a truthful predictor is strictly higher than the expected error of an uninformative predictor. For $\mathrm{ATB}$, the expected errors are the same.
Appendix B Deferred Proofs in Section 5
B.1 Decision-Theoretic Explanation of Blackwell Dominance
In this section, we explain the connection of proper losses and decision making. A proper loss evaluates the loss of a decision maker when they best responds to a prediction.
Definition B.1 (Decision problem).
A decision problem consists of an outcome space $\mathcal{Y}$, an action space $A$, and a loss
$L:\;A\times\mathcal{Y}\to\mathbb{R},\qquad(a,y)\mapsto L(a,y).$
Given a prediction $p\in\Delta(\mathcal{Y})$, the expected loss of an action $a\in A$ under $p$ is
$R_{L}(a;p)\;:=\;\operatorname{\mathbb{E}}_{y\sim p}\!\big[L(a,y)\big].$
A best response to $p$ is any action
$\text{BR}_{L}(p)\;\in\;\operatorname*{arg\,min}_{a\in A}R_{L}(a;p),$
and the best-responding loss, which we will refer to as the induced loss is a function of the prediction and the outcome
$l_{L}(p,y)\;:=\;L\big(\text{BR}_{L}(p),\,y\big),\qquad(p,y)\in\Delta(\mathcal{Y})\times\mathcal{Y}.$
Theorem B.2 shows that the space of induced best-responding loss is the same as the space of proper losses. The proof of this connection can be found in previous work, e.g., Frongillo and Kash (2014); Kleinberg et al. (2023); Hu and Wu (2024).
Theorem B.2 (Induced Losses $\Leftrightarrow$ Proper Losses).
Induced losses and proper losses are equivalent in the following sense:
  • For any decision problem $(A,\mathcal{Y},L)$, the induced loss $l_{L}(p,y)=L(\text{BR}_{L}(p),y)$ is proper.
  • Conversely, for any proper loss $l$, there exists a decision problem $(A,\mathcal{Y},L)$ whose induced loss equals $l$; in particular, one can take $A=\Delta(\mathcal{Y})$ and $L(a,y)=l(a,y)$, so that $l_{L}=l$. If $l$ is strictly proper, then $\text{BR}_{L}(p)=\{p\}$ for all $p$.
B.2 Proof of Theorem 5.2
The proof of Theorem 5.2 relies on a characterization of Blackwell dominance that constructs a coupling between the two predictors, shown in Lemma B.3. The error is comparable under the coupled joint distribution, shown in Lemma B.4.
Lemma B.3 (Blackwell’s Informativeness Theorem).
Assume predictors $A$ and $B$ are both calibrated. Let $J^{A}$ be the joint distribution of $(r^{A},y)\in\Delta(\mathcal{Y})\times\mathcal{Y}$ for the prediction of predictor $A$ and the state. Similarly, $J^{B}$ for predictor $B$. The following three conditions are equivalent:
  1. 1.
    AA Blackwell dominates $B$;
  2. 2.
    There exists a coupling $\Pi$ of the the random variables $r^{A}$, $r^{B}$, and state $y$, such that: $r^{B}$ is conditionally independent of $y$ given $r^{A}$;
  3. 3.
    For the same coupling above, $\operatorname{\mathbb{E}}[r^{A}|r^{B}]=r^{A}$ holds almost surely.
Lemma B.4.
Let Cal be an error measure that is ex-ante truthful (Definition 2.12). Let $\Pi$ be a joint distribution of $(r^{A},r^{B},y)\in[0,1]\times[0,1]\times\{0,1\}$ such that the distribution of $(r^{A},y)$ is calibrated, and $r^{B}$ is independent of $y$ conditioned on $r^{A}$. For $T$ i.i.d. examples $(r^{A}_{1},r^{B}_{1},y_{1}),\ldots,(r^{A}_{T},r^{B}_{T},y_{T})$ drawn from $\Pi$, defining $\bm{r}^{A}:=(r^{A}_{1},\ldots,r^{A}_{T}),\bm{r}^{B}:=(r^{B}_{1},\ldots,r^{B}_{T}),\bm{y}:=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{r}^{A},\bm{y}^{A})]\leq\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{r}^{B},\bm{y}^{B})].$ (53)
Proof of Lemma B.4.
By our assumption that $r^{B}$ is independent of $y$ given $r^{A}$, and that the distribution of $(r^{A},y)$ is calibrated, we have
$\operatorname{\mathbb{E}}[y|r^{A},r^{B}]=\operatorname{\mathbb{E}}[y|r^{A}]=r^{A}$
Define feature $x:=(r^{A},r^{B})$. The equation above becomes $r^{A}(x)=\operatorname{\mathbb{E}}[y|x]$. This means that $r^{A}$ can be viewed as the ground-truth predictor. For $n$ i.i.d. examples $(r^{A}_{1},r^{B}_{1},y_{1}),\ldots,(r^{A}_{n},r^{B}_{n},y_{n})$ drawn from $\Pi$, define $x_{t}:=(r^{A}_{t},r^{B}_{t})$ for $t=1,\ldots,T$ and define $S:=((r^{A}_{t},r^{B}_{t}))_{t\in[T]}$. By the truthfulness of Cal, we have
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{A})]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{B})],$
which is equivalent to (53). ∎
Proof of Theorem 5.2.
By Blackwell’s informativeness theorem (Lemma B.3), there exists a coupling between the two predictors such that $r^{\prime}(x)$ is conditionally independent of $y$ given $r(x)$. Thus, the proof is completed by Lemma B.4. ∎
B.3 Proof of Theorem 5.3
Proof of Theorem 5.3.
By Lemma B.4, it suffices to show that $v$ and $y$ are conditionally independent given $\widehat{v}$. Since $y\in\{0,1\}$ is binary, it suffices to show that $\operatorname{\mathbb{E}}[y|v,\widehat{v}$] is determined by $\widehat{v}$ alone and does not additionally depend on $v$, which is indeed true:
$\operatorname{\mathbb{E}}[y|v,\widehat{v}]=\operatorname{\mathbb{E}}[y|v]=\widehat{v}$$\qed$
Appendix C $\mathrm{ATB}$ and Brier Loss
For ATB as a special case of UBSE, we have the following stronger result than Lemma 3.6:
Lemma C.1 (ATB and Brier loss).
Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$. For $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ drawn i.i.d. from $J$, defining $\bm{v}:=(v_{1},\ldots,v_{T}),\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\mathrm{ATB}(\bm{v},\bm{y})]\geq\frac{1}{T}\operatorname{\mathbb{E}}_{(v,y)\sim J}[(v-y)^{2}].$ (54)
The inequality becomes an equality if $J$ is perfectly calibrated.
Proof.
For every fixed threshold $q\in[0,1]$, we have
$\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}<q}(v_{t}-y_{t})\right)^{2}\right]$ $=\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}(v_{t}-y_{t})\mathbb{I}\left[v_{t}<q\right]\right)^{2}\right]$
$=T\operatorname{\mathbb{E}}_{J}[((v-y)\mathbb{I}\left[v<q\right])^{2}]+T(T-1)\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]^{2}$ (because $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ are drawn i.i.d. from $J$)
$\geq T\operatorname{\mathbb{E}}_{J}[((v-y)\mathbb{I}\left[v<q\right])^{2}]$
$=T\operatorname{\mathbb{E}}_{J}[(v-y)^{2}\mathbb{I}\left[v<q\right]].$
Similarly,
$\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}\geq q}(v_{t}-y_{t})\right)^{2}\right]\geq T\operatorname{\mathbb{E}}_{J}[(v-y)^{2}\mathbb{I}\left[v\geq q\right]].$
Summing up the two inequalities above, for every $q\in[0,1]$ we have
$\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}<q}(v_{t}-y_{t})\right)^{2}+\left(\sum_{t:v_{t}\geq q}(v_{t}-y_{t})\right)^{2}\right]\geq\frac{1}{T}\operatorname{\mathbb{E}}_{J}[(v-y)^{2}].$
Taking expectation over $q\sim\mathrm{Unif}([0,1])$ proves (54). When $J$ is perfectly calibrated, all inequalities in this proof become equalities, so (54) also holds as an equality. ∎
Appendix D Standard Uniform Convergence Bounds
We include some standard notions and results on concentration inequalities and sample complexity bounds for uniform convergence. They are used when we prove the sample complexity bounds for estimating $\mathrm{ATB}$ and $\ell_{1}\textup{-}\mathrm{ATB}$ in Theorem 4.10.
We start with the definition of the Rademacher complexity.
Definition D.1 (Rademacher complexity).
Let $F$ be a family of real-valued functions $f:Z\to\mathbb{R}$ on some domain $Z$. Given $z_{1},\ldots,z_{n}\in Z$, we define the Rademacher complexity as follows:
$\mathcal{R}(F;z_{1,\ldots,n}):=\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right],$
where the expectation is over $s_{1},\ldots,s_{n}$ drawn uniformly at random from $\{-1,1\}^{n}$.
The following theorem is a standard application of the Rademacher complexity for proving uniform convergence bounds.
Proposition D.2 (Uniform convergence from Rademacher complexity).
Let $F$ be a family of functions $f:Z\to[a,b]$ on some domain $Z$ and with range bounded in $[a,b]$. Let $\Gamma$ be an arbitrary distribution over $Z$. Then for $n$ i.i.d. examples $z_{1},\ldots,z_{n}$ from $\Gamma$,
$\operatorname{\mathbb{E}}_{z_{1,\ldots,n}}\left[\sup_{f\in F}\left|\frac{1}{n}\sum_{i=1}^{n}f(z_{i})-\operatorname{\mathbb{E}}_{z\sim\Gamma}[f(z)]\right|\right]\leq 2\operatorname{\mathbb{E}}_{z_{1,\ldots,n}}[\mathcal{R}(F;z_{1,\ldots,n})].$
Moreover, for any $\delta\in(0,\frac{1}{3})$ and $n\in\mathbb{N}$, with probability at least $1-\delta$ over the random draw of $n$ i.i.d. examples $z_{1},\ldots,z_{n}$ from $\Gamma$, it holds that
$\sup_{f\in F}\left|\frac{1}{n}\sum_{i=1}^{n}f(z_{i})-\operatorname{\mathbb{E}}_{z\sim\Gamma}[f(z)]\right|\leq 2\mathcal{R}(F;z_{1,\ldots,n})+O\left((b-a)\sqrt{\frac{\log(1/\delta)}{n}}\right).$
Proposition D.3 (Rademacher Complexity after Lipschitz Postprocessing).
Let $F$ be a family of functions $f:Z\to\mathbb{R}$. For $i=1,\ldots,n$, let $z_{i}\in Z$ be an element of the domain $Z$ and let $\eta_{i}:\mathbb{R}\to\mathbb{R}$ be any $1$-Lipschitz function. It holds that
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}\eta_{i}(f(z_{i}))\right]\leq\mathcal{R}(F;z_{1,\ldots,n})=\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right].$
Proof.
By induction, it suffices to consider the case where all the $\eta_{i}$’s are the identity function except $\eta_{1}$.
Now we have
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}\eta_{i}(f(z_{i}))\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f\in F}\left(s_{1}\eta_{1}(f(z_{1}))+\sum_{i=2}^{n}s_{i}f(z_{i})\right)+\sup_{f\in F}\left(-s_{1}\eta_{1}(f(z_{1}))+\sum_{i=2}^{n}s_{i}f(z_{i})\right)\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(|\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))|+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right].$ (55)
Similarly,
$\operatorname{\mathbb{E}}\left[\sup_{f\in F}\frac{1}{n}\sum_{i=1}^{n}s_{i}f(z_{i})\right]$
$={}$ $\frac{1}{2n}\operatorname{\mathbb{E}}\left[\sup_{f_{+},f_{-}\in F}\left(|f_{+}(z_{1})-f_{-}(z_{1})|+\sum_{i=2}^{n}s_{i}(f_{+}(z_{i})+f_{-}(z_{i}))\right)\right].$ (56)
By the 11-Lipschitz property of $\eta_{1}$, we have
$|\eta_{1}(f_{+}(z_{1}))-\eta_{1}(f_{-}(z_{1}))|\leq|f_{+}(z_{1})-f_{-}(z_{1})|.$
This implies that (55) is a lower bound of (56), completing the proof. ∎
The following is the standard definition of the VC dimention for binary function families:
Definition D.4 (VC Dimension (Vapnik and Chervonenkis, 1971)).
The VC dimension of a family $F$ of binary functions $f:Z\to\{0,1\}$ is the largest size $d$ of a subset $Z^{\prime}=\{z_{1},\ldots,z_{d}\}\subseteq Z$ such that for each of the $2^{d}$ choices of $\bm{s}:=(s_{1},\ldots,s_{d})\in\{0,1\}^{n}$, there exists $f_{\bm{s}}\in F$ such that $f_{\bm{s}}(z_{i})=s_{i}$ for every $i=1,\ldots,d$.
The following standard result can be proved using Dudley’s chaining argument (see e.g. Theorem 8.3.23 of Vershynin (2018)):
Proposition D.5 (Rademacher Complexity from VC Dimension).
Let $F$ be a family of binary functions $f:Z\to\{0,1\}$ with VC dimension at most $d$. Then for any $n\in\mathbb{Z}_{>0}$ and any $z_{1},\ldots,z_{n}\in Z$, we have
$\mathcal{R}(F;z_{1,\ldots,n})\leq O\left(\sqrt{\frac{d}{n}}\right).$